### v1

Abstract
Modern multi-layer networks are commonly stored and analyzed in a local and distributed fashion because of the privacy, ownership, and communication costs. The literature on the model-based statistical methods for community detection based on these data is still limited. This paper proposes a new method for consensus community detection and estimation in a multi-layer stochastic block model using locally stored and computed network data with privacy protection. A novel algorithm named privacy-preserving Distributed Spectral Clustering (ppDSC) is developed. To preserve the edges’ privacy, we adopt the randomized response (RR) mechanism to perturb the network edges, which satisfies the strong notion of differential privacy. The ppDSC algorithm is performed on the squared RR-perturbed adjacency matrices to prevent possible cancellation of communities among different layers. To remove the bias incurred by RR and the squared network matrices, we develop a two-step bias-adjustment procedure. Then we perform eigen-decomposition on the debiased matrices, aggregation of the local eigenvectors using an orthogonal Procrustes transformation, and $k$-means clustering. We provide theoretical analysis on the statistical errors of ppDSC in terms of eigen-vector estimation. In addition, the blessings and curses of network heterogeneity are well-explained by our bounds.
Keywords: Multi-layer stochastic block model, Distributed SVD, Bias-correction, Differential privacy
1 Introduction
Advances in modern technology have facilitated the collection of multiple network data from various sources (e.g., Facebook, LinkedIn, and Twitter) that present disparate social ties (Kairouz et al., 2021). This data type can be well represented by multi-layer networks (Mucha et al., 2010; Boccaletti et al., 2014; Kivelä et al., 2014; Han et al., 2015), in which the nodes are the entities of interest, and the network layers reflect their relationships across different sources. As a result, the study of multi-layer networks has received remarkable attention recently. Most existing works assume that all network layers are stored in a single central machine, see, e.g., Paul and Chen (2016); Bhattacharyya and Chatterjee (2018); Liu et al. (2018); Paul and Chen (2020); Lei et al. (2020); Arroyo et al. (2021); MacDonald et al. (2022); Lei and Lin (2022); Huang et al. (2022). However, in practice, the multi-layer network is often locally stored because of the private ownership and the communication cost among others (Dwork et al., 2014; Kairouz et al., 2021). For example, consider a multi-layer friendship network obtained from different social platforms, where the nodes are the common users of these platforms, and the layers correspond to different social ties. In this network, neither the users would like others to know what kind of users they are following completely, nor are the platforms willing to share all the network information because of the ownership. Thus, aggregating all network layers into one central machine would be rather difficult in this scenario. To this end, the analysis of the multi-layer networks should be naturally based on a distributed and federated architecture.
In the literature, the multi-layer stochastic block model (multi-layer SBM) (Han et al., 2015; Paul and Chen, 2016; Barbillon et al., 2017) has been popularly used for community detection in multi-layer networks, in which each layer is assumed to share the same unobserved community membership (Greene and Cunningham, 2013; Liu et al., 2013). There are two major methodological challenges to apply the multi-layer SBM for locally stored networks with privacy-preserving constraints. First, because of the privacy budget, the randomized response (RR) mechanism (Warner, 1965; Karwa et al., 2017; Ayed et al., 2020) that simply flips all edges in the network with a certain probability is often adopted to achieve differential privacy (DP) (Dwork et al., 2006, 2014; Mironov, 2017; Dong et al., 2022; Bi and Shen, 2022). Although this method can greatly protect an individual’s privacy by ensuring that no individual significantly influences the algorithm’s output, the RR mechanism would bring bias into the multi-layer SBM. The existing community detection methods (e.g., Paul and Chen (2020); Arroyo et al. (2021); Lei and Lin (2022); Huang et al. (2022)) proposed for the exact multi-layer SBMs cannot be directly applied in this scenario. Therefore, a bias-correction procedure is needed for community detection and estimation in the composite and RR-perturbed multi-layer SBM.
Second, directly sending the RR-perturbed networks, either with or without bias-correction on the local machines to the central server, can bring large communication cost or storage cost on the server, especially for large networks. When the bias-correction is achieved on the local machines, the networks become dense; when it is performed on the central server, it needs to preserve all networks. Moreover, sending non-debiased sparse networks can be also communication expensive when the network is large. Therefore, a communication-efficient algorithm is of importance in our considered setting.
Motivated by the above problems, we propose a privacy-preserving Distributed Spectral Clustering (ppDSC) algorithm for detecting the communities of locally stored multi-layer networks with private preserving. To avoid possible cancellation of the communities among different layers, we adopt the squared adjacency matrix (Lei and Lin, 2022) as the input of ppDSC. Our proposed algorithm consists of four steps given below.
  • Step 1: we conduct RR on the network located in each local machine to achieve DP.
  • Step 2: we perform bias-correction on each local RR-perturbed network and then obtain a bias-corrected network on each machine.
  • Step 3: we perform eigen-decomposition on each bias-corrected network and send the top-eigenvectors of each network to the central server. Due to the orthogonal ambiguity of subspaces, we average the orthogonal transformed local eigenvectors using the orthogonal Procrustes transformation (OPT) (Schönemann, 1966; Cape, 2020; Charisopoulos et al., 2021) on the central server. This method ensures that the transformed local eigen-vectors are aligned with each other.
  • Step 4: we implement $k$-means clustering on the orthogonalized vector on the central server to obtain the estimated communities.
We provide rigorous non-asymptotic analysis of the statistical errors of ppDSC in terms of eigen-vector estimation and clustering. It turns out that both errors consist of three parts: the squared local, heterogeneity-induced and centralized errors. The privacy effect is inherent in the first and third parts.
1.1 Our contributions and related work
The contribution of this paper is multi-fold, summarized as follows.
First, we propose a novel community detection method for multiple networks collected from different sources with privacy concerns. Our algorithm satisfies the strong notion of DP. For network data, two notions of DP are often used (Rastogi et al., 2009); one is edge-DP, considered in our paper; the other one is node-DP, which conceals the inclusion or removal of a particular node with all its adjacent edges. Under the framework of these two notions, DP algorithms have been developed for several problems, including graph statistics releasing (Nissim et al., 2007; Karwa et al., 2011; Imola et al., 2021), synthetic graphs releasing (Qin et al., 2017), and generative graph models estimation (Mir and Wright, 2012; Borgs et al., 2015; Karwa and Slavković, 2016; Chang et al., 2021), among others. However, there are few studies on privacy-preserving community detection using statistical models, e.g., Nguyen et al. (2016); Ji et al. (2019); Hehir et al. (2021); Seif et al. (2022). In particular, to the best of our knowledge, privacy-preserving community detection for multi-layer networks has not been studied yet.
Second, we develop a bias-correction procedure to remove the bias brought by RR in the multi-layer SBM, and theoretically investigate the effect of the bias-correction on community detection. The previous works focused on deriving the information-theoretical limits (Seif et al., 2022) and misclassification rate of community detection in bias-corrected RR-perturbed SBMs (Hehir et al., 2021). However, how the bias-correction procedure can affect the clustering and estimation accuracy is still unclear, which is studied in our paper through rigorous theoretical justification.
Third, we develop a communication-efficient algorithm, ppDSC, by which we only need to send the top eigenvectors from the local machines to the central server. The method is inspired by the distributed PCA proposed for other scenarios in the divide-and-conquer paradigm (Fan et al., 2019; Bhaskara and Wijewardena, 2019; Charisopoulos et al., 2021) and iterative paradigm (Garber et al., 2017; Gang et al., 2019; Chen et al., 2021a; Li et al., 2021; Guo et al., 2021). However, we incorporate the perturbation and bias-correction procedure in our algorithm; as a result, it poses great challenges to derive the theories, including the estimation error bounds.
Last but not least, we provide a series of novel bounds and explicitly show how the bounds depend on the privacy parameters. These theoretical results have not yet been established. In the special case without the privacy and local computation demand, the error bounds of our algorithm reduce to that given in Lei and Lin (2022). The bounds also review the effect of RR on community detection. In particular, adding more network edges can lead to larger error bounds than removing edges. Our established bounds can also help to explain the blessings and curses of heterogeneity among the local networks.
1.2 Organizations
The rest of the paper is organized as follows. Section 2 introduces the preliminaries of the model and DP. Section 3 develops the algorithm for privacy-preserving community detection of locally stored networks. Section 4 studies the theoretical properties of ppDSC. Section 5 and 6 include the simulation and real data experiments. Section 7 concludes the paper.
2 Problem formulation
In this section, we first present the multi-layer SBMs for modeling the locally stored networks. Next, we introduce the DP and the RR mechanism for privacy protection.
2.1 The model for locally stored networks
Suppose that $L$ aligned networks with $n$ common nodes are locally stored in $L$ machines, respectively. The networks are binary and symmetric with the adjacency matrix denoted by $A_{l}\in\{0,1\}^{n\times n},l=1,...,L$. We consider the scenario that $A_{l}$’s share common communities but with possibly different edge densities.
To model the above locally stored networks, we assume that $A_{l}$’s are independently generated from the following multi-layer SBMs. The $n$ nodes are assigned to $K$ non-overlapping communities with the community assignment of node $i$ denoted by $g_{i}\in\{1,...,K\}$. Given the community assignments, each entry $A_{l,ij}(i<j)$ of $A_{l}$ is generated independently according to
$A_{l,ij}\sim{\mathrm{Bernoulli}}(B_{l,g_{i}g_{j}}),\quad i<j,\;l=1,...,L,$ (1)
where $B_{l}$’s are link probability matrices, and the diagonal entries $A_{l,ii}$’s are all 0. The heterogeneity among the networks is inherent in $B_{l}$’s, and we will quantify it in the next section. Note that we assume that $A_{l}$’s are generated locally at the machine $L$’s.
Model (1) is the multi-layer SBMs studied in Han et al. (2015); Paul and Chen (2016) and so forth. When the adjacency matrices $A_{l}$’s are generated and stored in the local machines, we consider DP and the RR mechanism for the adjacency matrices for private protection.
2.2 Differential privacy and randomized response
Though the local networks $A_{l}$’s are not directly shared by all the local machines, potential privacy breaches still exist (Dwork et al., 2017). DP provides a rigorous framework for privacy-preserving data analysis (Dwork et al., 2006). Generally, a DP algorithm pursues that its output appears similar in probability if one individual changes the input data. By definition, DP protects any individual’s privacy from an adversary who has access to the algorithm output and even sometimes the rest of the data. In addition, it can prevent other kind of attacks, say, the re-identification attack, the tracing attack, and the membership attack, among others; see Dwork et al. (2017) for a survey. In the context of networks, we consider the following notion of edge-DP (Rastogi et al., 2009; Karwa et al., 2017).
Definition 1 (Edge-DP)
Let $A\in\{0,1\}^{n\times n}$ be the network whose edges are sensitive. For any range $\mathcal{R}$, let $\mathcal{Q}(A)\in\mathcal{R}$ be a randomized mechanism that generates some synthetic information from $A$, indicated by a known parameter $\omega$. Then, $\mathcal{Q}(A)$ is said to be $\epsilon$-edge-DP if
$\underset{{\tilde{A}}\in\mathcal{R}}{\max}\,\underset{A^{\prime},A,\,\Delta(A,A^{\prime})=1}{\max}\;\log\frac{\mathbb{P}_{\omega}(\mathcal{Q}(A)={\tilde{A}}\,|\,A)}{\mathbb{P}_{\omega}(\mathcal{Q}(A^{\prime})={\tilde{A}}\,|\,A^{\prime})}\leq\epsilon,$ (2)
where $\Delta(A,A^{\prime})=1$ means that $A$ and $A^{\prime}$ are neighboring by differing only one edge.
Remark 1
We make the following remarks on the DP.
Edge-DP protects the privacy of network edges by ensuring that the algorithm’s output does not reveal the inclusion or removal of any certain edge in the network. In other words, edge-DP can protect against an adversary who knows the algorithm’s output and the existence or not of edges between all pairs but one pair of nodes.
The $\epsilon$ is often termed as the privacy budget. A smaller privacy budget indicates stronger protection of privacy. Note that DP is the worst-case definition, and the randomness in (2) only comes from the DP mechanism $\mathcal{Q}$.
Like the general DP, edge-DP enjoys the post-processing property (Dwork et al., 2014), that is, an $\epsilon$-edge-DP algorithm is still $\epsilon$-edge-DP after any post process that does not use the knowledge about the original network.
Edge-DP could be achieved by flipping the network edges with certain probabilities, which is called the randomized response (RR) mechanism (Warner, 1965; Karwa et al., 2017; Ayed et al., 2020). The algorithm outputs a perturbed adjacency matrix $\tilde{A}$ of $A$. Specifically, for each $i<j$, when $A_{ij}=1$, $\tilde{A}_{ij}$ keeps to be 1 with probability $q$ and flips to 0 otherwise. Similarly, when $A_{ij}=0$, ${\tilde{A}}_{ij}$ remains to be 0 with probability $q^{\prime}$ and changes to 1 otherwise. The diagonal entries ${\tilde{A}}_{ii}=0$. Note that $q$ and $q^{\prime}$ can be published publicly. The relationship between $q,q^{\prime}$ and the privacy budget $\epsilon$ is shown in the following proposition.
Proposition 1 (Karwa et al. (2017))
The randomized response mechanism satisfies $\epsilon$-edge-DP with
$\epsilon\geq\log\max\left\{\frac{q^{\prime}}{1-q},\frac{1-q}{q^{\prime}},\frac{1-q^{\prime}}{q},\frac{q}{1-q^{\prime}}\right\}.$ (3)
According to (3), a typical choice of $q$ and $q^{\prime}$ is that
$q=q^{\prime}=\frac{\mathrm{e}^{\epsilon}}{1+\mathrm{e}^{\epsilon}},$ (4)
which has been proven to be optimal in maximizing the mutual information between raw and RR-perturbed data (Ayed et al., 2020).
Although RR can achieve edge-DP, one can imagine that the resulting network largely deviates from the original network. In the next section, we will see how we can debias the perturbed networks so as to improve the algorithm’s utility under the multi-layer SBMs.
3 Privacy-preserving community detection for locally distributed networks
This section proposes the privacy-preserving community detection algorithm for locally distributed networks. Before that, we specify the centralized estimator and develop a bias-adjustment procedure to remove the bias inherent in the RR-perturbed networks.
Notes and notation. We use $\lambda_{\min}(A)$ to denote the smallest non-zero eigenvalue of a given matrix $A$. Moreover, we use the projection distance $\mathrm{dist}(U,U^{\prime}):=\|UU^{\intercal}-U^{\prime}U^{\prime{\intercal}}\|_{2}$ to measure the subspace distance between any pair of orthogonal basis $U,U^{\prime}\in\mathbb{R}^{n\times K}$. $\|A\|_{2}$ denotes the spectral norm of matrix $A$ or the Euclidian norm of vector $A$ and $\|A\|_{\tiny F}$ denote the Frobenius norm of matrix $A$. In addition, we use the following standard notation for asymptotics. We write $f(n)\asymp g(n)$ or $f(n)=\Theta(g(n))$ if $cg(n)\leq f(n)\leq Cg(n)$ for some constants $0<c<C<\infty$, $f(n)\lesssim g(n)$ or $f(n)=O(g(n))$ if $f(n)\leq Cg(n)$ for some constant $C<\infty$, and $f(n)\gtrsim g(n)$ or $f(n)=\Omega(g(n))$ if $f(n)\geq cg(n)$ for some constant $c>0$. Finally, the constants throughout this work may be different from place to place.
Recall that there are $L$ local machines, with each owning a symmetric binary network adjacency matrix $A_{l},l=1,...,L$. $A_{l}$’s are assumed to be generated from a multi-layer SBM. Denote the population of $A_{l}$ by $P_{l}:=\Theta B_{l}\Theta^{\intercal}$, where $\Theta\in\mathbb{R}^{n\times K}$ is the membership matrix which is formed by forcing $\Theta_{ig_{i}}=1$ and $\Theta_{ij}=0$ for $j\neq g_{i}$. We aim to recover the common membership matrix $\Theta$ using the network information.
Without the privacy constraint, a common centralized approach is to perform community detection algorithms on $\sum_{l}A_{l}$ (Bhattacharyya and Chatterjee, 2018; Paul and Chen, 2020). However, direct summation of $A_{l}$’s would lead to information loss when $A_{l}$’s are heterogeneous in the sense of community pattern. In other words, if some $P_{l}$’s are positive semidefinite, which present homophily social ties and others are negative semidefinite, which indicate heterophily social ties (Kim and Leskovec, 2012), then adding them up would cancel out some community signals. To alleviate this, we consider the matrix
$M:=\frac{1}{Ln}\sum_{l=1}^{L}A_{l}^{2}$ (5)
studied in Lei and Lin (2022), which could accommodate different community patterns. Note that Lei and Lin (2022) further developed a bias-correction procedure to modify the diagonal elements of $M$. The centralized estimator is then obtained by performing spectral clustering on $M$. Specifically, the eigen-decomposition is first computed on $M$ to obtain the first $K$ eigenvectors, and then $k$-means is conducted on the resulting eigenvectors to obtain $K$ communities. Hopefully, $M$ can approximate its population counterpart
${Q}:=\frac{1}{L}\sum_{l=1}^{L}\frac{P_{l}^{2}}{n}:=\frac{1}{L}\sum_{l=1}^{L}{Q}_{l},$ (6)
whose eigenvector reveals the true communities in the sense that two rows of the eigenvector matrix are identical if and only if the corresponding nodes are in the same underlying communities (Rohe et al., 2011; Lei and Rinaldo, 2015; Zhang et al., 2022).
However, with the privacy constraint, we could only use the RR-perturbed adjacency matrix instead of $A_{l}$ in (5), which is a biased estimator of $P_{l}$. In the next subsection, we will develop a bias-adjustment procedure to obtain a good estimator for $Q$.
3.1 Bias-adjustment
Our debiasing procedure involves two steps: in step 1, we build a debiased estimator for the link probability matrix upon the RR-perturbed adjacency matrix stored in each local machine; in step 2, we further remove the bias of the diagonals caused by taking the square of the debiased estimated matrix obtained in step 1.
Step 1. We now illustrate the necessity of bias-adjustment when dealing with RR-perturbed data.
For $1\leq l\leq L$, we denote the RR-perturbed $A_{l}$ by ${\tilde{A}}_{l}$. Specifically, for $i<j$, if $A_{l,ij}=1$,
${\tilde{A}}_{l,ij}=\begin{cases}1,\quad\mathrm{with\,probability}\,q,\\ 0,\quad\mathrm{otherwise},\end{cases}$ (7)
and ${\tilde{A}}_{l,ji}={\tilde{A}}_{l,ij}$. On the other hand, if $A_{l,ij}=0$,
[…]
and ${\tilde{A}}_{l,ji}={\tilde{A}}_{l,ij}$. By (7) and (8) and taking expectations for two times, we have
$\mathbb{E}({\tilde{A}}_{l,ij})=P_{l,ij}(q+q^{\prime}-1)+1-q^{\prime},$
which is biased from the true expectation $P_{l,ij}$ of $A_{l,ij}$. Therefore, for $i\neq j$, the local machine only needs to debias ${\tilde{A}}_{l,ij}$ to be
$\overline{A}_{l,ij}=\frac{{\tilde{A}}_{l,ij}-(1-q^{\prime})}{q+q^{\prime}-1}.$
We thus obtain the following debiased estimator $\overline{A}_{l}$ for $P_{l}$,
$\overline{A}_{l}:=[{\tilde{A}}_{l}-(1-q^{\prime})\mathbf{1}\mathbf{1}^{\intercal}+\mathrm{diag}(1-q^{\prime})]/(q+q^{\prime}-1),$ (9)
where $\mathbf{1}$ denotes the vector with all entries being 1.
Step 2. Though $\overline{A}_{l}$ is the unbiased estimator for $P_{l}$, the diagonal elements of $\overline{A}_{l}^{2}$ can cause bias for $P_{l}^{2}$. It turns out that the error would be smaller if we further subtract $\frac{\sum_{j}{\tilde{A}}_{l,ij}(q^{\prime})^{2}}{(q+q^{\prime}-1)^{2}}$ from the $i$th diagonal elements of $\overline{A}_{l}^{2}$ provided that $\max_{i,j,l}P_{l,ij}=o(1)$. In particular, this debiasing step is crucial when $L$ is increasing. For the sake of saving space, the details are relegated to Appendix A.
Combining the two steps of debiasing, we eventually obtain the following estimator $\overline{M}$ for $Q$ defined in (6),
$\overline{M}:=\frac{1}{L}\sum_{l=1}^{L}\overline{M}_{l}$ (10)
with
$\overline{M}_{l}:=\frac{\overline{A}_{l}^{2}}{n}-\frac{(q^{\prime})^{2}}{n(q+q^{\prime}-1)^{2}}G_{l},$ (11)
where $G_{l}$ is a diagonal matrix with $G_{l,ii}=\sum_{j}{\tilde{A}}_{l,ij}$, and $\overline{A_{l}}$ and ${\tilde{A}}_{l}$ are defined in (7)-(9). Note that the second step of debiasing is in a similar spirit of Lei and Lin (2022) except that their debiasing is performed on the original network matrix, while ours is on the debiased RR-perturbed network matrix.
Therefore, with the privacy constraint, the ideal centralized estimator is to perform spectral clustering on the debiased estimator $\overline{M}$. Noting that $\overline{M}_{l}$’s are separable across different local machines. We show in the next subsection how we can approximate the eigenvector of $\overline{M}$ with the local computation.
3.2 Privacy-preserving distributed spectral clustering
Since DP enjoys the post-processing property, in principle, the local machines can send the dense debiased estimators $\overline{M}_{l}$’s to the central server where the spectral clustering is further performed on the aggregated matrix $\overline{M}$. However, both the communication and computational cost can be very high for sending the whole dense matrices $\overline{M}_{l}$’s and performing the aggregation of these matrices on the central server, especially when the matrix dimensions are large. Instead, we consider a more efficient distributed algorithm described as follows.
We adopt the one-shot averaging framework in Charisopoulos et al. (2021) for distributed computing performed on our debiased RR-perturbed network matrices. That is, on the $l$th machine, we first compute the eigen-decomposition of $\overline{M}_{l}$, and obtain the matrix $\hat{V}_{l}\in\mathbb{R}^{n\times K}$ that consists of the first $K$ eigenvectors, for all $l=1,...,L$. On the central server, we average all the local solutions. Note that direct averaging is inferior due to the orthogonal ambiguity of the problem, i.e., for any orthogonal matrix $Z$, $\hat{V}_{l}Z$ also forms a valid orthonormal basis of the subspaces formed by $\hat{V}_{l}$. Hence, $\hat{V}_{l}$’s are not guaranteed to be close to each other, even if their corresponding subspaces are close. In light of this, the local $\hat{V}_{l}$’s must be aligned with each other before averaging.
To avoid the orthogonal ambiguity of the problem, we use the Procrustes fixing strategy (Schönemann, 1966; Cape, 2020; Charisopoulos et al., 2021; Guo et al., 2021). For two matrices $V_{1},V_{2}$ to be aligned, we can map $V_{2}$ to $V_{1}$ using an orthogonal transformation solved by the following optimization,
$Z:=\operatorname{arg\,min}_{Z\in\mathcal{O}_{K}}\|V_{1}-V_{2}Z\|_{F},$ (12)
where $\mathcal{O}_{k}$ denotes the set of $K\times K$ orthogonal matrices. Note that (12) has a closed-form solution. Let the SVD of $V_{2}^{\intercal}V_{1}$ be $LDR^{\intercal}$, then $Z:=LR^{\intercal}$ is the solution to (12). In such a way, we can align local eigenspaces $\hat{V}_{l}$ to an arbitrarily chosen reference eigenspace, for example, $\hat{V}_{o}$, after which the aligned eigenspaces can then be averaged and orthogonalized.
The proposed privacy-preserving Distributed Spectral Clustering (ppDSC) algorithm is summarized in Algorithm 1. It should be noted that, in a real scenario, the computation on the local machine may be fulfilled by a single party or cooperated by different parties. For example, a party with data ownership may desire to obtain good utility while preserving the data’s privacy. Thus all the local steps are conducted within the party. It can also happen that the RR perturbation is conducted by the data collection department of a company, while the data analysis department could do the remaining bias-correction and eigen-decomposition steps.
Algorithm 1 can achieve strong privacy protection in the DP’s framework.
Theorem 1
Define $A_{\mathrm{all}}:=(A_{1},...,A_{L})$. Denote the output of Algorithm 1 by $\mathcal{M}\in\mathcal{R}$. For $\epsilon$ satisfies (3), Algorithm 1 is $\epsilon$-edge-DP in the following sense
$\underset{\mathcal{M}\in\mathcal{R}}{\max}\,\underset{A_{\mathrm{all}},A_{\mathrm{all}}^{\prime},\ \Delta(A_{\mathrm{all}},A_{\mathrm{all}}^{\prime})=1}{\max}\;\log\frac{\mathbb{P}(\mathcal{M}(A_{\mathrm{all}})\,|\,A_{\mathrm{all}})}{\mathbb{P}(\mathcal{M}(A_{\mathrm{all}}^{\prime})\,|\,A_{\mathrm{all}}^{\prime})}\leq\epsilon,$
where $\Delta(A_{\mathrm{all}},A_{\mathrm{all}}^{\prime})=1$ means that all layers of the multi-layer network pairs $A_{\mathrm{all}}$ and $A_{\mathrm{all}}^{\prime}$ are equal except one edge of one layer.
Algorithm 1 Privacy-preserving distributed spectral clustering (ppDSC)
1:  Input: Network adjacency matrices $\{A_{l}\}_{l=1}^{L}$ owned by local machines, number of clusters $K$, privacy parameters $q,q^{\prime}$, index of reference solution $o$.
2:  The local machines do:
3:  for $l=1$ to $L$ do
4:     RR-perturbation: Perturb $A_{l}$ to obtain ${\tilde{A}}_{l}$ using (7) and (8);
5:     Bias-correction: Constructs the debiased estimator $\overline{M}_{l}$ according to (11);
6:     Eigen-decomposition: Performs eigen-decomposition on $\overline{M}_{l}$ and denote the resulting eigen-spaces as $\hat{V}_{l}$.
7:  end for
8:  The central server does:
9:  Procrustes fixing: $\tilde{V}_{l}:=\hat{V}_{l}Z_{l}$ with $Z_{l}:=\operatorname{arg\,min}_{Z\in\mathcal{O}_{K}}\|\hat{V}_{l}Z-\hat{V}_{o}\|_{F}$ for each $l=1,...,L$;
10:  Aggregation: Aggregates the local solutions according to $\bar{V}:=\frac{1}{L}\sum_{l=1}^{L}\tilde{V}_{l}$;
11:  Orthogonalization: Performs the QR decomposition to obtain $\tilde{V},\tilde{R}={\rm qr}(\bar{V})$.
12:  Clustering: Conduct $k$-means clustering on $\tilde{V}$ with cluster number being $K$. The resulting clusters are formulated by the membership matrix $\tilde{\Theta}$.
13:  Output: The approximated eigen-space $\tilde{V}$ and the estimated membership matrix $\tilde{\Theta}$.
4 Theoretical analysis
In this section, we theoretically study the ppDSC algorithm in terms of the estimation accuracy of the eigenvectors and the misclassification rate.
To fix the ideas, we now recall and introduce some notations. We assume that there are $L$ local networks whose adjacency matrices $A_{l}\in\{0,1\}^{n\times n}$’s are generated from SBMs with layer-specific link probability matrices $B_{l}$’s and a consensus membership matrix $\Theta\in\mathbb{R}^{n\times K}$. We here assume that the number of communities $K$ is known. When $K$ is unknown, we will investigate the method and theory for the estimation of $K$, which is a model selection problem, as a future research topic, as this current work focuses on estimation and clustering analysis for a given model based on RR-perturbed network matrices. The number of nodes within community $k$ is denoted by $n_{k}$. The privacy parameters are denoted by $q,q^{\prime}$ ($1/2<q,q^{\prime}\leq 1$); see also (7) and (8). Our goal is to use ppDSC (Algorithm 1) to recover the first $K$ eigenvectors $V$ of the underlying target matrix $Q$ defined in (6) and the membership matrix $\Theta$.
The following assumptions are generally required to derive the theoretical results for community detection.
Assumption 1 (Balanced communities)
The community is balanced such that for any $k\in\{1,...,K\}$,
$c\frac{n}{K}\leq n_{k}\leq C\frac{n}{K},$
for some constants $0\leq c\leq C$.
Assumption 2 (Sparsity)
The link probability matrix $B_{l}=\rho B_{l,0}$, where $\rho=o(1)$ and $B_{l,0}$ is a $K\times K$ symmetric matrix with constant entries in $[0,1]$.
Assumption 3 (Full rank)
$\sum_{l}B_{l,0}^{2}$ is of full rank $K$ with $\lambda_{{\min}}(\sum_{l}B_{l,0}^{2})\geq cL$.
Assumption 4 (Limited heterogeneity)
The heterogeneity among the link probability matrices $B_{l}$’s are limited as
$\max_{l}\|B_{l}\Delta^{2}B_{l}-\frac{1}{L}\sum_{l=1}^{L}B_{l}\Delta^{2}B_{l}\|_{2}\leq\mathcal{H}(n,\rho,L).$
with $\Delta:={\operatorname{diag}}(\sqrt{n_{1}},...,\sqrt{n_{K}})$.
Remark 2
Assumption 1 requires that the communities are balanced. Assumption 2 specifies the link probability matrices. In particular, we require that the network sparsity $\rho=o(1)$. Assumption 3 requires the linear growth of the aggregated squared link probability matrices in terms of its minimum eigenvalue, which is trivially satisfied when $B_{l,0}$’s are equal and is also used in Lei and Lin (2022). Assumption 3 combining with Assumptions 2 and 1 implies that $\sum_{l=1}^{L}B_{l}\Delta^{2}B_{l}$ is of full rank, and $Q$ is of rank $K$; see the proof of Corollary 1. Indeed, the full rank requirement could be extended to the rank-deficient case; see Zhang et al. (2022) for example. Assumption 4 bounds the maximum deviation of $B_{l}\Delta^{2}B_{l}$’s from its average by $\mathcal{H}(n,\rho,L)$. For simplicity, we call $\mathcal{H}$ the measure for the heterogeneity of locally stored networks.
We first provide the following error decomposition theorem to present the general error bound for the output eigen-space of ppDSC.
Theorem 2 (Error Decomposition)
Suppose Assumptions 1 and 4 hold, the output $\tilde{V}$ of Algorithm 1 satisfies
$\mathrm{dist}(\tilde{V},V)\lesssim{\max_{l}\frac{\|\overline{M}_{l}-Q_{l}\|_{2}^{2}}{\lambda_{{\min}}^{2}(Q)}}+{\frac{\mathcal{H}^{2}(n,\rho,L)}{\lambda_{{\min}}^{2}(Q)}}+\frac{\|\sum_{l}\overline{M}_{l}/L-Q\|_{2}}{\lambda_{{\min}}(Q)}:=\mathcal{I}_{1}+\mathcal{I}_{2}+\mathcal{I}_{3}.$
provided that both $\mathcal{I}_{1}$ and $\mathcal{I}_{2}$ are smaller than a small constant $c$.
Remark 3
The condition that $\mathcal{I}_{1}$ and $\mathcal{I}_{2}$ are smaller than a small constant $c$ is a standard condition (Damle and Sun, 2020; Charisopoulos et al., 2021); it ensures that the maximum local error $\max\|\overline{M}_{l}-Q\|_{2}$ is smaller than the eigen-gap $\lambda_{{\min}}(Q)$, which further indicates that the $K$-largest eigenvalues of $\overline{M}_{l}$ are separated from its remaining eigenvalues by Weyl’s inequality. In Corollary 1, we show that this condition is satisfied by our setting under mild conditions and further provide the explicit expressions for $\mathcal{I}_{1}$ and $\mathcal{I}_{2}$.
The decomposed three components $\mathcal{I}_{1},\mathcal{I}_{2},\mathcal{I}_{3}$ are explained as follows. $\mathcal{I}_{1}$ corresponds to the maximum squared error of the local estimators $\hat{V}_{l}$’s except that we use $Q$ in the denominator instead of $Q_{l}$; see the Davis-Kahan theorem on the perturbation of eigenvectors (Davis and Kahan (1970); see Lemma 3 in the Appendix). As a result, we only require the averaged population matrix $Q:=\sum Q_{l}/L$ be of rank $K$, and the local population matrix $Q_{l}$ can be rank deficient, which is one of the benefits of combining the local eigenvectors. $\mathcal{I}_{2}$ corresponds to the degree of heterogeneity. To distinguish, we call $\mathcal{H}$ the measure for heterogeneity and call $\mathcal{I}_{2}$ the measure for scaled heterogeneity. It is worth noting that the scaled heterogeneity $\mathcal{I}_{2}$ affects the theoretical accuracy of $\tilde{V}$ according to its magnitude compared to $\mathcal{I}_{1}$ and $\mathcal{I}_{3}$. Moreover, $\mathcal{I}_{3}$ corresponds to the error of the centralized estimator $\hat{V}:=\frac{1}{L}\sum_{l}\overline{M}_{l}$. $\mathcal{I}_{1}$ and $\mathcal{I}_{2}$ can be regarded as the price needed to pay for the distributed computation.
One of the contributions in Theorem 2 is that it explicitly explains how the heterogeneity $\mathcal{H}$’s effect could be a curse or blessing to $\mathcal{I}_{2}$ and possibly the final error bound. Note that $\mathcal{H}$ and $\lambda_{{\min}}(Q)$ jointly decide $\mathcal{I}_{2}$. Therefore, if the local $B_{l}$’s are noisy from the signal such that combining them could hide the true signal, i.e., $\mathcal{H}$ is large, and $\lambda_{{\min}}(Q)$ is small, then the heterogeneity would be a curse. On the other hand, if the local $B_{l}$’s are far from the signal but combining them would enhance the signal, i.e., $\mathcal{H}$ is large, but $\lambda_{{\min}}(Q)$ can also be large such that their ratio is small, then the heterogeneity would be a blessing. We refer to the two simulated examples in Appendix C for more illustrations on the aforementioned phenomena.
To make the bounds in Theorem 2 more specific, in the following we provide Theorem 3 which establishes the estimation bounds of $\|\sum_{l}\overline{M}_{l}/L-Q\|_{2}$. Note that we assume that the privacy parameters $q,q^{\prime}$ relate to the privacy budget $\epsilon$ via (4). For illustration, we interchange the use of $q,q^{\prime}$, and $\epsilon$. Note that our results also hold for $q=\frac{\mathrm{e}^{\epsilon_{1}}}{1+\mathrm{e}^{\epsilon_{1}}}$ and $q^{\prime}=\frac{\mathrm{e}^{\epsilon_{2}}}{1+\mathrm{e}^{\epsilon_{2}}}$ with $\epsilon:=\max\{\epsilon_{1},\epsilon_{2}\}$ and $\epsilon_{1}$ and $\epsilon_{2}$ being of the same order.
Theorem 3
Suppose Assumption 2 holds, and
$\epsilon\leq\alpha_{1}\cdot{\log}\,n\quad{and}\quad\log(L+n)\leq n^{\alpha_{2}}$ (13)
for some constants $10<\alpha_{1}<$ 1 and $0<\alpha_{2}\leq 1-\alpha_{1}$, the estimator $\overline{M}:=\frac{1}{L}\sum_{l}\overline{M}_{l}$ of $Q$ satisfies
$\small\|\overline{M}-Q\|_{2}\leq C\max$$\bigg{\{}\frac{(q\rho+1-q^{\prime})^{1/2}n^{1/2}\rho\log^{1/2}(L+n)}{(q+q^{\prime}-1)\sqrt{L}},$
$\quad\quad\quad\frac{(q\rho+1-q^{\prime})\log(L+n)}{(q+q^{\prime}-1)^{2}\sqrt{L}},\,\frac{((q+q^{\prime}-1)\rho+1-q^{\prime})^{2}}{(q+q^{\prime}-1)^{2}}\bigg{\}},$ (14)
with probability larger than $1-O((L+n)^{-\nu})$ for some constant $\nu>0$.
Theorem 3 provides the error bound for the centralized estimator $\sum_{l}\overline{M}_{l}/L$. The RHS of (3) indicates that the privacy parameters $q$ and $q^{\prime}$ affect the accuracy of the estimated eigenvectors. In addition, the influence of $q$ and $q^{\prime}$ on the final bound is not equivalent.
First, we observe that the privacy parameters $q,q^{\prime}$ appear in the numerators of the first two terms of (3) in the form $q\rho+1-q^{\prime}$. Considering that $\rho=o(1)$, we can see that a large level of $q^{\prime}$ is more desirable than a large level of $q$ for reducing the estimated error (3). In other words, perturbing 0 to 1 is more harmful than perturbing 1 to 0 for the model estimation. Second, when $\rho\leq C\frac{(q\rho+1-q^{\prime})^{1/2}}{(q+q^{\prime}-1)}\cdot\frac{\log^{1/2}(L+n)}{n^{1/2}}$, the RHS of (3) reduces to
$\max\left\{\frac{(q\rho+1-q^{\prime})\log(L+n)}{(q+q^{\prime}-1)^{2}\sqrt{L}},\,\frac{((q+q^{\prime}-1)\rho+1-q^{\prime})^{2}}{(q+q^{\prime}-1)^{2}}\right\},$
which matches the following bound without privacy-preserving up to log factors derived in Lei and Lin (2022),
$\max\left\{\frac{\rho\log^{1/2}(L+n)}{\sqrt{L}},\,\rho^{2}\right\},$ (15)
provided that $q$ and $q^{\prime}$ both approach to 1.
The derivation of the bound for the single matrix $\overline{M}_{l}$ is similar to that in Theorem 3. For the sake of space, we summarize the results in Theorem B.1 in the Appendix.
Now we are ready to specify the unknown terms in Theorem 2.
Corollary 1
Suppose Assumptions 1-4 and condition (13) hold. If
$\rho\gtrsim\frac{(q\rho+1-q^{\prime})^{1/2}}{q+q^{\prime}-1}\cdot\sqrt{\frac{\log(n)}{n}}\quad{and}\quad\mathcal{H}\lesssim n\rho^{2},$ (16)
then the output $\tilde{V}$ of Algorithm 1 satisfies
$\mathrm{dist}(\tilde{V},V)\lesssim\mathcal{I}_{1}+\mathcal{I}_{2}+\mathcal{I}_{3}$ (17)
with
$\mathcal{I}_{1}:=\frac{q\rho+1-q^{\prime}}{(q+q^{\prime}-1)^{2}}\cdot\frac{\log n}{n\rho^{2}},\quad\mathcal{I}_{2}:=\frac{\mathcal{H}^{2}(n,\rho,L)}{n^{2}\rho^{4}},$ (18)
and
$\small\mathcal{I}_{3}:=\max$$\bigg{\{}\frac{(q\rho+1-q^{\prime})^{1/2}\log^{1/2}(L+n)}{(q+q^{\prime}-1)n^{1/2}\rho\sqrt{L}},\frac{(q\rho+1-q^{\prime})\log(L+n)}{(q+q^{\prime}-1)^{2}n\rho^{2}\sqrt{L}},\,\frac{((q+q^{\prime}-1)\rho+1-q^{\prime})^{2}}{(q+q^{\prime}-1)^{2}n\rho^{2}}\bigg{\}},$ (19)
with probability larger than $1-O((L+n)^{-\nu})-O(Ln^{-\nu})$ for some constant $\nu>0$.
To make the upper bound (17) valid, $\mathcal{I}_{1}$ and $\mathcal{I}_{2}$ both should be small enough. To this end, the technical condition (16) is utilized to meet the requirement in Theorem 2. The results in Corollary 1 can be simplified in various ways. For example, if we further require $\rho\gtrsim\frac{(q\rho+1-q^{\prime})^{1/2}}{q+q^{\prime}-1}\cdot\sqrt{\frac{\log(L+n)}{n}}$, then the first term in $\mathcal{I}_{3}$ dominate the second one. And the bound would be further reduced to
$\frac{(q\rho+1-q^{\prime})^{1/2}\log^{1/2}(L+n)}{(q+q^{\prime}-1)n^{1/2}\rho\sqrt{L}}$ (20)
if $\mathcal{I}_{2}$ is suppressed by $\mathcal{I}_{1}$ and $\mathcal{I}_{3}$, and $L\lesssim\frac{(q+q^{\prime}-1)^{2}}{q\rho+1-q^{\prime}}\cdot\frac{n\rho^{2}\log(L+n)}{\log^{2}n}$. As expected, we see in (20) that more local networks (large $L$) and less stringent privacy-preserving demand (large $q$ and $q^{\prime}$) lead to a smaller projection distance.
With Corollary 1 at hand, it is convenient to derive the following misclassification error bound of ppDSC.
Theorem 4
Suppose the conditions in Corollary 1 all hold. Denote the set of misclassified nodes by $\mathcal{M}$. Then, with probability larger $1-O((L+n)^{-\nu})-O(Ln^{-\nu})$ for some constant $\nu>0$, Algorithm 1 has misclassification rate upper bounded as
$\frac{|\mathcal{M}|}{n}\leq(\mathcal{I}_{1}+\mathcal{I}_{2}+\mathcal{I}_{3})^{2}$
with $\mathcal{I}_{1},\mathcal{I}_{2},\mathcal{I}_{3}$ defined in (18) and (19).
Note that the misclassification error is just the squared error of the eigen-vectors $\tilde{V}$ in Corollary 1. Therefore, Theorem 4 can be similarly discussed and simplified. We omit the discussion for the sake of space.
Finally, we discuss the necessity of the bias-adjustment procedure in Appendix B. We investigate the deviation of the non-debiased matrix $\tilde{M}:=\frac{1}{L}\sum_{l}\frac{1}{n}\tilde{A}_{l}^{2}$ from population matrices $Q$ and $\tilde{Q}:=\frac{1}{L}\sum_{l}\left((q+q^{\prime}-1)P_{l}+(1-q^{\prime})\textbf{1}_{n}\textbf{1}_{n}^{\intercal}\right)^{2}/{n}$, respectively. It turns out that the bias-adjustment procedure can lead to a smaller error of eigenvector and misclassification, no matter which population matrix is used.
5 Simulations
In this section, we evaluate the finite sample performance of the proposed algorithm. We show how the privacy parameters, the number of networks, and the number of nodes affect the finite sample behavior of the resulting estimates. In Appendix C, we test the effect of the heterogeneity on the performance of the proposed algorithm.
Practical consideration.
To mimic the scenario considered in this work, we assume that all the generated network datasets are private and locally distributed. To preserve privacy, we assume that the data analyst can not use the original data directly but can only use the RR-perturbed networks. For the local computation, we generally adopt the distributed SVD framework (Charisopoulos et al., 2021) considered in this work.
Methods compared.
Our algorithm is denoted by ppDSC. We compare our algorithm, ppDSC, with the following algorithms.
  • ppDSC-1b: computed according to Algorithm 1 but only with one-step bias-correction (i.e., diagonal-debiasing) given in Lei and Lin (2022). Note that our bias-correction for the diagonal elements is not identical to Lei and Lin (2022);
  • ppDSC-2b: computed according to Algorithm 1 but without any bias-correction, which corresponds to Charisopoulos et al. (2021);
  • Oracle: a centralized algorithm. It is obtained from the spectral clustering based on the debiased sum of squared RR-perturbed matrices (i.e., $\overline{M}$ defined in (10)). Though violating the local computation constraint, this is used as a baseline algorithm.
Performance measures.
According to our theoretical analysis, we use the projection distance to measure the subspace distance between the estimated eigenvectors and the true eigenvectors and use the misclassification rate to measure the proportion of misclassified (up to permutations) nodes.
Network-generation.
The local networks are generated independently from multi-layer SBMs with the structure given in Section 3. Similar to Lei and Lin (2022), we fix $K=3$ with balanced communities and set $B_{l}:=0.8B^{(1)}$ for $l=1,...,\lfloor L/2\rfloor$ and $B_{l}:=0.6B^{(2)}$ for $l=\lfloor L/2\rfloor+1,...,L$ with
$\scriptsize B^{(1)}=W\begin{bmatrix}1.5&&\\ &0.2&\\ &&0.4\end{bmatrix}W^{\intercal}\approx\begin{bmatrix}0.62&0.22&0.46\\ 0.22&0.62&0.46\\ 0.46&0.46&0.85\end{bmatrix},$
and
$\scriptsize B^{(2)}=W\begin{bmatrix}1.5&&\\ &0.2&\\ &&-0.4\end{bmatrix}W^{\intercal}\approx\begin{bmatrix}0.22&0.62&0.46\\ 0.62&0.22&0.46\\ 0.46&0.46&0.85\end{bmatrix},$
where
$\scriptsize W=\begin{bmatrix}1/2&1/2&-\sqrt{2}/2\\ 1/2&1/2&\sqrt{2}/2\\ \sqrt{2}/2&-\sqrt{2}/2&0\end{bmatrix}.$
We evaluate the effect of $q,q^{\prime}$, $n$ and $L$ on the numerical performance. The parameter set-up is as follows.
  • Effect of $L$: $n=210$, $q=q^{\prime}=0.9$, and $L$ ranges from 2 to 20.
  • Effect of $n$: $L=10$, $q=q^{\prime}=0.9$, and $n$ ranges from 15 to 300.
  • Effect of $q,q^{\prime}$: $L=10$, $n=210$, and $q$ and $q^{\prime}$ are equal and range from 0.6 to 1.
Results.
The averaged results over 20 replications are displayed in Figure 1. We have the following observations. First, as expected, as $q,q^{\prime}$, $n$, and $L$ increase, the performance of all estimators improve. Second, our estimator ppDSC is superior to the biased ppDSC-1b and ppDSC-2b except when the privacy parameters $q,q^{\prime}$ become close to 1 and the sample size $n$ is large. It is because that ppDSC-1b and ppDSC-2b ignore the bias caused by privacy-preserving the networks and thus mis-specify the model. When $q,q^{\prime}$ become close to 1, the bias from privacy preservation does not matter much, and then ppDSC-1b performs similarly to ppDSC. Third, as with extra local computation demand, our method ppDSC is generally not better than the Oracle estimator. However, when $q,q^{\prime}$, $n$, and $L$ become large, they are getting close. This is consistent with our theoretical results.
(I) Projection distance
(a) Effect of $q,q^{\prime}$
(b) Effect of $n$
(c) Effect of $L$
(II) Misclassification rate
(d) Effect of $q,q^{\prime}$
(e) Effect of $n$
(f) Effect of $L$
Figure 1: Comparison of ppDSC, ppDSC-1b, ppDSC-2b and Oracle on the simulated data. The effect of the privacy parameters $q,q^{\prime}$, the number of nodes $n$, and the number of networks $L$ are shown, respectively. The projection distance and misclassification rate are evaluated.
6 Real data analysis
In this section, we analyze the AUCS dataset (Rossi and Magnani, 2015), which is a public dataset collected by the employees at the Department of Computer Science of Aarhus University, to demonstrate the utility of the proposed algorithm. Though privacy preservation is prominent nowadays, to the best of our knowledge, networks known to be noisy and perturbed by the RR mechanism are not currently available to the public. Hence, we regard the AUCS dataset as the original data without any perturbation. We will mimic the actual setup to see how the privacy and local computation affects the clustering performance of the proposed algorithm.
Data description.
There are 61 employees (out of the total number of 142) who decided to join the survey, consisting of professors, postdoctoral researchers, Ph.D. students, and administration staff. Five types of relationships among the employees are measured. Namely, current working relationships, repeated leisure activities, regularly eating lunch together, co-authorship of a publication, and friendship on Facebook. The first three relationships were collected via a questionnaire. All the respondents answered all questions indicating that the data was complete. And the last two relationships were acquired online via the DBLP database and Facebook. As a result, the original dataset consists of 5 undirected networks with 61 common nodes. Besides the relationships, the research group labels that 55 out of 61 employees are available. There are eight research groups ($K=8$) in total. Our analysis uses the subnetworks of 55 employees with known labels.
Evidence of communities and heterogeneity.
The topology of networks with respect to different relationships is shown in Figure 2 with colors indicating the research groups. It is clear to see that the network structures are quite different. To take a closer look at the community structures, we reorder the nodes according to their underlying research groups (i.e., nodes corresponding to the same research group are aligned next to each other). We further plot the corresponding network adjacency matrices in Figure 3, where the whole network (Figure 3(a)) corresponds to the network that contains the edges of all five relationships. The whole network shows conspicuous block structures in that nodes in the same research group have more links than those across different research groups. Hence, it is reliable to use the research group labels as the underlying community memberships. Distinct from the whole network, we see that the community structures of each network layer are heterogeneous in that some communities are not evident or even missing to different extents.
(a) Work
(b) Facebook
(c) Leisure
(d) Lunch
(e) Coauthor
Figure 2: A visualization of the AUCS network where vertex colors represent research groups. Each network corresponds to a specific relationship among the vertices.
Practical consideration.
Though the AUCS dataset is publicly available, we note that some information about individuals say the eating and working relationships are sometimes sensitive and private. In addition, different relationships may be investigated and owned by different parties. Therefore, we consider the setting that the networks corresponding to each kind of relationship are private and locally owned by different parties. We will see how the privacy and local computation constraints affect the clustering performance of the proposed algorithm.
Methods compared.
Consistent with the simulation experiments, we compare our method ppDSC with ppDSC-1b, ppDSC-2b and Oracle. The detailed description of each method can be found in Section 5. Generally, we study how the performance of each method alters with the network number $L$ and the privacy parameters $q$ and $q^{\prime}$. Regarding the effect of $L$, we mean that we use some of the five local networks to construct the estimator to see how the number of used networks affects the clustering performance. Regarding the effect of $q,q^{\prime}$, we consider three cases, including $q,q^{\prime}$ both varying, $q$ varying but $q^{\prime}$ being fixed, and $q^{\prime}$ varying being $q$ being fixed, where the latter two cases aim to verify the different effect of $q$ and $q^{\prime}$ indicated by the theory. The detailed setups are as follows.
  • Effect of $L$: $q=q^{\prime}=0.9$, and $L$ ranges from 2 to 5.
  • Effect of $q,q^{\prime}$: $L=5$, and $q$ and $q^{\prime}$ are equal and range from 0.6 to 1.
  • Effect of $q$: $L=5$, $q^{\prime}=0.95$, and $q$ ranges from 0.6 to 1.
  • Effect of $q^{\prime}$: $L=5$, $q=0.95$, and $q^{\prime}$ ranges from 0.6 to 1.
Results.
The averaged misclassification rates over 20 replications are displayed in Figure 4. We have the following observations. First, as $L$ increases, the misclassification rate of all four methods decreases, showing the benefits of aggregating the information of local networks. Note that when $L=5$, a small drop in terms of the clustering performance appears, partially because the fifth network (i.e., the coauthor network) is rather sparse and has a weak signal of communities. Second, as expected, as $q$ and $q^{\prime}$ increase, the clustering performance of all four methods becomes better. We note that the clustering performance with $q^{\prime}$ fixed to be 0.95 (See Figure 4(c)) is better than that with $q$ fixed to be 0.95 (See Figure 4(d)), which shows that large $q^{\prime}$ is more desirable than large $q$. In other words, perturbing 0 to 1 is more harmful than perturbing 1 to 0. This is consistent with the theoretical results. Third, consistent with the simulation results, the proposed method ppDSC is better than the biased method ppDSC-1b and ppDSC-2b and is close to the baseline method Oracle as $q$ and $q^{\prime}$ increase.
(a) Whole
(b) Work
(c) Facebook
(d) Leisure
(e) Lunch
(f) Coauthor
Figure 3: A visualization of the network adjacency matrices of the AUCS dataset. The nodes are permutated according to their underlying research groups. (a) corresponds to the network adjacency matrix that contains all the edges of the five relationships. (b)-(f) correspond to the adjacency matrices of the five relationships among vertices, respectively.
(a) Effect of $L$
(b) Effect of $q,q^{\prime}$
(c) Effect of $q$
(d) Effect of $q^{\prime}$
Figure 4: Comparison of ppDSC, ppDSC-1b, ppDSC-2b and Oracle on the AUCS network. The effect of the privacy parameters $q,q^{\prime}$ ($q,q^{\prime}$ synchronously vary, $q^{\prime}$ fixed but $q$ varies, $q$ fixed but $q^{\prime}$ varies), and the number of networks $L$ are shown, respectively. The misclassification rate with respect to the research group labels are evaluated.
7 Conclusion
Driven by practical data collection and the associated analytic challenges, we provide a powerful toolbox for community detection and estimation in multi-layer SBMs based on locally stored networks with privacy protection. In the toolbox, a new algorithm named ppDSC has been developed, and it is shown to be both communication-efficient and privacy-aware. We also propose a two-step bias-adjustment procedure to remove the biases caused by the RR perturbation and take the square of the private-preserved network matrices. The procedure is found to be effective both theoretically and numerically. Moreover, we have rigorously investigated the misclassification rate of the algorithm. We show how the privacy parameters affect the error bound. In particular, we demonstrate the effect of heterogeneity among population networks and the asymmetry roles of the privacy parameters $q$ and $q^{\prime}$ on the estimation accuracy through rigorous theoretical justifications. The numerical experiments well support our theoretical findings.
The proposed methodology can be extended to a variety of scenarios. First, the ppDSC algorithm involves multiple steps. The methods used in some steps are quite general and can be applied in other contexts. For example, the proposed bias-adjustment procedure can be used in other stochastic block models in different distributed SVD frameworks, e.g., Fan et al. (2019); Arroyo et al. (2021). Second, it is well-known that privacy could be amplified by sampling (Balle et al., 2018). With the same level of noise from privacy preservation, a DP algorithm with sampled data leaks less privacy than that with all data. Therefore, it is desirable to study the trade-off between privacy gain and information loss when sampling the network entries in network analysis. Third, for statistical convenience, we consider the local-DP (Duchi et al., 2018) protection of networks. We may extend the proposed methods in the context of central DP. Fourth, we focus on the distributed learning framework with only one round of communication. It is desirable to extend the results to the distributed framework with multiple rounds, say the communication-efficient distributed power method (Li et al., 2021). Last but not least, the method and theory for estimating $K$ can be an interesting topic to explore in the future. (Fishkind et al., 2013; Li et al., 2020; Ma et al., 2021).
Appendix
Appendix A provides the details of the bias-adjustment procedure. Appendix B contains the additional theoretical results and discussions. Appendix C contains the additional simulation experiment. Appendix D contains the technical proofs for Theorems 1-4 and Corollary 1. Several auxiliary lemmas are provided in Appendix E.
Appendix A Details of bias-adjustment
We illustrate how the diagonal elements of $\overline{A}_{l}^{2}$ would cause bias for $P_{l}^{2}$, even though $\overline{A}_{l}$ is the unbiased estimator of $P_{l}$. Denote the non-diagonal part of $P_{l}$ by $P_{l1}:=P_{l}-{{\operatorname{diag}}}(P_{l})$, and denote the deviation of $\overline{A}_{l}$ from $P_{l1}$ by $X_{l}:=\overline{A}_{l}-P_{l1}$. After simple calculation, we obtain
$\overline{A}_{l}^{2}$ $=(X_{l}+P_{l1})^{2}=X_{l}^{2}+P_{l1}^{2}+X_{l}P_{l1}+P_{l1}X_{l}.$ (A.1)
Plugging in $P_{l1}^{2}=(P_{l}-{\operatorname{diag}}(P_{l}))^{2}$, we have
$\overline{A}_{l}^{2}-P_{l}^{2}$ $=E_{l,0}+E_{l,1}+E_{l,2}+E_{l,3}$ (A.2)
with
$E_{l,0}$ $:={-P_{l}\operatorname{diag}(P_{l})-\operatorname{diag}(P_{l})\cdot P_{l}+(\operatorname{diag}(P_{l}))^{2}},$
$E_{l,1}$ $:=X_{l}(P_{l}-{\operatorname{diag}}(P_{l}))+(P_{l}-\operatorname{diag}(P_{l}))X_{l},$
$E_{l,2}$ $:=X_{l}^{2}-{\operatorname{diag}}(X_{l}^{2}),$
$E_{l,3}$ $:={\operatorname{diag}}(X_{l}^{2}).$
We would see in Appendix D that the first three terms are actually all small. While for the fourth term $E_{l,3}$, we now proceed to bound it. Note that
$(E_{l,3})_{ii}$ $=\sum_{j=1}^{n}X_{l,ij}^{2}=\sum_{j=1,j\neq i}^{n}(\overline{A}_{l,ij}-P_{l,ij})^{2}$
$=\frac{1}{(q+q^{\prime}-1)^{2}}\sum_{j=1}^{n}\mathbb{I}({\tilde{A}}_{l,ij}=1)(1-(1-q^{\prime})-(q+q^{\prime}-1)P_{l,ij})^{2}$
$\quad\quad\quad\quad\quad\quad+\frac{1}{(q+q^{\prime}-1)^{2}}\sum_{j=1}^{n}\mathbb{I}({\tilde{A}}_{l,ij}=0)(1-q^{\prime}+(q+q^{\prime}-1)P_{l,ij})^{2}$
$\leq\frac{\sum_{j=1}^{n}\mathbb{I}({\tilde{A}}_{l,ij}=1)(q^{\prime})^{2}}{(q+q^{\prime}-1)^{2}}+\frac{n(1-q^{\prime}+(q+q^{\prime}-1)\kappa)^{2}}{(q+q^{\prime}-1)^{2}}$
$=\frac{\sum_{j}{\tilde{A}}_{l,ij}(q^{\prime})^{2}}{(q+q^{\prime}-1)^{2}}+\frac{n(1-q^{\prime}+(q+q^{\prime}-1)\kappa)^{2}}{(q+q^{\prime}-1)^{2}},$ (A.3)
where in the first inequality we assume that $q,q^{\prime}>1/2$ and $\kappa=\max_{i,j,l}P_{l,ij}$. Hence, the error would be smaller if we further subtract $\frac{\sum_{j}{\tilde{A}}_{l,ij}(q^{\prime})^{2}}{(q+q^{\prime}-1)^{2}}$ from the $i$th diagonal elements of $\overline{A}_{l}^{2}$ and assume $\kappa=o(1)$. More importantly, as we aim to approximate $\sum_{l=1}^{L}P_{l}^{2}$, this debiasing step is crucial when $L$ is increasing.
Appendix B Additional theoretical results and discussions
B.1 Bound for $\overline{M}_{l}$
We provide the bound for the single matrix $\overline{M}_{l}$. As we will see, due to the absence of $L$, the condition and the dominating term in the resulting bound compared with the centralized bound is changed. The results are summarized in Theorem B.1.
Theorem B.1
Suppose Assumption 2 holds, and $\epsilon\lesssim{\log}\,n$, the estimator $\overline{M}_{l}$ for $Q_{l}$ satisfies
$\max_{l}\|\overline{M}_{l}-Q_{l}\|_{2}\leq C\max\left\{\frac{(q\rho+1-q^{\prime})^{1/2}n^{1/2}\rho\log^{1/2}(n)}{q+q^{\prime}-1},\,\frac{(q\rho+1-q^{\prime})\log n}{(q+q^{\prime}-1)^{2}}\right\},$ (B.1)
with probability larger than $1-O(Ln^{-\gamma})$, where $\gamma,C>0$ are some fixed constants.
The proof of Theorem B.1 is included in Appendix D.
B.2 The effect of bias-adjustment
We theoretically show the effect of bias-adjustment. Denote the non-debiased estimator by
$\tilde{M}:=\frac{1}{L}\sum_{l}\frac{1}{n}\tilde{A}_{l}^{2}$
and the corresponding population matrix by
$\tilde{Q}:=\frac{1}{L}\sum_{l}\left((q+q^{\prime}-1)P_{l}+(1-q^{\prime})\textbf{1}_{n}\textbf{1}_{n}^{\intercal}\right)^{2}/{n}.$
The following proposition provides the estimation error of $\tilde{M}$.
Proposition 2
Suppose Assumption 2 and condition (13) holds. Define
$\small\mathcal{I}:=\max\Big{\{}\frac{(q\rho+1-q^{\prime})^{1/2}n^{1/2}[\rho+(1-q^{\prime})/(q+q^{\prime}-1)]\log^{1/2}(L+n)}{(q+q^{\prime}-1)\sqrt{L}},$
$\quad\quad\quad\quad\quad\quad\quad\quad\frac{(q\rho+1-q^{\prime})\log(L+n)}{(q+q^{\prime}-1)^{2}\sqrt{L}},\,\frac{((q+q^{\prime}-1)\rho+1-q^{\prime})^{2}}{(q+q^{\prime}-1)^{2}}\Big{\}},$
then the non-debiased estimator $\tilde{M}$ of $\tilde{Q}$ satisfies
$\scriptsize\|\tilde{M}-\tilde{Q}\|_{2}\lesssim(q+q^{\prime}-1)^{2}\mathcal{I}$ (B.2)
with probability larger than $1-O((L+n)^{-\nu})$ for some constant $\nu>0$. Moreover, with the same probability,
$\|\tilde{M}-Q\|_{2}\lesssim(q+q^{\prime}-1)^{2}\mathcal{I}+\|\tilde{Q}-Q\|_{2}.$ (B.3)
The proof is similar to that of Theorem 3 provided in Appendix D; hence we omit it. It can be easily seen that $\mathcal{I}$ is larger than (3) in Theorem 3. Hence, if $q,q^{\prime}>c$ for some $c>1/2$, which implies that the multiplier $(q+q^{\prime}-1)^{2}$ is a constant, then (B.2) and (B.3) are both inferior than (3). To derive the bound for the eigen-vector and the misclassification error bound, we note that we can either use (B.2) or (B.3). What we need is only the minimum non-zero eigen-value of the corresponding population matrix. Specifically, if we use (B.3), then the population matrix is $Q$, which is the same as our setting. While if we use (B.2), then the minimum non-zero eigen-value of $\tilde{Q}$ is required. Here, we consider the two-parameter SBMs that $P_{l}$’s are homogeneous, with the inter-community and cross-community probabilities being $a$ and $b$, respectively. In this regime, the minimum non-zero eigen-value of $(q+q^{\prime}-1)P_{l}+(1-q^{\prime})\textbf{1}_{n}\textbf{1}_{n}^{\intercal}$ is identical to that of $(q+q^{\prime}-1)P_{l}$, which implies that the error for the eigen-vector (counterpart of Corollary 1) and the misclassification error (counterpart of Theorem 4) totally depend on (B.3). As a result, we can also see the accuracy gain in eigen-vector estimation and clustering.
Appendix C Additional numerical analysis and results
We conduct the following simulation to show the effect of heterogeneity.
Network-generation.
In Theorem 2, we show that the eigen-gap (i.e., $\lambda_{{\min}}(Q)$) divided by the heterogeneity of connectivity matrices (i.e., $\mathcal{H}$ in Assumption 4) can affect the estimation accuracy of the proposed algorithm theoretically. The heterogeneity can lead to either accuracy gain or loss, depending on the model structure of the network. To illustrate this phenomenon numerically, we consider the following two models for demonstrating the curses of heterogeneity and the blessings of heterogeneity, respectively.
  • Model I (Curses of heterogeneity). We fix $K=2$, $L=12$, $q,q^{\prime}=0.8$ and let $n$ vary from 20 to 400. The communities are balanced. The networks are generated from multi-layer SBMs and equally belong to the following four different classes,
    $B^{(i)}=0.25\cdot{\operatorname{diag}}(1,1)+\alpha C^{(k)}$
    with
    $C^{(1)}=\begin{pmatrix}0.1&0\\ 0&0.1\end{pmatrix},\quad C^{(2)}=-\begin{pmatrix}0.1&0\\ 0&0.1\end{pmatrix},\quad C^{(3)}=-\begin{pmatrix}0.1&0\\ 0&0\end{pmatrix},\quad C^{(4)}=-\begin{pmatrix}0&0\\ 0&0.1\end{pmatrix}.$
    For this model, one can imagine that larger $\alpha$ indicates larger heterogeneity but a similar level of signal strength (eigen-gap). Therefore, larger $\alpha$ leads to larger scaled heterogeneity (i.e., eigen-gap divided by heterogeneity; see the discussions after Theorem 2 for the details). This is verified in Figure C.1(a).
  • Model II (Blessings of heterogeneity). We fix $K=2$, $L=12$, $q,q^{\prime}=0.8$ and let $n$ vary from 20 to 400. The communities are balanced. Similar to Arroyo et al. (2021), the networks are generated from SBMs and equally belong to the following four different classes,
    $B^{(i)}=0.25(11^{\intercal})+\alpha C^{(k)}$
    with
    $C^{(1)}=\begin{pmatrix}0.1&0\\ 0&0.1\end{pmatrix},\quad C^{(2)}=-\begin{pmatrix}0.1&0\\ 0&0.1\end{pmatrix},\quad C^{(3)}=\begin{pmatrix}0.1&0\\ 0&0\end{pmatrix},\quad C^{(4)}=\begin{pmatrix}0&0\\ 0&0.1\end{pmatrix}.$
    For this model, larger $\alpha$ indicates larger heterogeneity but stronger signal strength (eigen-gap). Thus, the composite effects lead to larger $\alpha$ corresponding to smaller scaled heterogeneity. See Figure C.2(a) for the illustration.
Results.
The averaged results over 20 replications for Model I and II are displayed in Figure C.1 and C.2, respectively. The results are consistent with the theoretical results given in Theorem 2. For Model I, we see that large heterogeneity (i.e., large $\alpha$) leads to inferior performance of ppDSC, especially in terms of the projection distance, which is partially because that large $\alpha$ leads to large scaled heterogeneity (i.e., eigen-gap divided by heterogeneity) in Theorem 2; see Figure C.1. For Model II, the results are on the opposite side in that large heterogeneity (i.e., large $\alpha$) leads to performance gain of ppDSC, which is partially due to the fact that large $\alpha$ leads to large eigen-gap and small scaled heterogeneity; see Figure C.2. In addition, our estimator ppDSC is better than the biased ppDSC-1b and ppDSC-2b, and is closer to the Oracle estimator.
(a) Population parameters
(b) Projection distance
(c) Misclassification rate
Figure C.1: Illustration for the curse of heterogeneity (Model I). (a) corresponds to population parameters including the heterogeneity, eigen-gap and their ratios against $\alpha$. (b) and (c) correspond to the projection distance and misclassification rate of ppDSC, ppDSC-1b, ppDSC-2b and Oracle.
(a) Population parameters
(b) Projection distance
(c) Misclassification rate
Figure C.2: Illustration for the blessing of heterogeneity (Model II). (a) corresponds to population parameters including the heterogeneity, eigen-gap and their ratios against $\alpha$. (b) and (c) correspond to the projection distance and misclassification rate of ppDSC, ppDSC-1b, ppDSC-2b and Oracle.
Appendix D Main proofs
D.1 Proof of Theorem 1
For $i\in\{1,...,L\}$, denote the RR-perturbed $A_{i}$ by ${\tilde{A}}_{i}$. By Proposition 1, for each $i$, ${\tilde{A}}_{i}$ satisfies $\epsilon$-edge-DP. Further, by the parallel composition theorem (McSherry, 2009), $({\tilde{A}}_{1},...,{\tilde{A}}_{L})$ also satisfies $\epsilon$-edge-DP. Finally, by the post-processing property of DP (Proposition 1), Algorithm 1 can achieve $\epsilon$-edge-DP. $\blacksquare$
D.2 Proof of Theorem 2
First, we note that we have a degree of freedom in choosing $V$. For simplicity, we assume that the reference solution $\hat{V}_{o}$ in Algorithm 1 is already aligned with $V$ in that $\arg\min_{Z\in\mathcal{O}_{K}}\|\hat{V}_{o}Z-V\|_{F}=I_{K}$. Otherwise, we could work with $VZ$ for some $Z\in\mathcal{O}_{K}$ instead.
We will make use of the following result to prove, which is modified from Charisopoulos et al. (2021) using the notation of this paper.
Proposition 3
Suppose
$\lambda_{{\min}}(Q)\geq\delta,\quad{for\;some\;}\delta>0$ (D.1)
and
$\|\overline{M}_{l}-Q\|_{2}\leq\frac{\delta}{8},\quad{for\;all\;}l=1,...,L.$ (D.2)
Then the output $\tilde{V}$ of Algorithm 1 satisfies that
${\mathrm{dist}}(\tilde{V},V)\lesssim\frac{1}{\delta^{2}}\max_{l}\|\overline{M}_{l}-Q\|_{2}^{2}+\frac{1}{\delta}\|\frac{1}{L}\sum_{l=1}^{L}\overline{M}_{l}-Q\|_{2}.$
We first have that
$\|\overline{M}_{l}-Q\|_{2}$ $\leq\|\overline{M}_{l}-Q_{l}\|_{2}+\|Q_{l}-Q\|_{2}$
$\leq\|\overline{M}_{l}-Q_{l}\|_{2}+C^{\prime}\mathcal{H}(n,\rho,L),$
where the last inequality follows from
$\|Q_{l}-Q\|_{2}$ $=\left\|\frac{P_{l}^{2}}{n}-\frac{1}{L}\sum_{i=1}^{L}\frac{P_{i}^{2}}{n}\right\|_{2}\leq\frac{1}{n}\|\Delta\|_{2}^{2}\|B_{l}\Delta^{2}B_{l}-\frac{1}{L}\sum_{i=1}^{L}B_{i}\Delta^{2}B_{i}\|_{2}$
$\leq C\frac{1}{K}\|B_{l}\Delta^{2}B_{l}-\frac{1}{L}\sum_{i=1}^{L}B_{i}\Delta^{2}B_{i}\|_{2}\leq C^{\prime}\mathcal{H}(n,\rho,L),$
by using the fact that
$P_{l}^{2}=\Theta B_{l}\Theta^{\intercal}\Theta B_{l}\Theta^{\intercal}=\bar{\Theta}\Delta(B_{l}\Delta^{2}B_{l})\Delta\bar{\Theta}^{\intercal}$
with $\Delta:={\operatorname{diag}}(\sqrt{n_{1}},...,\sqrt{n_{K}})$, $\bar{\Theta}:=\Theta\Delta^{-1}$, and $\bar{\Theta}^{\intercal}\bar{\Theta}=\mathbb{I}$, and Assumptions 1 and 4. To meet (D.1) and (D.2), we let $\delta\asymp\lambda_{\min}(Q)$ and enforce
$\|\overline{M}_{l}-Q_{l}\|_{2}\leq\frac{\lambda_{\min}(Q)}{16}\quad{and}\quad\mathcal{H}\leq\frac{\lambda_{\min}(Q)}{16C^{\prime}},$
which follows by our condition in Theorem 2. We can thus use the results in Proposition 3 to obtain that,
$\mathrm{dist}(\tilde{V},V)\lesssim{\max_{l}\frac{\|\overline{M}_{l}-Q_{l}\|_{2}^{2}}{\lambda_{\min}^{2}(Q)}}+{\frac{\mathcal{H}^{2}(n,\rho,L)}{\lambda_{\min}^{2}(Q)}}+\frac{\|\sum_{l}\overline{M}_{l}/L-Q\|_{2}}{\lambda_{\min}(Q)}.$
The results of Theorem 2 are proved. $\blacksquare$
D.3 Proof of Theorem 3
We make use of the concentration tools developed in Lei and Lin (2022) to prove. Recall that $\overline{M}=\frac{1}{L}\sum_{l=1}^{L}\overline{M}_{l}$ and (11), then we can decompose the difference between $\overline{M}$ and $Q$ as
$n\overline{M}-nQ=\sum_{l=1}^{L}\frac{1}{L}(n\overline{M}_{l}-nQ_{l}):=E_{0}+E_{1}+E_{2}+E_{3}-\sum_{l}\frac{1}{L}\frac{G_{l}(q^{\prime})^{2}}{(q+q^{\prime}-1)^{2}},$
with $E_{i}:=\frac{1}{L}\sum_{l=1}^{L}E_{l,i}$ for $i=0,1,2,3$ and $E_{l,i}$ is the same with those in (A.2). Now, we bound these error terms, respectively.
For $E_{0}:=\sum_{l}\frac{1}{L}\left({-P_{l}}\cdot{\operatorname{diag}}(P_{l})-{\operatorname{diag}}(P_{l})\cdot P_{l}+({\operatorname{diag}}(P_{l}))^{2}\right)$, we have
$\|E_{0}\|_{2}\leq 2\max_{l}\|{\operatorname{diag}}(P_{l})\|_{2}\max_{l}\|P_{l}\|_{{F}}+\max_{l}\|{\operatorname{diag}}(P_{l})\|_{2}^{2}\leq Cn\rho^{2}.$
For $E_{1}:=\sum_{l}\frac{1}{L}\left(X_{l}(P_{l}-\operatorname{diag}(P_{l}))+(P_{l}-\operatorname{diag}(P_{l}))X_{l}\right)$, we would use the result in Lemma 1. To that end, we first show that each entry of $(q+q^{\prime}-1)X_{l}$ satisfies the Bernstein tail condition (see Definition 2). Given $i\neq j$, with slight abuse of notation, define
$Y_{l}:=(q+q^{\prime}-1)X_{l,ij}=(q+q^{\prime}-1)(\overline{A}_{l,ij}-P_{l,ij})={\tilde{A}}_{l,ij}-(1-q^{\prime})-(q+q^{\prime}-1)P_{l,ij}.$
Then
$Y_{l}=\begin{cases}q^{\prime}-(q+q^{\prime}-1)P_{l,ij}\;\,\mathrm{with\;probability}\,\;qP_{l,ij}+(1-P_{l,ij})(1-q^{\prime}),\\ q^{\prime}-1-(q+q^{\prime}-1)P_{l,ij}\;\,\mathrm{with\;probability}\,\;q^{\prime}(1-P_{l,ij})+P_{l,ij}(1-q),\end{cases}$
and for $k\geq 2$, we have
${\mathbb{E}}|Y_{l}^{k}|$ $\leq|q^{\prime}-(q+q^{\prime}-1)P_{l,ij}|^{k}\Big{[}qP_{l,ij}+(1-P_{l,ij})(1-q^{\prime})\Big{]}$
$\quad\quad\quad+|(q+q^{\prime}-1)P_{l,ij}+1-q^{\prime}|^{k}\Big{[}q^{\prime}(1-P_{l,ij})+P_{l,ij}(1-q)\Big{]}$
$\leq\left(qP_{l,ij}+(1-P_{l,ij})(1-q^{\prime})\right)+\left((q+q^{\prime}-1)P_{l,ij}+1-q^{\prime}\right)$
$\leq 2\left(q\rho+1-q^{\prime}\right)\leq\frac{v_{1}}{2}k!R_{1}^{k-2},$ (D.3)
where $v_{1}:=4\left(q\rho+1-q^{\prime}\right)$ and $R_{1}:=1$, and we used the fact that $P_{l,ij}=o(1)$. Hence, $(q+q^{\prime}-1)X_{l,ij}$ is $(v_{1},R_{1})$-Bernstein. In addition, it holds that $\|\sum_{l}(P_{l}-{\operatorname{diag}}(P_{l}))^{2}\|_{2}\leq\sum_{l}\|P_{l}-{\operatorname{diag}}(P_{l})\|_{{F}}^{2}\leq Ln^{2}\rho^{2}$ and $\max_{l}\|P_{l}-{\operatorname{diag}}(P_{l})\|_{2,\infty}\leq\sqrt{n}\rho$. Applying Lemma 1, we obtain
${\mathbb{P}}((q+q^{\prime}-1)L\|E_{1}\|_{2}\geq t)\leq 4n\cdot\exp\left\{\frac{-t^{2}/2}{4(q\rho+(1-q^{\prime}))Ln^{3}\rho^{2}+\sqrt{n}\rho t}\right\}.$
Choosing $t=c(q\rho+1-q^{\prime})^{1/2}n^{3/2}L^{1/2}\rho\log^{1/2}(L+n)$, it turns out that
$\sqrt{n}\rho t\leq c4(q\rho+(1-q^{\prime}))Ln^{3}\rho^{2},$
which is $\log^{1/2}(L+n)\leq c(q\rho+1-q^{\prime})^{1/2}nL^{1/2}$ and can be certainly met if $\log^{1/2}(L+n)\leq c(1-q^{\prime})^{1/2}nL^{1/2}$, which can be satisfied by the fact that $q^{\prime}={\mathrm{e}}^{\epsilon}/(1+{\mathrm{e}}^{\epsilon})$ and the condition (13). As a result, we have with probability larger than $1-(L+n)^{-\beta}$ for some $\beta>0$ that
$\|E_{1}\|_{2}\leq c\frac{(q\rho+1-q^{\prime})^{1/2}}{q+q^{\prime}-1}\cdot n^{3/2}\rho\log^{1/2}(L+n)/\sqrt{L}.$
For $E_{2}:=\sum_{l}\frac{1}{L}\left(X_{l}^{2}-{\operatorname{diag}}(X^{2}_{l})\right)$, we would use the results in Lemma 2. For this purpose, we first show that given an independent copy $(q+q^{\prime}-1)\tilde{X}_{l}$ of $(q+q^{\prime}-1)X_{l}$, each $(q+q^{\prime}-1)^{2}X_{l,ij}\tilde{X}_{l,ij}$ satisfies the Bernstein tail condition. Actually, similar to the derivation of (D.3), we can easily show that $(q+q^{\prime}-1)^{2}X_{l,ij}\tilde{X}_{l,ij}$ is $(v^{\prime}_{2},R^{\prime}_{2})$-Bernstein with $v^{\prime}_{2}=8\left(q\rho+1-q^{\prime}\right)^{2}$ and $R^{\prime}_{2}=1$. Then by Lemma 2, we obtain with probability larger than $1-O((L+n)^{-1})$ that
$\|E_{2}\|_{2}\leq c\max\left\{\frac{(q\rho+1-q^{\prime})n\log(L+n),(q\rho+1-q^{\prime})^{1/2}\sqrt{n}\log^{3/2}(L+n),\log^{2}(L+n)/\sqrt{L}}{\sqrt{L}(q+q^{\prime}-1)^{2}}\right\}.$ (D.4)
Further, by the condition (13), we have $\log(L+n)\leq cn(1-q^{\prime})\leq cn(q\rho+1-q^{\prime})$ and (D.4) reduces to
$\|E_{2}\|_{2}\leq c\frac{(q\rho+1-q^{\prime})n\log(L+n)}{\sqrt{L}(q+q^{\prime}-1)^{2}}.$
For $E_{3}-\sum_{l}\frac{1}{L}\frac{G_{l}(q^{\prime})^{2}}{(q+q^{\prime}-1)^{2}}$, recalling (A), it is easy to see that
$\|E_{3}-\sum_{l}\frac{1}{L}\frac{G_{l}(q^{\prime})^{2}}{(q+q^{\prime}-1)^{2}}\|_{2}\leq C\frac{n((q+q^{\prime}-1)\rho+1-q^{\prime})^{2}}{(q+q^{\prime}-1)^{2}}.$
Consequently, $E_{1}$, $E_{2}$ and $E_{3}$ are the main effects and we obtained the results of Theorem 3. $\blacksquare$
D.4 Proof of Theorem B.1
The proof is similar to the that of Theorem 3. We here go through the main steps.
Recalling (A.2) and (11), we can write
$n\overline{M}_{l}-nQ_{l}:=n\overline{M}_{l}-nP_{l}^{2}:=E_{l,0}+E_{l,1}+E_{l,2}+E_{l,3}-\frac{G_{l}(q^{\prime})^{2}}{(q+q^{\prime}-1)^{2}}.$
Next, we bound these error terms, respectively.
For $E_{l,0}:={-P_{l}}\cdot{\operatorname{diag}}(P_{l})-{\operatorname{diag}}(P_{l})\cdot P_{l}+({\operatorname{diag}}(P_{l}))^{2}$, we can easily show that
$\|E_{l,0}\|_{2}\leq cn\rho^{2}.$
For $E_{l,1}:=X_{l}(P_{l}-{\operatorname{diag}}(P_{l}))+(P_{l}-{\operatorname{diag}}(P_{l}))X_{l}$, using Lemma 1 with the same Bernstein tail condition as those in the proof of Theorem 3, we can obtain that with probability larger than $1-n^{-\beta}$ for some $\beta>0$,
$\|E_{l,1}\|_{2}\leq c\frac{(q\rho+1-q^{\prime})^{1/2}}{q+q^{\prime}-1}n^{3/2}\rho\log^{1/2}(n),$
provided that $\log^{1/2}\,n\leq C(q\rho+1-q^{\prime})^{1/2}n$, which can be implied by $\log n\leq Cn(1-q^{\prime})=C{n}/({1+{\mathrm{e}}^{\epsilon}})$ by noting our condition $\epsilon\leq C{\log}\,n$.
For $E_{l,2}:=X_{l}^{2}-{\operatorname{diag}}(X^{2}_{l})$, using Lemma 2 with the same Bernstein tail condition as the one used in the proof of Theorem 3, we observe that with probability larger than $1-O(n^{-1})$,
$\|E_{l,2}\|_{2}\leq c\max\left\{vn\log(n),\,\sqrt{v}\sqrt{n}\log^{3/2}(n),\,\log^{2}(n)\right\}/(q+q^{\prime}-1)^{2},$ (D.5)
where $v:=q\rho+1-q^{\prime}$. Note that by the condition that $\epsilon\leq C{\log}\,n$, we have ${\log}\,n\leq Cn(1-q^{\prime})\leq Cn(q\rho+1-q^{\prime})$, and thus the first term in (D.5) dominates and
$\|E_{l,2}\|_{2}\leq c\frac{(q\rho+1-q^{\prime})n\log(n)}{(q+q^{\prime}-1)^{2}}.$
For the last two error terms $E_{l,3}-\frac{G_{l}(q^{\prime})^{2}}{(q+q^{\prime}-1)^{2}}$, it is easy to see that
$\|E_{l,3}-\frac{G_{l}(q^{\prime})^{2}}{(q+q^{\prime}-1)^{2}}\|_{2}\leq c\frac{(1-q^{\prime}+(q+q^{\prime}-1)\rho)^{2}}{(q+q^{\prime}-1)^{2}}\cdot n.$
It is easy to see that $E_{l,0}\leq CE_{l,1}$. Now we show that $E_{3}\leq CE_{2}$, for which we only need to show
$(1-q^{\prime}+(q+q^{\prime}-1)\rho)^{2}\leq C(q\rho+1-q^{\prime}){\log}\,n.$
Actually, we have $(1-q^{\prime})^{2}\leq C\log\,n(1-q^{\prime})\leq C\log\,n(q\rho+1-q^{\prime})$, and $((q+q^{\prime}-1)\rho)^{2}\leq Cq\rho\log\,n\leq C(q\rho+1-q^{\prime}){\log}\,n$ by noting that $q>1/2$ and $q+q^{\prime}-1\leq 1$. As a result, $E_{l,1}$ and $E_{l,2}$ dominate. After using the union bound, we finally have
$\max_{l}\|\overline{M}_{l}-Q_{l}\|_{2}\leq C\max\left\{\frac{(q\rho+1-q^{\prime})^{1/2}n^{1/2}\rho\log^{1/2}(n)}{q+q^{\prime}-1},\,\frac{(q\rho+1-q^{\prime})\log n}{(q+q^{\prime}-1)^{2}}\right\},$
with probability larger than $1-O(Ln^{-\gamma})$ for some constant $\gamma>0$. The results of Theorem B.1 are arrived. $\blacksquare$
D.5 Proof of Corollary 1
The proof of Corollary 1 follows directly from that of Theorem 2. First, we need to show that $Q$ is of rank $K$ and provide the lower bound of $\lambda_{K}(Q)$. Indeed, define $\Delta={\operatorname{diag}}(\sqrt{n_{1}},...,\sqrt{n_{K}})$ and denote $\bar{\Theta}:=\Theta\Delta^{-1}$, then $\bar{\Theta}$ is orthogonal and we can write $Q$ as
$Q$ $=\frac{\rho^{2}}{nL}\bar{\Theta}\Delta(\sum_{l=1}^{L}B_{l,0}\Delta^{2}B_{l,0})\Delta\bar{\Theta}^{\intercal}$
$\succeq c\frac{\rho^{2}}{L}\bar{\Theta}\Delta(\sum_{l=1}^{L}B_{l,0}B_{l,0})\Delta\bar{\Theta}^{\intercal}$
$\succeq c{\rho^{2}}\bar{\Theta}\Delta^{2}\bar{\Theta}^{\intercal}$
$\succeq c{n\rho^{2}}\bar{\Theta}\bar{\Theta}^{\intercal},$ (D.6)
where we used Assumption 2 in the first equality, Assumption 1 in the first and last inequality, and Assumption 3 in the second inequality. (D.5) implies that $Q$’s rank is $K$ and $\lambda_{K}(Q)\geq cn\rho^{2}$.
Next, we specify the $\mathcal{I}_{1},\mathcal{I}_{2}$ and $\mathcal{I}_{3}$ and the corresponding condition in Theorem 2 that $\mathcal{I}_{1}$ and $\mathcal{I}_{2}$ need to be smaller enough. Recall the bound for the single matrix $\overline{M}_{l}$ in Theorem B.1, we find that when
$\rho\gtrsim\frac{(q\rho+1-q^{\prime})^{1/2}}{q+q^{\prime}-1}\cdot\sqrt{\frac{\log(n)}{n}},$ (D.7)
the bound in (B.1) reduces to the first term $\frac{(q\rho+1-q^{\prime})^{1/2}n^{1/2}\rho\log^{1/2}(n)}{q+q^{\prime}-1}$ which is automatically smaller than $\lambda_{K}(Q)$ provided with (D.7). As a result, $\mathcal{I}_{1}$ is specified under (D.7). For $\mathcal{I}_{2}$, we do not specify it but only require the condition that
$\mathcal{H}\lesssim n\rho^{2}.$
For $\mathcal{I}_{3}$, the bound directly follows from Theorem 3.
As a result, the results of Corollary 1 have been arrived. $\blacksquare$
D.6 Proof of Theorem 4
The general proving strategy is same with Lei and Rinaldo (2015); Rohe et al. (2011); Zhang et al. (2022), among others. To fix ideas, we now recall and introduce some notation. $V$ denotes $K$ leading eigenvectors of $Q$ and $\tilde{V}$ denotes the $K$ leading eigenvectors produced by Algorithm 1. Likewise, $\tilde{V}^{\prime}={{\tilde{\Theta}}}{{X}}$ denotes the estimated eigenvectors by the $k$-means with $\tilde{V}$ being its input, where ${\tilde{\Theta}}$ is the estimated membership matrix and ${{X}}$ denotes the centriods.
First, we have that there exists $O\in\mathcal{O}_{K}$ such that
$\|{\tilde{V}^{\prime}}-VO\|_{\tiny F}^{2}$ $=\|{\tilde{V}^{\prime}}-\tilde{V}+\tilde{V}-VO\|_{\tiny F}^{2}$
$\leq 2\|VO-\tilde{V}\|_{\tiny F}^{2}+2\|\tilde{V}-VO\|_{\tiny F}^{2}$
$=4\|\tilde{V}-VO\|_{\tiny F}^{2}$
$\leq C\|\tilde{V}-VO\|_{2}^{2}\leq C\mathcal{S}^{2},$
where $\mathcal{S}:=\mathcal{I}_{1}+\mathcal{I}_{2}+\mathcal{I}_{3}$ with $\mathcal{I}_{i}$’s defined in Corollary 1, the first inequality follows because we assume that ${\tilde{V}^{\prime}}$ is the global solution minimum of the following $k$-means objective and $VO$ is a feasible solution by Lemma 4,
$(\tilde{\Theta},X)=\underset{{{\Theta}\in\mathbb{M}_{n,K},{X}\in\mathbb{R}^{K\times K}}}{{\mathrm{arg\;min}}}\;\|{\Theta}{X}-\tilde{V}\|_{\tiny F}^{2},$
and the last inequality follows from Corollary 1 and the fact that $\min_{O\in\mathcal{O}_{K}}\|\tilde{V}-VO\|_{2}^{2}\lesssim\mathrm{dist}(\tilde{V},V)$.
Then, we begin to bound the fraction of misclassified nodes. Define
=$\mathcal{M}=\{i\in\{1,...,n\}:\;\|{\tilde{V}^{\prime}}_{i\ast}-({V}O)_{i\ast}\|_{\tiny\mathrm{2}}\geq\frac{\tau}{2}\},$
with
$\tau=\min_{k\neq l}\sqrt{(n_{k})^{-1}+(n_{l})^{-1}}.$
Then we can see obviously that
$\frac{|\mathcal{M}|}{n}\leq C\mathcal{S}^{2}/(n\tau^{2})\leq C^{\prime}\mathcal{S}^{2}.$
where the last inequality follows from Assumption 1.
Finally, we show that the nodes outside $\mathcal{M}$ are correctly clustered to their underlying communities. Define $T_{k}:=G_{k}\backslash\mathcal{M}$, where $G_{k}$ denotes the set of nodes within the true community $k$. $T_{k}$ is not an empty set provided that $|\mathcal{M}|<n_{k}$ for any $k$ by our condition. Let $T=\cup_{k=1}^{K}T_{k}$. We will rightly see that the rows in $(VO)_{T\ast}$ has a one to one correspondence with those in ${\tilde{V}^{\prime}}_{T\ast}$. On the one hand, for $i\in T_{k}$ and $j\in T_{l}$ with $l\neq k$, ${\tilde{V}^{\prime}}_{i\ast}\neq{\tilde{V}^{\prime}}_{j\ast}$, otherwise the following contradiction follows
$\tau\leq\|(VO)_{i\ast}-(VO)_{j\ast}\|_{2}\leq\|(VO)_{i\ast}-{\tilde{V}^{\prime}}_{i\ast}\|_{2}+\|(VO)_{j\ast}-{\tilde{V}^{\prime}}_{j\ast}\|_{2}<\frac{\tau}{2}+\frac{\tau}{2},$
where the last inequality is implied by the definition of $\mathcal{M}$ and the first inequality is implied by Lemma 4 by noting
$Q=\frac{1}{L}\sum_{l}P_{l}^{2}=\frac{1}{L}\sum_{l}\Theta B_{l}\Theta^{\intercal}\Theta B_{l}\Theta^{\intercal}=\bar{\Theta}\Delta\frac{1}{L}\sum_{l}(B_{l}\Delta^{2}B_{l})\Delta\bar{\Theta}^{\intercal}:={\Theta}\overline{B}{\Theta}^{\intercal},$
where $\Delta={\operatorname{diag}}(\sqrt{n_{1}},...,\sqrt{n_{K}})$, $\bar{\Theta}:=\Theta\Delta^{-1}$, $\bar{\Theta}^{\intercal}\bar{\Theta}=\mathbb{I}$ and $\bar{B}$ is of full rank by our condition (see (D.5) for explanations). On the other hand, for $i,j\in T_{k}$, ${\tilde{V}^{\prime}}_{i\ast}={\tilde{V}^{\prime}}_{j\ast}$, because otherwise ${\tilde{V}^{\prime}}_{T\ast}$ has more than $K$ distinct rows which is contradict with the fact that the community size is $K$.
So far, we have arrived at the conclusion of Theorem 4. $\blacksquare$
Appendix E Auxiliary lemmas
We provide auxiliary lemmas that are used for proving the theorems in the paper.
Definition 2 (Bernstein tail condition)
A random variable $Y$ is said to be $(v,R)$-Bernstein if ${\mathbb{E}}[|Y|^{k}]\leq\frac{v}{2}k!R^{k-2}$ for all integers $k\geq 2$.
Lemma 1 (Theorem 3 in Lei and Lin (2022))
Let $X_{1},...,X_{L}$ be a sequence of independent $n\times r$ random matrices with zero-mean independent entries being $(v_{1},R_{1})$-Bernstein. Let $H_{1},...H_{L}$ be any sequence of $r\times n$ non-random matrices. Then for all $t>0$,
${\mathbb{P}}(\|\sum_{l=1}^{L}X_{l}H_{l}\|_{2}\geq t)\leq 4n\exp\left(-\frac{t^{2}/2}{v_{1}n\|\sum_{l}H_{l}^{\intercal}H_{l}\|_{2}+R_{1}\max_{l}\|H_{l}\|_{2,\infty}t}\right).$
$\blacksquare$
Lemma 2 (Theorem 4 in Lei and Lin (2022))
Let $X_{1},X_{2},...,X_{L}$ be independent $n\times n$ random symmetric matrix with independent diagonal and upper diagonal entries. For any $l\in[L]$ and any pair of $(i,j)(i\leq j)$, $X_{l,ij}$ is $(v_{1},R_{1})$-Bernstein, and $X_{l,ij}\tilde{X}_{l,ij}$ is $(v^{\prime}_{2},R^{\prime}_{2})$-Bernstein where $\tilde{X}_{l,ij}$ is an independent copy of $X_{l,ij}$. Then with probability larger than $1-O((n+L)^{-1})$,
$\|\sum_{l}X_{l}^{2}-{\operatorname{diag}}(\sum_{l}X_{l}^{2})\|_{2}\leq C$ $\Big{[}v_{1}n\log(L+n)\sqrt{L}+\sqrt{v_{1}}R_{1}\sqrt{Ln}\log^{3/2}(L+n)$
$\quad\quad+\sqrt{v^{\prime}_{2}}\sqrt{Ln}\log(L+n)+(R_{1}^{2}+R^{\prime}_{2})\log^{2}(L+n)\Big{]}.$
Moreover, if the squared entry $X_{l,ij}^{2}$ is $(v_{2},R_{2})$-Bernstein, then with probability larger than $1-O((n+L)^{-1})$.
$\|{\operatorname{diag}}(\sum_{l}X_{l}^{2})-\mathbb{E}({\operatorname{diag}}(\sum_{l}X_{l}^{2}))\|_{2}\leq C\left[\sqrt{v_{2}}\log(L+n)\sqrt{Ln}+R_{2}\log(L+n)\right].$
$\blacksquare$
Lemma 3 (Davis-Kahan theorem)
Let $Q\in\mathbb{R}^{n\times n}$ be a rank $K$ symmetric matrix with smallest nonzero singular value $\gamma_{n}$. Let $M$ be any symmetric matrix and $\hat{U},U\in\mathbb{R}^{n\times K}$ be the $K$ leading eigenvectors of $M$ and $Q$, respectively. Then there exists a $K\times K$ orthogonal matrix $O$ such that
$\|\hat{U}-UO\|_{F}\leq\frac{2\sqrt{2K}\|M-Q\|_{2}}{\gamma_{n}}.$
Proof. See Lei and Rinaldo (2015); Chen et al. (2021b), among others. $\blacksquare$
Lemma 4
Define the population matrix $Q=\Theta\bar{B}\Theta^{\intercal}$. Suppose $\bar{B}$ is of full rank and denote the eigen-decomposition of $Q$ by ${V}_{n\times K}{\Sigma}_{K\times K}{V}^{\intercal}_{K\times n}$. Then, $V=\Theta X$ for some given matrix $X$. Specifically, for $\Theta_{i\ast}=\Theta_{j\ast}$, we have ${V}_{i\ast}={V}_{j\ast}$; while for $\Theta_{i\ast}\neq\Theta_{j\ast}$, we have $\|{V}_{i\ast}-{V}_{j\ast}\|_{2}=\sqrt{(n_{g_{i}})^{-1}+(n_{g_{j}})^{-1}}$.
Proof. See Lei and Rinaldo (2015); Guo et al. (2020); Zhang et al. (2022), among others. $\blacksquare$

### v2

@@
-
+ppDSC_JCGS_SM
 Abstract
-Modern multi-layer networks are commonly stored and analyzed in a local and distributed fashion because of the privacy, ownership, and communication costs.
-The literature on the model-based statistical methods for community detection based on these data is still limited.
-This paper proposes a new method for consensus community detection and estimation in a multi-layer stochastic block model using locally stored and computed network data with privacy protection.
-A novel algorithm named privacy-preserving Distributed Spectral Clustering (ppDSC) is developed.
-To preserve the edges’ privacy, we adopt the randomized response (RR) mechanism to perturb the network edges, which satisfies the strong notion of differential privacy.
-The ppDSC algorithm is performed on the squared RR-perturbed adjacency matrices to prevent possible cancellation of communities among different layers.
-To remove the bias incurred by RR and the squared network matrices, we develop a two-step bias-adjustment procedure.
-Then we perform eigen-decomposition on the debiased matrices, aggregation of the local eigenvectors using an orthogonal Procrustes transformation, and $k$-means clustering.
-We provide theoretical analysis on the statistical errors of ppDSC in terms of eigen-vector estimation.
-In addition, the blessings and curses of network heterogeneity are well-explained by our bounds.
-Keywords: Multi-layer stochastic block model, Distributed SVD, Bias-correction, Differential privacy
+This paper proposes a new efficient and scalable consensus community detection approach for multi-layer stochastic block models in a federated learning framework, where network layers are locally stored across several machines and have privacy-preserving needs.
+Specifically, we develop a spectral clustering-based algorithm named privacy-preserving Distributed Spectral Clustering (ppDSC).
+To reduce the bias incurred by the randomized response (RR) mechanism for achieving differential privacy, we develop a two-step bias adjustment procedure.
+For federated learning, we perform the eigen-decomposition of each privacy-preserved network locally and then aggregate the local eigenvectors using orthogonal Procrustes transformation on the central server.
+We establish a novel bound on the misclassification rate of ppDSC.
+The new bound reveals the asymmetric roles of the two edge-flipping probabilities of the RR in the misclassification rate.
+Through the bound, we can also find the optimal choices for the flipping probabilities given a fixed privacy budget.
+Moreover, we show that ppDSC enjoys the same statistical error rate as its centralized counterpart when the number of machines satisfies a polynomial order with the sample size on each local machine and the effective heterogeneity is well controlled.
+Keywords: Bias-correction, Differential privacy, Distributed SVD, Multi-layer stochastic block models
 1 Introduction
 Advances in modern technology have facilitated the collection of multiple network data from various sources (e.g., Facebook, LinkedIn, and Twitter) that present disparate social ties (Kairouz et al., 2021).
-This data type can be well represented by multi-layer networks (Mucha et al., 2010; Boccaletti et al., 2014; Kivelä et al., 2014; Han et al., 2015), in which the nodes are the entities of interest, and the network layers reflect their relationships across different sources.
-As a result, the study of multi-layer networks has received remarkable attention recently.
-Most existing works assume that all network layers are stored in a single central machine, see, e.g., Paul and Chen (2016); Bhattacharyya and Chatterjee (2018); Liu et al.
+This type of data can be well represented by multi-layer networks (Mucha et al., 2010), where the nodes are the entities of interest, and the network layers reflect their relationships across different local sources.
+In practice, a multi-layer network is often stored in a distributed manner because of ownership and communication costs, among others (Kairouz et al., 2021).
+Moreover, the network edges from local sources often contain sensitive and private information (Dwork et al., 2014).
+Therefore, the analysis of the multi-layer networks should be naturally based on a distributed and federated architecture with additional privacy-preserving guarantees.
+In the literature, the multi-layer stochastic block model (multi-layer SBM) (Han et al., 2015; Paul and Chen, 2016; Barbillon et al., 2017) has been popularly used for the community detection task in multi-layer networks, in which each layer is assumed to share the same unobserved community membership.
+Among the various methods for community detection under multi-layer SBMs, spectral clustering-based methods stand out because of their computational tractability.
+There are two problems in developing the spectral clustering-based method for multi-layer SBMs in a privacy-preserving and distributed learning paradigm.
+The first problem is how to obtain theoretically sound statistics for each layer of networks in a privacy-preserving way.
+The randomized response (RR) mechanism (Warner, 1965; Karwa et al., 2017) that simply flips all edges in the network with a certain probability has been developed to achieve the rigorous differential privacy (DP) notion (Dwork et al., 2006).
+However, it undoubtedly brings bias into the multi-layer SBMs.
+Thus, a bias-correction procedure is crucially needed for community detection in the composite and RR-perturbed multi-layer SBMs.
+The second problem is how to develop a communication-efficient distributed learning algorithm for detecting communities under multi-layer SBMs.
+Recent years have seen various spectral clustering-based methods developed for community detection in multi-layer networks, say, the methods based on the aggregated adjacency matrix (e.g., Paul and Chen, 2020; Lei and Lin, 2022) and the tensor-based methods (e.g., Jing et al., 2021).
+However, these methods are developed in the centralized paradigm, where all network layers are stored on a single central machine.
+In the federated learning paradigm, directly sending the RR-perturbed networks stored on the local machines to the central server can bring enormous communication costs or storage costs to the server, especially for large networks.
+Therefore, performing the eigen-decomposition step of spectral clustering on each local machine is a natural way to reduce the communication cost.
+A follow-up problem is how to combine the eigenvectors to cope with subspaces’ orthogonal ambiguity.
+In this paper, we propose a privacy-preserving Distributed Spectral Clustering (ppDSC) algorithm with only one-round of communication for detecting the communities of locally stored multi-layer network.
+In particular, we assume that the layers of an $L$-layer network are distributed across $m\ (m\leq L)$ machines.
+In our algorithm, each machine first performs a carefully designed two-step bias-correction procedure on the squared RR-perturbed local networks, where we adopt the squared matrix to avoid possible cancellation of communities among different layers (Lei and Lin, 2022).
+Then, the eigen-decomposition is performed locally.
+Finally, the central server collects the local eigenspaces and conducts the spectral clustering using the average of orthogonal Procrustes transformed (OPT) (Schönemann, 1966; Cape, 2020; Charisopoulos et al., 2021) eigenspaces.
+The main contributions of this paper are summarized below.
+First, we propose a one-shot-based distributed spectral clustering algorithm for detecting the communities of RR-perturbed multi-layer SBMs.
+In particular, we develop a debiased estimator to accommodate the noise incurred by the RR mechanism and the square transformation.
+The debiased estimator can consistently estimate the original population matrix, which leads to smaller eigenvector and misclassification error bounds than the non-debiased estimator (with respect to its population matrix or the original population matrix).
+To the best of our knowledge, the theoretical examination of the bias-correction is rarely addressed in existing literature.
+We propose to aggregate the local eigenspaces instead of the debiased matrices for communication efficiency.
+Due to orthogonal ambiguity, we propose aligning the local eigenspaces with OPT (Cape (2020); Charisopoulos et al.
+(2021)).
+It is worth mentioning that the algorithm is general because it includes the special cases where the number of local machines is one or the privacy-preserving constraint disappears.
+Second, we provide novel bounds of the distributed algorithm ppDSC regarding the eigenvector estimation and clustering, which explicitly show how the misclassification rate depends on the privacy parameters.
+In particular, when the privacy budget in DP increases logarithmically with the number of nodes, the error bounds coincide with those without the privacy-preserving constraint.
+Two interesting findings are observed from the privacy-aware bounds.
+First, without a pre-specified privacy budget, perturbing the network entries from 0 to 1 is more harmful than perturbing the network from 1 to 0 to the upper error bound.
+Second, with a given privacy budget, from our theoretical bounds, one can find the optimal edge-flipping probabilities in the RR mechanism.
+To the best of our knowledge, these new theoretical results for RR-perturbed DP algorithms have not been investigated in the SBM literature.
+In addition, the conditions for the bounds reveal that denser networks can allow a more strict privacy-preserving level, i.e., a smaller privacy budget.
+Last but not least, the bounds show how local computation and heterogeneity affect the performance of ppDSC.
+Note that in the regression set-up, the one-shot-based distributed algorithm can attain the centralized bound when the number of machines satisfies a polynomial order of the sample size on each local machine; see Zhang et al.
+(2013); Lee et al.
+(2017), among many others.
+Similarly, we show that the error bound for our ppDSC decomposes into a term that corresponds to the centralized error and two remainder terms corresponding to the squared local error and the effective heterogeneity.
+The two remaining terms are negligible when the number of machines satisfies a condition similar to the one in the regression setting, and the effective heterogeneity is not severe.
+Then, the estimator converges at the same rate as a centralized estimator.
+Our work is the first to derive an error bound for a privacy-preserving and distributed learning algorithm in multi-layer SBMs.
+1.1 Related work
+Recent years have seen increasing work on multi-layer SBMs including Han et al.
 (2015); Paul and Chen (2016); Lei et al.
-(2020); Arroyo et al.
-(2021); MacDonald et al.
-(2022); Lei and Lin (2022); Huang et al.
-(2022).
-However, in practice, the multi-layer network is often locally stored because of the private ownership and the communication cost among others (Dwork et al., 2014; Kairouz et al., 2021).
-For example, consider a multi-layer friendship network obtained from different social platforms, where the nodes are the common users of these platforms, and the layers correspond to different social ties.
-In this network, neither the users would like others to know what kind of users they are following completely, nor are the platforms willing to share all the network information because of the ownership.
-Thus, aggregating all network layers into one central machine would be rather difficult in this scenario.
-To this end, the analysis of the multi-layer networks should be naturally based on a distributed and federated architecture.
-In the literature, the multi-layer stochastic block model (multi-layer SBM) (Han et al., 2015; Paul and Chen, 2016; Barbillon et al., 2017) has been popularly used for community detection in multi-layer networks, in which each layer is assumed to share the same unobserved community membership (Greene and Cunningham, 2013; Liu et al., 2013).
-There are two major methodological challenges to apply the multi-layer SBM for locally stored networks with privacy-preserving constraints.
-First, because of the privacy budget, the randomized response (RR) mechanism (Warner, 1965; Karwa et al., 2017; Ayed et al., 2020) that simply flips all edges in the network with a certain probability is often adopted to achieve differential privacy (DP) (Dwork et al., 2006, 2014; Mironov, 2017; Dong et al., 2022; Bi and Shen, 2022).
-Although this method can greatly protect an individual’s privacy by ensuring that no individual significantly influences the algorithm’s output, the RR mechanism would bring bias into the multi-layer SBM.
-The existing community detection methods (e.g., Paul and Chen (2020); Arroyo et al.
-(2021); Lei and Lin (2022); Huang et al.
-(2022)) proposed for the exact multi-layer SBMs cannot be directly applied in this scenario.
-Therefore, a bias-correction procedure is needed for community detection and estimation in the composite and RR-perturbed multi-layer SBM.
-Second, directly sending the RR-perturbed networks, either with or without bias-correction on the local machines to the central server, can bring large communication cost or storage cost on the server, especially for large networks.
-When the bias-correction is achieved on the local machines, the networks become dense; when it is performed on the central server, it needs to preserve all networks.
-Moreover, sending non-debiased sparse networks can be also communication expensive when the network is large.
-Therefore, a communication-efficient algorithm is of importance in our considered setting.
-Motivated by the above problems, we propose a privacy-preserving Distributed Spectral Clustering (ppDSC) algorithm for detecting the communities of locally stored multi-layer networks with private preserving.
-To avoid possible cancellation of the communities among different layers, we adopt the squared adjacency matrix (Lei and Lin, 2022) as the input of ppDSC.
-Our proposed algorithm consists of four steps given below.
-  • Step 1: we conduct RR on the network located in each local machine to achieve DP.
-  • Step 2: we perform bias-correction on each local RR-perturbed network and then obtain a bias-corrected network on each machine.
-  • Step 3: we perform eigen-decomposition on each bias-corrected network and send the top-eigenvectors of each network to the central server.
-Due to the orthogonal ambiguity of subspaces, we average the orthogonal transformed local eigenvectors using the orthogonal Procrustes transformation (OPT) (Schönemann, 1966; Cape, 2020; Charisopoulos et al., 2021) on the central server.
-This method ensures that the transformed local eigen-vectors are aligned with each other.
-  • Step 4: we implement $k$-means clustering on the orthogonalized vector on the central server to obtain the estimated communities.
-We provide rigorous non-asymptotic analysis of the statistical errors of ppDSC in terms of eigen-vector estimation and clustering.
-It turns out that both errors consist of three parts: the squared local, heterogeneity-induced and centralized errors.
-The privacy effect is inherent in the first and third parts.
-1.1 Our contributions and related work
-The contribution of this paper is multi-fold, summarized as follows.
-First, we propose a novel community detection method for multiple networks collected from different sources with privacy concerns.
-Our algorithm satisfies the strong notion of DP.
-For network data, two notions of DP are often used (Rastogi et al., 2009); one is edge-DP, considered in our paper; the other one is node-DP, which conceals the inclusion or removal of a particular node with all its adjacent edges.
-Under the framework of these two notions, DP algorithms have been developed for several problems, including graph statistics releasing (Nissim et al., 2007; Karwa et al., 2011; Imola et al., 2021), synthetic graphs releasing (Qin et al., 2017), and generative graph models estimation (Mir and Wright, 2012; Borgs et al., 2015; Karwa and Slavković, 2016; Chang et al., 2021), among others.
-However, there are few studies on privacy-preserving community detection using statistical models, e.g., Nguyen et al.
-(2016); Ji et al.
-(2019); Hehir et al.
-(2021); Seif et al.
-(2022).
-In particular, to the best of our knowledge, privacy-preserving community detection for multi-layer networks has not been studied yet.
-Second, we develop a bias-correction procedure to remove the bias brought by RR in the multi-layer SBM, and theoretically investigate the effect of the bias-correction on community detection.
-The previous works focused on deriving the information-theoretical limits (Seif et al., 2022) and misclassification rate of community detection in bias-corrected RR-perturbed SBMs (Hehir et al., 2021).
-However, how the bias-correction procedure can affect the clustering and estimation accuracy is still unclear, which is studied in our paper through rigorous theoretical justification.
-Third, we develop a communication-efficient algorithm, ppDSC, by which we only need to send the top eigenvectors from the local machines to the central server.
-The method is inspired by the distributed PCA proposed for other scenarios in the divide-and-conquer paradigm (Fan et al., 2019; Bhaskara and Wijewardena, 2019; Charisopoulos et al., 2021) and iterative paradigm (Garber et al., 2017; Gang et al., 2019; Chen et al., 2021a; Li et al., 2021; Guo et al., 2021).
-However, we incorporate the perturbation and bias-correction procedure in our algorithm; as a result, it poses great challenges to derive the theories, including the estimation error bounds.
-Last but not least, we provide a series of novel bounds and explicitly show how the bounds depend on the privacy parameters.
-These theoretical results have not yet been established.
-In the special case without the privacy and local computation demand, the error bounds of our algorithm reduce to that given in Lei and Lin (2022).
-The bounds also review the effect of RR on community detection.
-In particular, adding more network edges can lead to larger error bounds than removing edges.
-Our established bounds can also help to explain the blessings and curses of heterogeneity among the local networks.
+(2020); Bhaskara and Wijewardena (2019); Paul and Chen (2020); Arroyo et al.
+(2021); Lei and Lin (2022); Jing et al.
+(2021), among many others.
+A common assumption in these works is that all network layers are centrally stored without considering privacy concerns.
+As a result, the community detection is performed directly on the original data on the central server.
+Different from the aforementioned works, we consider the federated learning setting in which the network layers are stored locally on each machine with privacy preservation need.
+To improve communication efficiency, we propose to work with the locally computed eigenvectors instead of the original data, which results in both methodological and theoretical differences between our methods and the aforementioned works.
+There have been various DP algorithms developed for analyzing network data, including graph statistics releasing (Nissim et al., 2007; Karwa et al., 2011; Imola et al., 2021), synthetic graphs releasing (Qin et al., 2017), generative graph models estimation (Mir and Wright, 2012; Borgs et al., 2015; Karwa and Slavković, 2016; Chang et al., 2021), community detection (Nguyen et al., 2016; Ji et al., 2019) among others.
+These algorithms are not developed based on SBMs.
+Under the single SBM, two recent works Hehir et al.
+(2021) and Mohamed et al.
+(2022) studied respectively the misclassification rate and information-theoretical limits of community detection under the RR mechanism.
+The differences between our work and these two works are as follows.
+First, we study the multi-layer SBM while they studied the single SBM.
+Second, we derive several new privacy-aware bounds that reflect the asymmetric roles of the two edge-flipping probabilities in RR and indicate the best choice of the edge-flipping probabilities given a fixed privacy budget.
+These observations are new to the literature.
+In addition, we theoretically show the effect of bias-correction, which was previously unclear.
+Driven by large decentralized datasets, there has been increasing work on distributed algorithms for PCA, including the divide-and-conquer paradigm (Fan et al., 2019; Bhaskara and Wijewardena, 2019; Charisopoulos et al., 2021) and iterative paradigm (Garber et al., 2017; Gang et al., 2019; Chen et al., 2021; Guo et al., 2024).
+However, we work with privacy-preserved network data and incorporate the perturbation and bias-correction procedure in our algorithm; as a result, it poses great challenges to derive the statistical theories, including the estimation error bounds.
+1.2 Notation
+We use $\lambda_{\min}(A)$ to denote the smallest non-zero eigenvalue of a given matrix $A$.
+Moreover, we use the projection distance $\mathrm{dist}(U,U^{\prime}):=\|UU^{\intercal}-U^{\prime}U^{\prime{\intercal}}\|_{2}$ to measure the subspace distance between any pair of orthogonal basis $U,U^{\prime}\in\mathbb{R}^{n\times K}$.
+$\|A\|_{2}$ denotes the spectral norm of the matrix $A$ or the Euclidean norm of vector $A$ and $\|A\|_{\tiny F}$ denotes the Frobenius norm of matrix $A$.
+In addition, we use the following standard notation for asymptotics.
+We write $f(n)\asymp g(n)$ or $f(n)=\Theta(g(n))$ if $cg(n)\leq f(n)\leq Cg(n)$ for some constants $0<c<C<\infty$, $f(n)\lesssim g(n)$ or $f(n)=O(g(n))$ if $f(n)\leq Cg(n)$ for some constant $C<\infty$, and $f(n)\gtrsim g(n)$ or $f(n)=\Omega(g(n))$ if $f(n)\geq cg(n)$ for some constant $c>0$.
+Finally, the constants throughout this work may be different from place to place.
 1.3 Organizations
 The rest of the paper is organized as follows.
-Section 2 introduces the preliminaries of the model and DP.
+Section 2 provides the preliminaries of the model and DP.
 Section 3 develops the algorithm for privacy-preserving community detection of locally stored networks.
@@
 2 Problem formulation
-In this section, we first present the multi-layer SBMs for modeling the locally stored networks. Next, we introduce the DP and the RR mechanism for privacy protection.
+In this section, we first present the multi-layer SBMs for generating the locally stored networks. Next, we provide the formal definition of the DP and RR mechanisms.
 2.1 The model for locally stored networks
-Suppose that $L$ aligned networks with $n$ common nodes are locally stored in $L$ machines, respectively.
-The networks are binary and symmetric with the adjacency matrix denoted by $A_{l}\in\{0,1\}^{n\times n},l=1,...,L$.
-We consider the scenario that $A_{l}$’s share common communities but with possibly different edge densities.
-To model the above locally stored networks, we assume that $A_{l}$’s are independently generated from the following multi-layer SBMs.
-The $n$ nodes are assigned to $K$ non-overlapping communities with the community assignment of node $i$ denoted by $g_{i}\in\{1,...,K\}$.
+Suppose there are $L$ aligned networks with $n$ common nodes.
+The networks are binary and symmetric with the adjacency matrix denoted by $A_{l}\in\{0,1\}^{n\times n},\;l\in[L]:=\{1,2,\dots,L\}$.
+In the multi-layer SBMs (Han et al., 2015; Paul and Chen, 2016), each $A_{l}$ is independently generated as follows.
+The $n$ nodes are assigned to $K$ non-overlapping communities with the community assignment of node $i$ denoted by $g_{i}\in[K]$.
 Given the community assignments, each entry $A_{l,ij}\ (i<j)$ of $A_{l}$ is generated independently according to
-$A_{l,ij}\sim{\mathrm{Bernoulli}}(B_{l,g_{i}g_{j}}),\quad i<j,\;l=1,...,L,$ (1)
-where $B_{l}$’s are link probability matrices, and the diagonal entries $A_{l,ii}$’s are all 0.
-The heterogeneity among the networks is inherent in $B_{l}$’s, and we will quantify it in the next section.
-Note that we assume that $A_{l}$’s are generated locally at the machine $L$’s.
-Model (1) is the multi-layer SBMs studied in Han et al.
-(2015); Paul and Chen (2016) and so forth.
-When the adjacency matrices $A_{l}$’s are generated and stored in the local machines, we consider DP and the RR mechanism for the adjacency matrices for private protection.
+$A_{l,ij}\sim{\mathrm{Bernoulli}}(B_{l,g_{i}g_{j}}),\quad i<j,\;l\in[L],$ (1)
+where $B_{l}$’s are connectivity matrices, and the diagonal entries $A_{l,ii}$’s are all 0. The heterogeneity among the networks is inherent in $B_{l}$’s, and we will quantify it in the next section.
+We assume $A_{l\in[L]}$’s are locally generated and stored at $m$ different machines.
+For notational simplicity, we assume each machine evenly owns $s:=L/m$ networks.
+We use $S_{i}:=\{(i-1)s+1,...,i\cdot s\}$ to denote the index set of networks stored in the $i$th machine.
+Given the generation process of $A_{l}$’s, we consider the RR mechanism for the adjacency matrices for private protection in the context of DP (Dwork et al., 2014).
 2.2 Differential privacy and randomized response
-Though the local networks $A_{l}$’s are not directly shared by all the local machines, potential privacy breaches still exist (Dwork et al., 2017).
+Though the local networks $A_{l\in S_{i}}$’s are not directly shared by all the local machines, potential privacy breaches still exist (Dwork et al., 2017).
 DP provides a rigorous framework for privacy-preserving data analysis (Dwork et al., 2006).
@@
 By definition, DP protects any individual’s privacy from an adversary who has access to the algorithm output and even sometimes the rest of the data.
-In addition, it can prevent other kind of attacks, say, the re-identification attack, the tracing attack, and the membership attack, among others; see Dwork et al.
+In addition, it can prevent other kinds of attacks, say, the re-identification attack, the tracing attack, and the membership attack, among others; see Dwork et al.
 (2017) for a survey.
@@
 Edge-DP protects the privacy of network edges by ensuring that the algorithm’s output does not reveal the inclusion or removal of any certain edge in the network.
-In other words, edge-DP can protect against an adversary who knows the algorithm’s output and the existence or not of edges between all pairs but one pair of nodes.
+In other words, edge-DP can protect against an adversary who knows the algorithm’s output and the existence or absence of edges between all pairs but one pair of nodes.
 The $\epsilon$ is often termed as the privacy budget.
@@
 The algorithm outputs a perturbed adjacency matrix $\tilde{A}$ of $A$.
-Specifically, for each $i<j$, when $A_{ij}=1$, $\tilde{A}_{ij}$ keeps to be 1 with probability $q$ and flips to 0 otherwise.
-Similarly, when $A_{ij}=0$, ${\tilde{A}}_{ij}$ remains to be 0 with probability $q^{\prime}$ and changes to 1 otherwise.
+Specifically, for each $i<j$, when $A_{ij}=1$, $\tilde{A}_{ij}$ remains 1 with probability $q$ and flips to 0 otherwise.
+Similarly, when $A_{ij}=0$, ${\tilde{A}}_{ij}$ remains 0 with probability $q^{\prime}$ and changes to 1 otherwise.
 The diagonal entries ${\tilde{A}}_{ii}=0$.
@@
 $\epsilon\geq\log\max\left\{\frac{q^{\prime}}{1-q},\frac{1-q}{q^{\prime}},\frac{1-q^{\prime}}{q},\frac{q}{1-q^{\prime}}\right\}.$ (3)
-According to (3), a typical choice of $q$ and $q^{\prime}$ is that
+According to (3), we can easily derive that for a given $\epsilon$,
+[…]
+We plot the feasible region of $q$ and $q^{\prime}$ for a given $\epsilon$ in Figure 1, where the formulation of the four boundaries is shown in red.
+For fixed $\epsilon$, larger $q$ and $q^{\prime}$ will lead to less accuracy loss.
+Hence, the upper boundaries, i.e., $q^{\prime}=1-\frac{q}{\mathrm{e}^{\epsilon}}$ when $q\leq\frac{\mathrm{e}^{\epsilon}}{1+\mathrm{e}^{\epsilon}}$ and $q^{\prime}=\mathrm{e}^{\epsilon}(1-q)$ when $q\geq\frac{\mathrm{e}^{\epsilon}}{1+\mathrm{e}^{\epsilon}}$, are more desirable.
+A typical choice of $q$ and $q^{\prime}$ is that
 $q=q^{\prime}=\frac{\mathrm{e}^{\epsilon}}{1+\mathrm{e}^{\epsilon}},$ (4)
 which has been proven to be optimal in maximizing the mutual information between raw and RR-perturbed data (Ayed et al., 2020).
+More importantly, it turns out that this choice is also optimal in the context of this work in the sense that it leads to the smallest misclassification upper bound; see discussions after Theorem 3.
+Figure 1: The feasible region (in blue) of $q$ and $q^{\prime}$ under a given privacy budget $\epsilon$.
 Although RR can achieve edge-DP, one can imagine that the resulting network largely deviates from the original network.
-In the next section, we will see how we can debias the perturbed networks so as to improve the algorithm’s utility under the multi-layer SBMs.
+In the next section, we will see how we can de-bias the perturbed networks to improve the algorithm’s utility under the multi-layer SBMs.
 3 Privacy-preserving community detection for locally distributed networks
 This section proposes the privacy-preserving community detection algorithm for locally distributed networks.
-Before that, we specify the centralized estimator and develop a bias-adjustment procedure to remove the bias inherent in the RR-perturbed networks.
-Notes and notation.
-We use $\lambda_{\min}(A)$ to denote the smallest non-zero eigenvalue of a given matrix $A$.
-Moreover, we use the projection distance $\mathrm{dist}(U,U^{\prime}):=\|UU^{\intercal}-U^{\prime}U^{\prime{\intercal}}\|_{2}$ to measure the subspace distance between any pair of orthogonal basis $U,U^{\prime}\in\mathbb{R}^{n\times K}$.
-$\|A\|_{2}$ denotes the spectral norm of matrix $A$ or the Euclidian norm of vector $A$ and $\|A\|_{\tiny F}$ denote the Frobenius norm of matrix $A$.
-In addition, we use the following standard notation for asymptotics.
-We write $f(n)\asymp g(n)$ or $f(n)=\Theta(g(n))$ if $cg(n)\leq f(n)\leq Cg(n)$ for some constants $0<c<C<\infty$, $f(n)\lesssim g(n)$ or $f(n)=O(g(n))$ if $f(n)\leq Cg(n)$ for some constant $C<\infty$, and $f(n)\gtrsim g(n)$ or $f(n)=\Omega(g(n))$ if $f(n)\geq cg(n)$ for some constant $c>0$.
-Finally, the constants throughout this work may be different from place to place.
-Recall that there are $L$ local machines, with each owning a symmetric binary network adjacency matrix $A_{l},l=1,...,L$.
-$A_{l}$’s are assumed to be generated from a multi-layer SBM.
+In particular, we develop a bias-adjustment procedure to remove the bias inherent in the RR-perturbed networks.
+Recall that there are $m$ local machines, and the $i$th machine owns $s$ adjacency matrices $A_{l\in S_{i}}$’s.
+$A_{l\in[L]}$’s are assumed to be generated from a multi-layer SBM.
 Denote the population of $A_{l}$ by $P_{l}:=\Theta B_{l}\Theta^{\intercal}$, where $\Theta\in\mathbb{R}^{n\times K}$ is the membership matrix which is formed by forcing $\Theta_{ig_{i}}=1$ and $\Theta_{ij}=0$ for $j\neq g_{i}$.
 We aim to recover the common membership matrix $\Theta$ using the network information.
-Without the privacy constraint, a common centralized approach is to perform community detection algorithms on $\sum_{l}A_{l}$ (Bhattacharyya and Chatterjee, 2018; Paul and Chen, 2020).
-However, direct summation of $A_{l}$’s would lead to information loss when $A_{l}$’s are heterogeneous in the sense of community pattern.
-In other words, if some $P_{l}$’s are positive semidefinite, which present homophily social ties and others are negative semidefinite, which indicate heterophily social ties (Kim and Leskovec, 2012), then adding them up would cancel out some community signals.
-To alleviate this, we consider the matrix
-$M:=\frac{1}{Ln}\sum_{l=1}^{L}A_{l}^{2}$ (5)
-studied in Lei and Lin (2022), which could accommodate different community patterns.
-Note that Lei and Lin (2022) further developed a bias-correction procedure to modify the diagonal elements of $M$.
-The centralized estimator is then obtained by performing spectral clustering on $M$.
-Specifically, the eigen-decomposition is first computed on $M$ to obtain the first $K$ eigenvectors, and then $k$-means is conducted on the resulting eigenvectors to obtain $K$ communities.
-Hopefully, $M$ can approximate its population counterpart
-${Q}:=\frac{1}{L}\sum_{l=1}^{L}\frac{P_{l}^{2}}{n}:=\frac{1}{L}\sum_{l=1}^{L}{Q}_{l},$ (6)
-whose eigenvector reveals the true communities in the sense that two rows of the eigenvector matrix are identical if and only if the corresponding nodes are in the same underlying communities (Rohe et al., 2011; Lei and Rinaldo, 2015; Zhang et al., 2022).
-However, with the privacy constraint, we could only use the RR-perturbed adjacency matrix instead of $A_{l}$ in (5), which is a biased estimator of $P_{l}$.
-In the next subsection, we will develop a bias-adjustment procedure to obtain a good estimator for $Q$.
+Since $P_{l}$’s are not necessarily positive semidefinite, the top-$K$ singular vectors of $P_{l}$’s are a good choice to reveal the communities with both assortative and disassortative structures (Rohe et al., 2011), which is equivalent to the top-$K$ eigenvectors of $P_{l}^{2}$.
+In the following, we will develop a bias-adjustment procedure to obtain a good estimator for $P_{l}^{2}$ when we only have access to the RR-perturbed adjacency matrices.
 3.1 Bias-adjustment
-Our debiasing procedure involves two steps: in step 1, we build a debiased estimator for the link probability matrix upon the RR-perturbed adjacency matrix stored in each local machine; in step 2, we further remove the bias of the diagonals caused by taking the square of the debiased estimated matrix obtained in step 1.
-Step 1. We now illustrate the necessity of bias-adjustment when dealing with RR-perturbed data.
+Our debiasing procedure involves two steps: in step 1, we build a debiased estimator for the connectivity matrix upon the RR-perturbed adjacency matrix stored in each local machine; in step 2, we further remove the bias of the diagonals caused by taking the square of the debiased estimated matrix obtained in step 1.
+Step 1. We now illustrate the necessity of bias adjustment when dealing with RR-perturbed data.
 For $1\leq l\leq L$, we denote the RR-perturbed $A_{l}$ by ${\tilde{A}}_{l}$. Specifically, for $i<j$, if $A_{l,ij}=1$,
@@
 […]
-and ${\tilde{A}}_{l,ji}={\tilde{A}}_{l,ij}$. By (7) and (8) and taking expectations for two times, we have
+and ${\tilde{A}}_{l,ji}={\tilde{A}}_{l,ij}$. By (5) and (6), taking expectations for two times, we have
 $\mathbb{E}({\tilde{A}}_{l,ij})=P_{l,ij}(q+q^{\prime}-1)+1-q^{\prime},$
-which is biased from the true expectation $P_{l,ij}$ of $A_{l,ij}$. Therefore, for $i\neq j$, the local machine only needs to debias ${\tilde{A}}_{l,ij}$ to be
+which is biased from the true expectation $P_{l,ij}$ of $A_{l,ij}$. Therefore, for $i\neq j$, the local machine only needs to de-bias ${\tilde{A}}_{l,ij}$ to be
 $\overline{A}_{l,ij}=\frac{{\tilde{A}}_{l,ij}-(1-q^{\prime})}{q+q^{\prime}-1}.$
@@
 In particular, this debiasing step is crucial when $L$ is increasing.
-For the sake of saving space, the details are relegated to Appendix A.
-Combining the two steps of debiasing, we eventually obtain the following estimator $\overline{M}$ for $Q$ defined in (6),
-$\overline{M}:=\frac{1}{L}\sum_{l=1}^{L}\overline{M}_{l}$ (10)
-with
+To save space, the details are relegated to Section A of Supplementary Materials.
+Combining the two steps of debiasing, we eventually obtain the following estimator $\overline{M}_{l}$ for $P_{l}^{2}/n$,
 $\overline{M}_{l}:=\frac{\overline{A}_{l}^{2}}{n}-\frac{(q^{\prime})^{2}}{n(q+q^{\prime}-1)^{2}}G_{l},$ (8)
@@
 Note that the second step of debiasing is in a similar spirit of Lei and Lin (2022) except that their debiasing is performed on the original network matrix, while ours is on the debiased RR-perturbed network matrix.
-Therefore, with the privacy constraint, the ideal centralized estimator is to perform spectral clustering on the debiased estimator $\overline{M}$.
-Noting that $\overline{M}_{l}$’s are separable across different local machines.
-We show in the next subsection how we can approximate the eigenvector of $\overline{M}$ with the local computation.
 3.2 Privacy-preserving distributed spectral clustering
-Since DP enjoys the post-processing property, in principle, the local machines can send the dense debiased estimators $\overline{M}_{l}$’s to the central server where the spectral clustering is further performed on the aggregated matrix $\overline{M}$.
-However, both the communication and computational cost can be very high for sending the whole dense matrices $\overline{M}_{l}$’s and performing the aggregation of these matrices on the central server, especially when the matrix dimensions are large.
-Instead, we consider a more efficient distributed algorithm described as follows.
+Since DP enjoys the post-processing property, in principle, each local machine can send the averaged estimator $\sum_{l\in S_{i}}\overline{M}_{l}/s$ to the central server where the spectral clustering is further performed on the aggregated matrix
+$\overline{M}:=\frac{1}{m}\sum_{i=1}^{m}\sum_{l\in S_{i}}\frac{\overline{M}_{l}}{s}.$ (9)
+However, both the communication and computational cost can be very high for sending the whole dense matrices $\sum_{l\in S_{i}}\overline{M}_{l}/s$’s and performing further aggregation of these matrices on the central server, especially when the matrix dimensions are large.
+Instead, we consider a more efficient distributed algorithm, as described below.
 We adopt the one-shot averaging framework in Charisopoulos et al.
 (2021) for distributed computing performed on our debiased RR-perturbed network matrices.
-That is, on the $l$th machine, we first compute the eigen-decomposition of $\overline{M}_{l}$, and obtain the matrix $\hat{V}_{l}\in\mathbb{R}^{n\times K}$ that consists of the first $K$ eigenvectors, for all $l=1,...,L$.
+That is, on the $i$th machine, we first compute the eigen-decomposition on $\sum_{l\in S_{i}}\overline{M}_{l}/s$, and obtain the matrix $\hat{V}_{i}\in\mathbb{R}^{n\times K}$ that consists of the first $K$ eigenvectors, for all $i\in[m]$.
 On the central server, we average all the local solutions.
-Note that direct averaging is inferior due to the orthogonal ambiguity of the problem, i.e., for any orthogonal matrix $Z$, $\hat{V}_{l}Z$ also forms a valid orthonormal basis of the subspaces formed by $\hat{V}_{l}$.
-Hence, $\hat{V}_{l}$’s are not guaranteed to be close to each other, even if their corresponding subspaces are close.
-In light of this, the local $\hat{V}_{l}$’s must be aligned with each other before averaging.
-To avoid the orthogonal ambiguity of the problem, we use the Procrustes fixing strategy (Schönemann, 1966; Cape, 2020; Charisopoulos et al., 2021; Guo et al., 2021).
+Note that direct averaging is inferior due to the orthogonal ambiguity of the problem, i.e., for any orthogonal matrix $Z$, $\hat{V}_{i}Z$ also forms a valid orthonormal basis of the subspaces formed by $\hat{V}_{i}$.
+Hence, $\hat{V}_{i}$’s are not guaranteed to be close to each other, even if their corresponding subspaces are close.
+In light of this, the local $\hat{V}_{i}$’s must be aligned with each other before averaging.
+To avoid the orthogonal ambiguity of the problem, we use the Procrustes fixing strategy (Schönemann, 1966; Cape, 2020; Charisopoulos et al., 2021; Guo et al., 2024).
 For two matrices $V_{1},V_{2}$ to be aligned, we can map $V_{2}$ to $V_{1}$ using an orthogonal transformation solved by the following optimization,
@@
 Let the SVD of $V_{2}^{\intercal}V_{1}$ be $LDR^{\intercal}$, then $Z:=LR^{\intercal}$ is the solution to (10).
-In such a way, we can align local eigenspaces $\hat{V}_{l}$ to an arbitrarily chosen reference eigenspace, for example, $\hat{V}_{o}$, after which the aligned eigenspaces can then be averaged and orthogonalized.
+In such a way, we can align local eigenspaces $\hat{V}_{i}$ to an arbitrarily chosen reference eigenspace, for example, $\hat{V}_{o}$, after which the aligned eigenspaces can then be averaged and orthogonalized.
+As we will see in Section 4, the selection of the reference eigenspace does not affect the theoretical bounds.
+Algorithm 1 Privacy-preserving distributed spectral clustering (ppDSC)
+1:  Input: Network adjacency matrices $\{A_{l}\}_{l=1}^{L}$, number of machines $m$, number of clusters $K$, privacy parameters $q,q^{\prime}$, index of reference solution $o$.
+2:  The local machines do:
+3:  for $i=1$ to $m$ do
+4:     for $l$ in $S_{i}$ do
+5:        RR-perturbation: Perturb $A_{l}$ to obtain ${\tilde{A}}_{l}$ using (5) and (6);
+6:        Bias-correction: Constructs the debiased estimator $\overline{M}_{l}$ according to (8);
+7:     end for
+8:     Eigen-decomposition: Performs eigen-decomposition on $\sum_{l\in S_{i}}\overline{M}_{l}/s$ and denote the resulting eigen-spaces as $\hat{V}_{i}$.
+9:  end for
+10:  The central server does:
+11:  Procrustes fixing: $\tilde{V}_{i}:=\hat{V}_{i}Z_{i}$ with $Z_{i}:=\operatorname{arg\,min}_{Z\in\mathcal{O}_{K}}\|\hat{V}_{i}Z-\hat{V}_{o}\|_{F}$ for each $i=1,...,m$;
+12:  Aggregation: Aggregates the local solutions according to $\bar{V}:=\frac{1}{m}\sum_{i=1}^{m}\tilde{V}_{i}$;
+13:  Orthogonalization: Performs the QR decomposition to obtain $\tilde{V},\tilde{R}={\rm qr}(\bar{V})$.
+14:  Clustering: Conduct $k$-means clustering on $\tilde{V}$ with cluster number being $K$. The resulting clusters are formulated by the membership matrix $\tilde{\Theta}$.
+15:  Output: The approximated eigen-space $\tilde{V}$ and the estimated membership matrix $\tilde{\Theta}$.
 The proposed privacy-preserving Distributed Spectral Clustering (ppDSC) algorithm is summarized in Algorithm 1.
@@
 For example, a party with data ownership may desire to obtain good utility while preserving the data’s privacy.
-Thus all the local steps are conducted within the party.
+Thus, all the local steps are conducted within the party.
 It can also happen that the RR perturbation is conducted by the data collection department of a company, while the data analysis department could do the remaining bias-correction and eigen-decomposition steps.
@@
 where $\Delta(A_{\mathrm{all}},A_{\mathrm{all}}^{\prime})=1$ means that all layers of the multi-layer network pairs $A_{\mathrm{all}}$ and $A_{\mathrm{all}}^{\prime}$ are equal except one edge of one layer.
-Algorithm 1 Privacy-preserving distributed spectral clustering (ppDSC)
-1:  Input: Network adjacency matrices $\{A_{l}\}_{l=1}^{L}$ owned by local machines, number of clusters $K$, privacy parameters $q,q^{\prime}$, index of reference solution $o$.
-2:  The local machines do:
-3:  for $l=1$ to $L$ do
-4:     RR-perturbation: Perturb $A_{l}$ to obtain ${\tilde{A}}_{l}$ using (7) and (8);
-5:     Bias-correction: Constructs the debiased estimator $\overline{M}_{l}$ according to (11);
-6:     Eigen-decomposition: Performs eigen-decomposition on $\overline{M}_{l}$ and denote the resulting eigen-spaces as $\hat{V}_{l}$.
-7:  end for
-8:  The central server does:
-9:  Procrustes fixing: $\tilde{V}_{l}:=\hat{V}_{l}Z_{l}$ with $Z_{l}:=\operatorname{arg\,min}_{Z\in\mathcal{O}_{K}}\|\hat{V}_{l}Z-\hat{V}_{o}\|_{F}$ for each $l=1,...,L$;
-10:  Aggregation: Aggregates the local solutions according to $\bar{V}:=\frac{1}{L}\sum_{l=1}^{L}\tilde{V}_{l}$;
-11:  Orthogonalization: Performs the QR decomposition to obtain $\tilde{V},\tilde{R}={\rm qr}(\bar{V})$.
-12:  Clustering: Conduct $k$-means clustering on $\tilde{V}$ with cluster number being $K$. The resulting clusters are formulated by the membership matrix $\tilde{\Theta}$.
-13:  Output: The approximated eigen-space $\tilde{V}$ and the estimated membership matrix $\tilde{\Theta}$.
 4 Theoretical analysis
@@
 To fix the ideas, we now recall and introduce some notations.
-We assume that there are $L$ local networks whose adjacency matrices $A_{l}\in\{0,1\}^{n\times n}$’s are generated from SBMs with layer-specific link probability matrices $B_{l}$’s and a consensus membership matrix $\Theta\in\mathbb{R}^{n\times K}$.
+We assume that there are $L$ networks with adjacency matrices being $A_{l}\in\{0,1\}^{n\times n}$’s evenly distributed on $m$ local machines; each machine owns $s$ networks.
+For $i\in[m]$, the $i$th machine has networks indexed by set $S_{i}$.
+The adjacency matrices $A_{l}$’s are generated from SBMs with layer-specific connectivity matrices $B_{l}$’s and a consensus membership matrix $\Theta\in\mathbb{R}^{n\times K}$.
 We here assume that the number of communities $K$ is known.
@@
 The privacy parameters are denoted by $q,q^{\prime}$ ($1/2<q,q^{\prime}\leq 1$); see also (5) and (6).
-Our goal is to use ppDSC (Algorithm 1) to recover the first $K$ eigenvectors $V$ of the underlying target matrix $Q$ defined in (6) and the membership matrix $\Theta$.
-The following assumptions are generally required to derive the theoretical results for community detection.
+Our goal is to use ppDSC (Algorithm 1) to recover the first $K$ eigenspaces $V$ of the underlying target matrix $Q$ defined by
+${Q}:=\frac{1}{L}\sum_{l=1}^{L}\frac{P_{l}^{2}}{n}:=\frac{1}{L}\sum_{l=1}^{L}{Q}_{l},$ (11)
+and the membership matrix $\Theta$.
+The following assumptions and definition are generally required to derive the theoretical results for community detection.
 Assumption 1 (Balanced communities)
-The community is balanced such that for any $k\in\{1,...,K\}$,
-$c\frac{n}{K}\leq n_{k}\leq C\frac{n}{K},$
-for some constants $0\leq c\leq C$.
+The community is balanced such that for any $k\in\{1,...,K\}$, $c\frac{n}{K}\leq n_{k}\leq C\frac{n}{K}$ for some constants $0\leq c\leq C$.
 Assumption 2 (Sparsity)
-The link probability matrix $B_{l}=\rho B_{l,0}$, where $\rho=o(1)$ and $B_{l,0}$ is a $K\times K$ symmetric matrix with constant entries in $[0,1]$.
+The connectivity matrix $B_{l}=\rho B_{l,0}$, where $\rho=o(1)$ and $B_{l,0}$ is a $K\times K$ symmetric matrix with constant entries in $[0,1]$.
 Assumption 3 (Full rank)
 $\sum_{l}B_{l,0}^{2}$ is of full rank $K$ with $\lambda_{{\min}}(\sum_{l}B_{l,0}^{2})\geq cL$.
-Assumption 4 (Limited heterogeneity)
-The heterogeneity among the link probability matrices $B_{l}$’s are limited as
-$\max_{l}\|B_{l}\Delta^{2}B_{l}-\frac{1}{L}\sum_{l=1}^{L}B_{l}\Delta^{2}B_{l}\|_{2}\leq\mathcal{H}(n,\rho,L).$
-with $\Delta:={\operatorname{diag}}(\sqrt{n_{1}},...,\sqrt{n_{K}})$.
+Definition 2 (Heterogeneity)
+Suppose the connectivity matrix $B_{l}$’s are heterogeneous and define a heterogeneous measurement as
+$\mathcal{H}(n,\rho,L):=\max_{i\in[m]}\|\frac{1}{s}\sum_{l\in S_{i}}B_{l}\Delta^{2}B_{l}-\frac{1}{L}\sum_{l\in[L]}B_{l}\Delta^{2}B_{l}\|_{2},$
+where $\Delta:={\operatorname{diag}}(\sqrt{n_{1}},...,\sqrt{n_{K}})$.
 Remark 2
 Assumption 1 requires that the communities are balanced.
-Assumption 2 specifies the link probability matrices.
-In particular, we require that the network sparsity $\rho=o(1)$.
-Assumption 3 requires the linear growth of the aggregated squared link probability matrices in terms of its minimum eigenvalue, which is trivially satisfied when $B_{l,0}$’s are equal and is also used in Lei and Lin (2022).
-Assumption 3 combining with Assumptions 2 and 1 implies that $\sum_{l=1}^{L}B_{l}\Delta^{2}B_{l}$ is of full rank, and $Q$ is of rank $K$; see the proof of Corollary 1.
+Assumption 2 specifies the connectivity matrices.
+In particular, we require the network sparsity $\rho=o(1)$.
+Assumption 3 requires the linear growth of the aggregated squared connectivity matrices in terms of its minimum eigenvalue, which is trivially satisfied when $B_{l,0}$’s are equal and is also used in Lei and Lin (2022).
+Assumption 3 combining with Assumptions 2 and 1 implies that $\sum_{l=1}^{L}B_{l}\Delta^{2}B_{l}$ is of full rank, and $Q$ is of rank $K$; see the proof of Theorem 4.
 Indeed, the full rank requirement could be extended to the rank-deficient case; see Zhang et al.
 (2022) for example.
-Assumption 4 bounds the maximum deviation of $B_{l}\Delta^{2}B_{l}$’s from its average by $\mathcal{H}(n,\rho,L)$.
-For simplicity, we call $\mathcal{H}$ the measure for the heterogeneity of locally stored networks.
+Definition 2 measures how far the local average $\sum_{l\in S_{i}}B_{l}\Delta^{2}B_{l}/s$’s deviates from their global average $\sum_{l\in[L]}B_{l}\Delta^{2}B_{l}/L$.
+In what follows, we call $\mathcal{H}$ the measure for the heterogeneity.
 We first provide the following error decomposition theorem to present the general error bound for the output eigen-space of ppDSC.
 Theorem 2 (Error Decomposition)
-Suppose Assumptions 1 and 4 hold, the output $\tilde{V}$ of Algorithm 1 satisfies
-$\mathrm{dist}(\tilde{V},V)\lesssim{\max_{l}\frac{\|\overline{M}_{l}-Q_{l}\|_{2}^{2}}{\lambda_{{\min}}^{2}(Q)}}+{\frac{\mathcal{H}^{2}(n,\rho,L)}{\lambda_{{\min}}^{2}(Q)}}+\frac{\|\sum_{l}\overline{M}_{l}/L-Q\|_{2}}{\lambda_{{\min}}(Q)}:=\mathcal{I}_{1}+\mathcal{I}_{2}+\mathcal{I}_{3}.$
+Suppose Assumption 1 holds, then the output $\tilde{V}$ of Algorithm 1 satisfies
+$\mathrm{dist}(\tilde{V},V)\lesssim{\max_{i\in[m]}\frac{\|\sum_{l\in S_{i}}(\overline{M}_{l}-Q_{l})/s\|_{2}^{2}}{\lambda_{{\min}}^{2}(Q)}}+{\frac{\mathcal{H}^{2}(n,\rho,L)}{\lambda_{{\min}}^{2}(Q)}}+\frac{\|\sum_{l}\overline{M}_{l}/L-Q\|_{2}}{\lambda_{{\min}}(Q)}:=\mathcal{I}_{1}+\mathcal{I}_{2}+\mathcal{I}_{3}.$
 provided that both $\mathcal{I}_{1}$ and $\mathcal{I}_{2}$ are smaller than a small constant $c$.
 Remark 3
-The condition that $\mathcal{I}_{1}$ and $\mathcal{I}_{2}$ are smaller than a small constant $c$ is a standard condition (Damle and Sun, 2020; Charisopoulos et al., 2021); it ensures that the maximum local error $\max\|\overline{M}_{l}-Q\|_{2}$ is smaller than the eigen-gap $\lambda_{{\min}}(Q)$, which further indicates that the $K$-largest eigenvalues of $\overline{M}_{l}$ are separated from its remaining eigenvalues by Weyl’s inequality.
-In Corollary 1, we show that this condition is satisfied by our setting under mild conditions and further provide the explicit expressions for $\mathcal{I}_{1}$ and $\mathcal{I}_{2}$.
+The condition that $\mathcal{I}_{1}$ and $\mathcal{I}_{2}$ are smaller than a small constant $c$ is a standard condition (Damle and Sun, 2020; Charisopoulos et al., 2021); it ensures that the maximum error for the local average $\max_{i\in[m]}\|\sum_{l\in S_{i}}\overline{M}_{l}/s-Q\|_{2}$ is smaller than the eigen-gap $\lambda_{{\min}}(Q)$, which further indicates that the $K$th largest eigenvalues of $\sum_{l\in S_{i}}\overline{M}_{l}/s$ are separated from its remaining eigenvalues by Weyl’s inequality.
+In Theorem 4, we show that this condition is satisfied by our setting under mild conditions and further provide the explicit expressions for $\mathcal{I}_{1}$ and $\mathcal{I}_{2}$.
+The results of Theorem 2 hold for any reference eigenvector $\hat{V}_{o}$ provided that we choose the true eigenspace $V$ to be aligned with $\hat{V}_{o}$ in the sense that $\operatorname{arg\,min}_{Z\in\mathcal{O}_{K}}\|\hat{V}_{o}Z-V\|_{F}=\mathbb{I}_{K}$ by noting that we have a degree to choose the version of eigenspace $V$ (Charisopoulos et al., 2021).
+Empirically, since it is easy to see that each local machine has the same population-wise eigenspaces, the experimental results would be similar when choosing a different sample-wise reference solution $\hat{V}_{o}$.
+Remark 4
 The decomposed three components $\mathcal{I}_{1},\mathcal{I}_{2},\mathcal{I}_{3}$ are explained as follows.
-$\mathcal{I}_{1}$ corresponds to the maximum squared error of the local estimators $\hat{V}_{l}$’s except that we use $Q$ in the denominator instead of $Q_{l}$; see the Davis-Kahan theorem on the perturbation of eigenvectors (Davis and Kahan (1970); see Lemma 3 in the Appendix).
-As a result, we only require the averaged population matrix $Q:=\sum Q_{l}/L$ be of rank $K$, and the local population matrix $Q_{l}$ can be rank deficient, which is one of the benefits of combining the local eigenvectors.
-$\mathcal{I}_{2}$ corresponds to the degree of heterogeneity.
-To distinguish, we call $\mathcal{H}$ the measure for heterogeneity and call $\mathcal{I}_{2}$ the measure for scaled heterogeneity.
-It is worth noting that the scaled heterogeneity $\mathcal{I}_{2}$ affects the theoretical accuracy of $\tilde{V}$ according to its magnitude compared to $\mathcal{I}_{1}$ and $\mathcal{I}_{3}$.
+$\mathcal{I}_{1}$ corresponds to the maximum squared error of the local estimators $\hat{V}_{i}$’s except that we use $Q$ in the denominator instead of $\sum_{l\in S_{i}}Q_{l}/s$’s; see the Davis-Kahan theorem on the perturbation of eigenvectors (Davis and Kahan (1970); see Lemma 3 in the Supplementary Materials).
+As a result, we only require the averaged population matrix $Q:=\sum Q_{l}/L$ be of rank $K$, and the local averaged population matrix $\sum_{l\in S_{i}}Q_{l}/s$ can be rank deficient, which is one of the benefits of combining the local eigenvectors.
+$\mathcal{I}_{2}$ corresponds to the effect of heterogeneity.
+To distinguish, we call $\mathcal{H}$ the measure for heterogeneity and call $\mathcal{I}_{2}$ the measure for effective heterogeneity.
+Note that $\mathcal{H}$ and $\lambda_{{\min}}(Q)$ jointly decide $\mathcal{I}_{2}$.
+Therefore, if the local averages are noisy from the signal such that combining them across machines could hide the true signal, i.e., $\mathcal{H}$ is large, and $\lambda_{{\min}}(Q)$ is small, then the effective heterogeneity $\mathcal{I}_{2}$ would be large.
+On the other hand, if the local averages are far from the signal but combining them would enhance the signal, i.e., $\mathcal{H}$ is large, but $\lambda_{{\min}}(Q)$ can also be large such that their ratio is small, then the effective heterogeneity $\mathcal{I}_{2}$ would be small.
+We refer to the two simulated examples in Section C of the Supplementary Materials for more illustrations on the aforementioned phenomena.
+$\mathcal{I}_{1}$ and $\mathcal{I}_{2}$ can be regarded as the price needed to pay for the distributed computation.
 Moreover, $\mathcal{I}_{3}$ corresponds to the error of the centralized estimator $\hat{V}:=\frac{1}{L}\sum_{l}\overline{M}_{l}$.
-$\mathcal{I}_{1}$ and $\mathcal{I}_{2}$ can be regarded as the price needed to pay for the distributed computation.
-One of the contributions in Theorem 2 is that it explicitly explains how the heterogeneity $\mathcal{H}$’s effect could be a curse or blessing to $\mathcal{I}_{2}$ and possibly the final error bound.
-Note that $\mathcal{H}$ and $\lambda_{{\min}}(Q)$ jointly decide $\mathcal{I}_{2}$.
-Therefore, if the local $B_{l}$’s are noisy from the signal such that combining them could hide the true signal, i.e., $\mathcal{H}$ is large, and $\lambda_{{\min}}(Q)$ is small, then the heterogeneity would be a curse.
-On the other hand, if the local $B_{l}$’s are far from the signal but combining them would enhance the signal, i.e., $\mathcal{H}$ is large, but $\lambda_{{\min}}(Q)$ can also be large such that their ratio is small, then the heterogeneity would be a blessing.
-We refer to the two simulated examples in Appendix C for more illustrations on the aforementioned phenomena.
-To make the bounds in Theorem 2 more specific, in the following we provide Theorem 3 which establishes the estimation bounds of $\|\sum_{l}\overline{M}_{l}/L-Q\|_{2}$.
-Note that we assume that the privacy parameters $q,q^{\prime}$ relate to the privacy budget $\epsilon$ via (4).
-For illustration, we interchange the use of $q,q^{\prime}$, and $\epsilon$.
-Note that our results also hold for $q=\frac{\mathrm{e}^{\epsilon_{1}}}{1+\mathrm{e}^{\epsilon_{1}}}$ and $q^{\prime}=\frac{\mathrm{e}^{\epsilon_{2}}}{1+\mathrm{e}^{\epsilon_{2}}}$ with $\epsilon:=\max\{\epsilon_{1},\epsilon_{2}\}$ and $\epsilon_{1}$ and $\epsilon_{2}$ being of the same order.
+Before developing our main result, namely, the specific version of Theorem 2, we first provide Theorem 3 to study how the privacy parameters $q,q^{\prime}$ affect the upper bound for the centralized estimator $\|\sum_{l}\overline{M}_{l}/L-Q\|_{2}$.
+We need the following definition of $\xi(q,q^{\prime})$ and $C(q,q^{\prime})$ through out the remainder of this section to simplify the notation,
+$\xi(q,q^{\prime})=\frac{1-q^{\prime}}{(q+q^{\prime}-1)^{2}}\quad{\rm and}\quad C(q,q^{\prime})=\frac{q}{(q+q^{\prime}-1)^{2}},$ (12)
+where we note that when $q,q^{\prime}>1/2+\tau$ with a small constant $\tau$, $C(q,q^{\prime})$ can be regarded as a constant.
+We further use $\xi_{q,q^{\prime}}$ and $C_{q,q^{\prime}}$ to simplify the notation.
 Theorem 3
@@
 for some constants $10<\alpha_{1}<$ 1 and $0<\alpha_{2}\leq 1-\alpha_{1}$, the estimator $\overline{M}:=\frac{1}{L}\sum_{l}\overline{M}_{l}$ of $Q$ satisfies
-$\small\|\overline{M}-Q\|_{2}\leq C\max$$\bigg{\{}\frac{(q\rho+1-q^{\prime})^{1/2}n^{1/2}\rho\log^{1/2}(L+n)}{(q+q^{\prime}-1)\sqrt{L}},$
-$\quad\quad\quad\frac{(q\rho+1-q^{\prime})\log(L+n)}{(q+q^{\prime}-1)^{2}\sqrt{L}},\,\frac{((q+q^{\prime}-1)\rho+1-q^{\prime})^{2}}{(q+q^{\prime}-1)^{2}}\bigg{\}},$ (14)
+$\small\|\overline{M}-Q\|_{2}\lesssim\max$$\bigg{\{}\frac{(C_{q,q^{\prime}}\rho+\xi_{q,q^{\prime}})^{1/2}n^{1/2}\rho\log^{1/2}(L+n)}{\sqrt{L}},$
+$\quad\quad\quad\frac{(C_{q,q^{\prime}}\rho+\xi_{q,q^{\prime}})\log(L+n)}{\sqrt{L}},\,{(\rho+(q+q^{\prime}-1)\xi_{q,q^{\prime}})^{2}}\bigg{\}},$ (14)
 with probability larger than $1-O((L+n)^{-\nu})$ for some constant $\nu>0$.
+Remark 5
+Condition (13) implicitly assumes that $q,q^{\prime}$ relate to the privacy budget $\epsilon$ via (4).
+Nonetheless, the proof of Theorem 3 indicates that the result (3) also holds under (3), i.e., the more general region of $q,q^{\prime}$, with a modification of condition (13).
 Theorem 3 provides the error bound for the centralized estimator $\sum_{l}\overline{M}_{l}/L$.
-The RHS of (3) indicates that the privacy parameters $q$ and $q^{\prime}$ affect the accuracy of the estimated eigenvectors.
-In addition, the influence of $q$ and $q^{\prime}$ on the final bound is not equivalent.
-First, we observe that the privacy parameters $q,q^{\prime}$ appear in the numerators of the first two terms of (3) in the form $q\rho+1-q^{\prime}$.
-Considering that $\rho=o(1)$, we can see that a large level of $q^{\prime}$ is more desirable than a large level of $q$ for reducing the estimated error (3).
-In other words, perturbing 0 to 1 is more harmful than perturbing 1 to 0 for the model estimation.
-Second, when $\rho\leq C\frac{(q\rho+1-q^{\prime})^{1/2}}{(q+q^{\prime}-1)}\cdot\frac{\log^{1/2}(L+n)}{n^{1/2}}$, the RHS of (3) reduces to
-$\max\left\{\frac{(q\rho+1-q^{\prime})\log(L+n)}{(q+q^{\prime}-1)^{2}\sqrt{L}},\,\frac{((q+q^{\prime}-1)\rho+1-q^{\prime})^{2}}{(q+q^{\prime}-1)^{2}}\right\},$
-which matches the following bound without privacy-preserving up to log factors derived in Lei and Lin (2022),
-$\max\left\{\frac{\rho\log^{1/2}(L+n)}{\sqrt{L}},\,\rho^{2}\right\},$ (15)
-provided that $q$ and $q^{\prime}$ both approach to 1.
-The derivation of the bound for the single matrix $\overline{M}_{l}$ is similar to that in Theorem 3. For the sake of space, we summarize the results in Theorem B.1 in the Appendix.
-Now we are ready to specify the unknown terms in Theorem 2.
+When the privacy parameter $q^{\prime}$ is large in the sense that $\xi_{q,q^{\prime}}\lesssim C_{q,q^{\prime}}\rho$, the upper bound in (3) matches that in Lei and Lin (2022) up to log terms, where they did not impose the privacy-preserving constraint.
+Similar to Theorem 3, we can easily derive the upper bound for $\max_{i}\|\sum_{l\in S_{i}}(\overline{M}_{l}-Q_{l})/s\|_{2}$.
+Putting together the results, we provide in the following Theorem 4 the specific error bound for the estimated eigen-spaces by Algorithm 1.
+It is well-known that large network density will help to recover the communities.
+Therefore, in what follows, we focus on the more challenging regime with $n{\rho^{2}}/(C_{q,q^{\prime}}\rho+\xi_{q,q^{\prime}})\lesssim{{\log(s+n)}}$, though our results can also be extended to $n{\rho^{2}}/(C_{q,q^{\prime}}\rho+\xi_{q,q^{\prime}})\gtrsim{{\log(s+n)}}$.
+Theorem 4
+Suppose Assumptions 1-3 and condition (13) hold. If $n{\rho^{2}}/(C_{q,q^{\prime}}\rho+\xi_{q,q^{\prime}})\lesssim{{\log(s+n)}}$,
+$n\sqrt{s}\cdot\frac{\rho^{2}}{C_{q,q^{\prime}}\rho+\xi_{q,q^{\prime}}}\gtrsim{{\log(s+n)}};$ (15)
+$(q+q^{\prime}-1)\cdot\frac{\xi_{q,q^{\prime}}}{\rho}\lesssim n^{1/2},$ (16)
+then the output $\tilde{V}$ of Algorithm 1 satisfies
+$\mathrm{dist}(\tilde{V},V)\lesssim\mathcal{E}_{1}+\mathcal{E}_{2}+\mathcal{E}_{3};$
+$\mathcal{E}_{1}:=\max\bigg{\{}\frac{\log^{2}(s+n)(C_{q,q^{\prime}}\rho+\xi_{q,q^{\prime}})^{2}}{n^{2}{s}\rho^{4}},\;\frac{(1+(q+q^{\prime}-1)\xi_{q,q^{\prime}}/\rho)^{4}}{n^{2}}\bigg{\}},\;\mathcal{E}_{2}:=\frac{\mathcal{H}^{2}(n,\rho,L)}{n^{2}\rho^{4}};$
+$\mathcal{E}_{3}:=\max\bigg{\{}\frac{\log(L+n)(C_{q,q^{\prime}}\rho+\xi_{q,q^{\prime}})}{n{\sqrt{L}}\rho^{2}},\;\frac{(1+(q+q^{\prime}-1)\xi_{q,q^{\prime}}/\rho)^{2}}{n}\bigg{\}}$
+with probability larger than $1-O((L+n)^{-\nu})-O(m(s+n)^{-\nu}))$ for some constant $\nu>0$.
+In Theorem 4, $\mathcal{E}_{1},\mathcal{E}_{2}$ and $\mathcal{E}_{3}$ correspond to the squared local error, effective heterogeneity, and centralized error, respectively.
+The first two errors are the price to pay for the unavailability of full network layers.
+Condition (15) and (16) are required to make the results in Theorem 2 valid, which ensures $\mathcal{I}_{1}\lesssim 1$ in Theorem 2.
+$\mathcal{I}_{2}\lesssim 1$ is automatically met as we have shown in the proof.
+Several remarks are in order to explain the conditions and results in Theorem 4.
+Remark 6 (Effect of network sparsity)
+Condition (15) on $\rho$ depends on the number of network layers $s$ in each local machine.
+It reduces to $L^{1/2}n\rho\gtrsim\log(L+n)$ when $m=1$ and $\xi_{q,q^{\prime}}\lesssim C_{q,q^{\prime}}\rho$, which is identical to the condition in Theorem 1 of Lei and Lin (2022) up to log terms.
+It is worth mentioning that the squared local error $\mathcal{E}_{1}$ and the follow-up condition on $\rho$ is inherent in the distributed learning algorithm with only one round of communication.
+See Zhang et al.
+(2013); Lee et al.
+(2017) for example in the context of distributed empirical risk minimization and distributed sparse regression.
+In the context of SBMs, as far as we are aware, this work is the first result to obtain a regression-like result for the one-shot-based distributed algorithm.
+It is also worth mentioning that Arroyo et al.
+(2021) studied community detection on multi-layer networks, with each layer distributively stored in a separate machine.
+The requirement on $\rho$ turns out to be $\Omega(\log n/n)$, independent of network layers, where $n$ denotes the number of nodes.
+The condition on $\rho$ may be alleviated if we consider multi-round communications as in the context of regression; see Jordan et al.
+(2018); Chen et al.
+(2020), among others.
+The power method-based approach provides a possible basis Guo et al.
+(2024), which we left as our future work.
+Condition (15) and (16) both reveal that denser networks (i.e., larger $\rho$) can allow a more strict privacy-preserving level (i.e., small $\epsilon$).
+To see this, suppose $q,q^{\prime}$ relate to $\epsilon$ via (4), then (15) reduces to $\gamma(\epsilon):=\frac{(\mathrm{e}^{\epsilon}-1)^{2}}{\mathrm{e}^{\epsilon}+1}\gtrsim\frac{1}{\rho^{2}}\cdot\frac{\log(s+n)}{{ns^{1/2}}}$, where $\gamma(\epsilon)$ increases with $\epsilon$; (16) reduces to $\mathrm{e}^{\epsilon}-1\gtrsim{1}/{(\rho n^{1/2})}$.
+Remark 7 (Asymmetric roles of $(q,q^{\prime})$)
+From the definition of $C_{q,q^{\prime}}$ and $\xi_{q,q^{\prime}}$, it is undoubtedly that in general, the error bound of the eigenspace decreases when $q$ and $q^{\prime}$ values increase.
+Interestingly, we find that without a pre-specified privacy budget $\epsilon$, $q$ and $q^{\prime}$ play an asymmetric role in the error bound when the network is sparse with $\rho=o(1)$.
+The asymmetric role can be observed through the term $C_{q,q^{\prime}}\rho+\xi_{q,q^{\prime}}$.
+When $q,q^{\prime}>1/2+\tau$ for an arbitrarily small constant $\tau>0$, $C_{q,q^{\prime}}$ is a constant and $\xi_{q,q^{\prime}}\asymp 1-q^{\prime}$.
+Thus, $C_{q,q^{\prime}}\rho+\xi_{q,q^{\prime}}\asymp\rho+1-q^{\prime}$, so we see that larger $1-q^{\prime}$ would make the error bound larger, i.e., perturbing 0 to 1 is more harmful to the upper bound than perturbing 1 to 0.
+This interesting phenomenon can be explained as follows.
+We see that the sparsity level of the RR-perturbed network has the order of $\rho+1-q^{\prime}$, so perturbing 1 to 0 does not change the order of the network sparsity so it does not affect the error bound with a large sample.
+However, perturbing 0 to 1 can make the network denser if $1-q^{\prime}>\rho$, but the inflated “1” values are the wrong and harmful ones, making the error bound worse.
+Remark 8 (Optimal choice of $(q,q^{\prime})$)
+With a pre-specified specified privacy budget $\epsilon$, Theorem 4 indicates that $q=q^{\prime}=\frac{\mathrm{e}^{\epsilon}}{1+\mathrm{e}^{\epsilon}}$ is the best choice to yield the smallest statistical error bound for the eigenspace estimation.
+In particular, we can easily show the three terms $C_{q,q^{\prime}}$, $\xi_{q,q^{\prime}}$ and $(q+q^{\prime}-1)\xi_{q,q^{\prime}}$ involving $q,q^{\prime}$ all attain their minimum at $q=q^{\prime}=\frac{\mathrm{e}^{\epsilon}}{1+\mathrm{e}^{\epsilon}}$.
+For example, consider $\xi_{q,q^{\prime}}$, plugging in the upper boundaries shown in Figure 1, i.e., $q^{\prime}=1-\frac{q}{\mathrm{e}^{\epsilon}}$ when $q\leq\frac{\mathrm{e}^{\epsilon}}{1+\mathrm{e}^{\epsilon}}$ and $q^{\prime}=\mathrm{e}^{\epsilon}(1-q)$ when $q\geq\frac{\mathrm{e}^{\epsilon}}{1+\mathrm{e}^{\epsilon}}$, we have
+$\xi_{q,q^{\prime}}=\frac{1-q^{\prime}}{(q+q^{\prime}-1)^{2}}=\left\{\begin{array}[]{ll}\frac{1}{q\mathrm{e}^{\epsilon}(1-\mathrm{e}^{-\epsilon})^{2}}&\text{if }q\leq\frac{\mathrm{e}^{\epsilon}}{1+\mathrm{e}^{\epsilon}},\\ \frac{1-\mathrm{e}^{\epsilon}+\mathrm{e}^{\epsilon}q}{(1-q)^{2}(1-\mathrm{e}^{\epsilon})^{2}}&\text{if }q\geq\frac{\mathrm{e}^{\epsilon}}{1+\mathrm{e}^{\epsilon}},\end{array}\right.$
+which attains minimum at the $q=\frac{\mathrm{e}^{\epsilon}}{1+\mathrm{e}^{\epsilon}}$ and at this time $q^{\prime}$ equals $q$.
+The terms $C_{q,q^{\prime}}$ and $(q+q^{\prime}-1)\xi_{q,q^{\prime}}$ can be derived similarly.
+The next corollary shows that under certain extra conditions, the bound of our distributed estimator reduces to its centralized counterpart.
 Corollary 1
-Suppose Assumptions 1-4 and condition (13) hold. If
-$\rho\gtrsim\frac{(q\rho+1-q^{\prime})^{1/2}}{q+q^{\prime}-1}\cdot\sqrt{\frac{\log(n)}{n}}\quad{and}\quad\mathcal{H}\lesssim n\rho^{2},$ (16)
-then the output $\tilde{V}$ of Algorithm 1 satisfies
-$\mathrm{dist}(\tilde{V},V)\lesssim\mathcal{I}_{1}+\mathcal{I}_{2}+\mathcal{I}_{3}$ (17)
-with
-$\mathcal{I}_{1}:=\frac{q\rho+1-q^{\prime}}{(q+q^{\prime}-1)^{2}}\cdot\frac{\log n}{n\rho^{2}},\quad\mathcal{I}_{2}:=\frac{\mathcal{H}^{2}(n,\rho,L)}{n^{2}\rho^{4}},$ (18)
-and
-$\small\mathcal{I}_{3}:=\max$$\bigg{\{}\frac{(q\rho+1-q^{\prime})^{1/2}\log^{1/2}(L+n)}{(q+q^{\prime}-1)n^{1/2}\rho\sqrt{L}},\frac{(q\rho+1-q^{\prime})\log(L+n)}{(q+q^{\prime}-1)^{2}n\rho^{2}\sqrt{L}},\,\frac{((q+q^{\prime}-1)\rho+1-q^{\prime})^{2}}{(q+q^{\prime}-1)^{2}n\rho^{2}}\bigg{\}},$ (19)
-with probability larger than $1-O((L+n)^{-\nu})-O(Ln^{-\nu})$ for some constant $\nu>0$.
-To make the upper bound (17) valid, $\mathcal{I}_{1}$ and $\mathcal{I}_{2}$ both should be small enough.
-To this end, the technical condition (16) is utilized to meet the requirement in Theorem 2.
-The results in Corollary 1 can be simplified in various ways.
-For example, if we further require $\rho\gtrsim\frac{(q\rho+1-q^{\prime})^{1/2}}{q+q^{\prime}-1}\cdot\sqrt{\frac{\log(L+n)}{n}}$, then the first term in $\mathcal{I}_{3}$ dominate the second one.
-And the bound would be further reduced to
-$\frac{(q\rho+1-q^{\prime})^{1/2}\log^{1/2}(L+n)}{(q+q^{\prime}-1)n^{1/2}\rho\sqrt{L}}$ (20)
-if $\mathcal{I}_{2}$ is suppressed by $\mathcal{I}_{1}$ and $\mathcal{I}_{3}$, and $L\lesssim\frac{(q+q^{\prime}-1)^{2}}{q\rho+1-q^{\prime}}\cdot\frac{n\rho^{2}\log(L+n)}{\log^{2}n}$.
-As expected, we see in (20) that more local networks (large $L$) and less stringent privacy-preserving demand (large $q$ and $q^{\prime}$) lead to a smaller projection distance.
-With Corollary 1 at hand, it is convenient to derive the following misclassification error bound of ppDSC.
+Suppose the assumptions and conditions in Theorem 4 all hold. If we further have
+$m\lesssim\frac{n\rho^{2}}{(C_{q,q^{\prime}}\rho+\xi_{q,q^{\prime}})\log(s+n)}\cdot L^{1/2};$ (17)
+$\mathcal{H}\lesssim\max\bigg{\{}\frac{n^{1/2}\rho^{2}(C_{q,q^{\prime}}\rho+\xi_{q,q^{\prime}})^{1/2}\log^{1/2}(L+n)}{L^{1/4}\rho},\,{n^{1/2}\rho}(\rho+(q+q^{\prime}-1)\xi_{q,q^{\prime}})\bigg{\}},$ (18)
+then $\mathcal{E}_{3}$ dominates $\mathcal{E}_{1}$ and $\mathcal{E}_{2}$; see their definitions in Theorem 4. The output of $\tilde{V}$ of Algorithm 1 satisfies
+$\mathrm{dist}(\tilde{V},V)\lesssim\max\bigg{\{}\frac{\log(L+n)(C_{q,q^{\prime}}\rho+\xi_{q,q^{\prime}})}{n{\sqrt{L}}\rho^{2}},\;\frac{(1+(q+q^{\prime}-1)\xi_{q,q^{\prime}}/\rho)^{2}}{n}\bigg{\}}$
+with probability larger than $1-O((L+n)^{-\nu})-O(m(s+n)^{-\nu}))$ for some constant $\nu>0$.
+We make the following remarks on Corollary 1.
+Remark 9
+Condition (17) requires the number of machines should not be large given other parameters.
+In particular, when ${n}\rho^{2}/(C_{q,q^{\prime}}\rho+\xi_{q,q^{\prime}})\asymp{{\log(s+n)}}$, (17) reduces to $m\lesssim\sqrt{L}$, which is a crucial condition for the distributed estimator to match the error rate of its centralized counterpart in one-shot-based distributed learning literature, see Zhang et al.
+(2013); Lee et al.
+(2017), among others.
+The requirement for $m$ might be alleviated if one considers multi-round communications; see Jordan et al.
+(2018); Chen et al.
+(2020) in the context of regression.
+Condition (18) is mild.
+For example, suppose $\xi_{q,q^{\prime}}\lesssim C_{q,q^{\prime}}\rho$ and $\rho\asymp\frac{\log(s+n)m}{n\sqrt{L}}$ (i.e., the boundary in (17)), then (18) requires $\mathcal{H}\lesssim\frac{m^{2}\log^{1/2}(L+n)\log^{3/2}(s+n)}{nL}$.
+On the other hand, we naturally have $\mathcal{H}\lesssim n\rho^{2}\asymp\frac{m^{2}\log^{2}(s+n)}{nL}$ (see the proof of Theorem 4).
+Remark 10
+The error bound is comparable to Lei and Lin (2022) when $\xi_{q,q^{\prime}}\lesssim C_{q,q^{\prime}}\rho$.
+In particular, if $\rho\asymp\frac{\log(s+n)m}{n\sqrt{L}}$ and $\epsilon\asymp\log(\frac{n\sqrt{L}}{m})$ (suppose $q,q^{\prime}$ relates to $\epsilon$ via (4)), then $\mathcal{E}_{3}$ reduces to $\max\{\frac{\log\,(L+n)}{n\rho\sqrt{L}},\,\frac{1}{n}\}$, which is identical (up to log terms) to that in Lei and Lin (2022).
+Here, we remark that the large $\epsilon$ in the context of local-DP (Duchi et al., 2018) (i.e., DP at the level of original data rather than aggregated data) is reasonable and has also been used in other literature.
+We note that Mohamed et al.
+(2022) chose $\epsilon\gtrsim\log\,n$ when they studied the strong consistency of community detection algorithms under RR-perturbed SBMs.
+We also note that the work by Bhowmick et al.
+(2018) revisited the protections of local-DP (Duchi et al., 2018) and it is said that algorithms that only protect against the accurate reconstruction of (functions of) an individual’s data could allow a larger privacy budget ($\epsilon\gg 1$).
+The following theorem shows that the misclassification error bound for ppDSC turns out to be the squared error of the eigen-vectors $\tilde{V}$.
 Theorem 5
@@
 Denote the set of misclassified nodes by $\mathcal{M}$.
-Then, with probability larger $1-O((L+n)^{-\nu})-O(Ln^{-\nu})$ for some constant $\nu>0$, Algorithm 1 has misclassification rate upper bounded as
-$\frac{|\mathcal{M}|}{n}\leq(\mathcal{I}_{1}+\mathcal{I}_{2}+\mathcal{I}_{3})^{2}$
-with $\mathcal{I}_{1},\mathcal{I}_{2},\mathcal{I}_{3}$ defined in (18) and (19).
-Note that the misclassification error is just the squared error of the eigen-vectors $\tilde{V}$ in Corollary 1.
-Therefore, Theorem 4 can be similarly discussed and simplified.
-We omit the discussion for the sake of space.
-Finally, we discuss the necessity of the bias-adjustment procedure in Appendix B.
+Then, with probability larger $1-O((L+n)^{-\nu})-O(m(s+n)^{-\nu}))$ for some constant $\nu>0$, Algorithm 1 has misclassification rate upper bounded as $\frac{|\mathcal{M}|}{n}\lesssim\mathcal{E}_{3}^{2};$ see the definition of $\mathcal{E}_{3}$ in Theorem 4.
+Finally, we discuss the necessity of the bias-adjustment procedure in Section B of the Supplementary Materials.
 We investigate the deviation of the non-debiased matrix $\tilde{M}:=\frac{1}{L}\sum_{l}\frac{1}{n}\tilde{A}_{l}^{2}$ from population matrices $Q$ and $\tilde{Q}:=\frac{1}{L}\sum_{l}\left((q+q^{\prime}-1)P_{l}+(1-q^{\prime})\textbf{1}_{n}\textbf{1}_{n}^{\intercal}\right)^{2}/{n}$, respectively.
-It turns out that the bias-adjustment procedure can lead to a smaller error of eigenvector and misclassification, no matter which population matrix is used.
+It turns out that the bias-adjustment procedure can lead to a smaller eigenvector and misclassification error bound, no matter which population matrix is used.
 5 Simulations
 In this section, we evaluate the finite sample performance of the proposed algorithm.
-We show how the privacy parameters, the number of networks, and the number of nodes affect the finite sample behavior of the resulting estimates.
-In Appendix C, we test the effect of the heterogeneity on the performance of the proposed algorithm.
+We show how the number of networks, the number of nodes, and the number of local machines affect the finite sample behavior of the resulting estimates.
+We also justify our theory in terms of how to choose the edge-flipping probabilities in RR under a given privacy budget and how the network sparsity affects the performance under different privacy budgets.
+In Section C of the Supplementary Materials, we test the effect of heterogeneity and network sparsity on the performance of the proposed algorithm.
 Practical consideration.
@@
 Methods compared.
-Our algorithm is denoted by ppDSC. We compare our algorithm, ppDSC, with the following algorithms.
-  • ppDSC-1b: computed according to Algorithm 1 but only with one-step bias-correction (i.e., diagonal-debiasing) given in Lei and Lin (2022).
-Note that our bias-correction for the diagonal elements is not identical to Lei and Lin (2022);
-  • ppDSC-2b: computed according to Algorithm 1 but without any bias-correction, which corresponds to Charisopoulos et al. (2021);
-  • Oracle: a centralized algorithm.
-It is obtained from the spectral clustering based on the debiased sum of squared RR-perturbed matrices (i.e., $\overline{M}$ defined in (10)).
-Though violating the local computation constraint, this is used as a baseline algorithm.
+Our algorithm is denoted by ppDSC.
+We compare our algorithm with the following two distributed algorithms denoted by ppDSC-1b and ppDSC-2b, both of which follow the above practical restrictions.
+Hence, ppDSC-1b and ppDSC-2b suffer from the model misspecification problem.
+We also compare two baseline centralized algorithms.
+One is the centralized counterpart of the proposed algorithm, denoted by ppSC.
+The other is the algorithm in Lei and Lin (2022), denoted by Oracle because it knows the unperturbed network like an oracle.
+  • ppDSC-1b: computed according to Algorithm 1 but with line 6 replaced by one-step bias-correction (i.e., diagonal-debiasing) given in Lei and Lin (2022) (i.e., $\tilde{A}_{l}^{2}-\mbox{diag}(\tilde{A}_{l}^{2})$);
+  • ppDSC-2b: computed according to Algorithm 1 but without any bias-correction, which corresponds to Charisopoulos et al. (2021) with RR-perturbed networks as input;
+  • ppSC: obtained from the spectral clustering based on the debiased sum of squared RR-perturbed matrices (i.e., $\overline{M}$ defined in (9)).
+  • Oracle: the algorithm in Lei and Lin (2022); obtained from the spectral clustering based on the diagonal-debiased sum of squared original adjacency matrices (i.e., $\sum_{l}(A_{l}^{2}-{\rm diag}(A_{l}^{2})))$.
 Performance measures.
@@
 Similar to Lei and Lin (2022), we fix $K=3$ with balanced communities and set $B_{l}:=0.8B^{(1)}$ for $l=1,...,\lfloor L/2\rfloor$ and $B_{l}:=0.6B^{(2)}$ for $l=\lfloor L/2\rfloor+1,...,L$ with
-$\scriptsize B^{(1)}=W\begin{bmatrix}1.5&&\\ &0.2&\\ &&0.4\end{bmatrix}W^{\intercal}\approx\begin{bmatrix}0.62&0.22&0.46\\ 0.22&0.62&0.46\\ 0.46&0.46&0.85\end{bmatrix},$
-and
+$\scriptsize B^{(1)}=W\begin{bmatrix}1.5&&\\ &0.2&\\ &&0.4\end{bmatrix}W^{\intercal}\approx\begin{bmatrix}0.62&0.22&0.46\\ 0.22&0.62&0.46\\ 0.46&0.46&0.85\end{bmatrix};$
 $\scriptsize B^{(2)}=W\begin{bmatrix}1.5&&\\ &0.2&\\ &&-0.4\end{bmatrix}W^{\intercal}\approx\begin{bmatrix}0.22&0.62&0.46\\ 0.62&0.22&0.46\\ 0.46&0.46&0.85\end{bmatrix},$
@@
 $\scriptsize W=\begin{bmatrix}1/2&1/2&-\sqrt{2}/2\\ 1/2&1/2&\sqrt{2}/2\\ \sqrt{2}/2&-\sqrt{2}/2&0\end{bmatrix}.$
-We evaluate the effect of $q,q^{\prime}$, $n$ and $L$ on the numerical performance. The parameter set-up is as follows.
-  • Effect of $L$: $n=210$, $q=q^{\prime}=0.9$, and $L$ ranges from 2 to 20.
-  • Effect of $n$: $L=10$, $q=q^{\prime}=0.9$, and $n$ ranges from 15 to 300.
-  • Effect of $q,q^{\prime}$: $L=10$, $n=210$, and $q$ and $q^{\prime}$ are equal and range from 0.6 to 1.
-Results.
+Experiment 1.
+We evaluate the effect of $L$, $n$ and $m$ on the numerical performance. The parameter setup is as follows.
+  • Effect of $L$: $n=210$, $q=q^{\prime}=0.8$, $m=L$, and $L$ from 2 to 17.
+  • Effect of $n$: $L=m=12$, $q=q^{\prime}=0.8$, and $n$ ranges from 60 to 230.
+  • Effect of $m$: $L=12$, $n=210$, $q=q^{\prime}=0.8$, and $m$ varies in $\{2,4,6,12\}$.
 The averaged results over 20 replications are displayed in Figure 2.
 We have the following observations.
-First, as expected, as $q,q^{\prime}$, $n$, and $L$ increase, the performance of all estimators improve.
-Second, our estimator ppDSC is superior to the biased ppDSC-1b and ppDSC-2b except when the privacy parameters $q,q^{\prime}$ become close to 1 and the sample size $n$ is large.
-It is because that ppDSC-1b and ppDSC-2b ignore the bias caused by privacy-preserving the networks and thus mis-specify the model.
-When $q,q^{\prime}$ become close to 1, the bias from privacy preservation does not matter much, and then ppDSC-1b performs similarly to ppDSC.
-Third, as with extra local computation demand, our method ppDSC is generally not better than the Oracle estimator.
-However, when $q,q^{\prime}$, $n$, and $L$ become large, they are getting close.
+First, as expected, as $L,n$ increases and $m$ decreases, the performance of all estimators improves.
+Second, our estimator ppDSC is superior to the biased ppDSC-1b and ppDSC-2b.
+It is because ppDSC-1b and ppDSC-2b ignore the bias caused by privacy-preserving the networks and thus mis-specify the model.
+Third, as with extra local computation and privacy preservation demand, our method ppDSC is generally not better than the centralized methods ppSC and Oracle.
+However, when $L$ and $n$ become large, and $m$ becomes small, they are getting close.
 This is consistent with our theoretical results.
 (I) Projection distance
-(a) Effect of $q,q^{\prime}$
+(a) Effect of $L$
 (b) Effect of $n$
-(c) Effect of $L$
+(c) Effect of $m$
 (II) Misclassification rate
-(d) Effect of $q,q^{\prime}$
+(d) Effect of $L$
 (e) Effect of $n$
-(f) Effect of $L$
-Figure 1: Comparison of ppDSC, ppDSC-1b, ppDSC-2b and Oracle on the simulated data.
-The effect of the privacy parameters $q,q^{\prime}$, the number of nodes $n$, and the number of networks $L$ are shown, respectively.
+(f) Effect of $m$
+Figure 2: Comparison of ppDSC, ppDSC-1b, ppDSC-2b, ppSC and Oracle on the simulated data.
+The effect of the number of networks $L$, the number of nodes $n$, and the number of local machines $m$ are shown, respectively.
 The projection distance and misclassification rate are evaluated.
+Experiment 2.
+We evaluate the performance of ppDSC with the privacy budget $\epsilon=1$ fixed and the edge-fillping probabilities $q,q^{\prime}$ varying in the feasible region of DP shown in Figure 1.
+Other parameters are $n=210$ and $L=m=12$.
+The averaged misclassification rates over 10 replications under different couples of $q$ and $q^{\prime}$ are displayed in Figure 3, where the infeasible region is filled by pure white.
+When $q,q^{\prime}$ are getting close to the upper boundaries, the performance of ppDSC becomes better.
+In particular, the misclassification rate attains the empirical minimum at the position marked by the pink cross, which is quite close to the theoretically best choice of $q=q^{\prime}=\frac{{\rm e}^{\epsilon}}{1+{\rm e}^{\epsilon}}$ (see the discussions after Theorem 3) marked by the red star.
+Figure 3: The clustering performance of ppDSC with a pre-specified $\epsilon=1$ and $q,q^{\prime}$ varying in the feasible region of DP (see Figure 1).
+The empirically and theoretically best combination of $q,q^{\prime}$ are marked with red star and pink cross, respectively.
 6 Real data analysis
@@
 Hence, we regard the AUCS dataset as the original data without any perturbation.
-We will mimic the actual setup to see how the privacy and local computation affects the clustering performance of the proposed algorithm.
+We will mimic the actual setup to see how the privacy and local computation affect the clustering performance of the proposed algorithm.
 Data description.
@@
 The first three relationships were collected via a questionnaire.
-All the respondents answered all questions indicating that the data was complete.
+All the respondents answered all questions, indicating that the data was complete.
 And the last two relationships were acquired online via the DBLP database and Facebook.
@@
 Evidence of communities and heterogeneity.
-The topology of networks with respect to different relationships is shown in Figure 2 with colors indicating the research groups.
+The topology of networks with respect to different relationships is shown in Figure C.4 in the Supplementary Materials with colors indicating the research groups.
 It is clear to see that the network structures are quite different.
 To take a closer look at the community structures, we reorder the nodes according to their underlying research groups (i.e., nodes corresponding to the same research group are aligned next to each other).
-We further plot the corresponding network adjacency matrices in Figure 3, where the whole network (Figure 3(a)) corresponds to the network that contains the edges of all five relationships.
+We further plot the corresponding network adjacency matrices (see Figure C.5 in the Supplementary Materials), where the whole network (Figure C.5(a)) corresponds to the network that contains the edges of all five relationships.
 The whole network shows conspicuous block structures in that nodes in the same research group have more links than those across different research groups.
@@
 Distinct from the whole network, we see that the community structures of each network layer are heterogeneous in that some communities are not evident or even missing to different extents.
-(a) Work
-(b) Facebook
-(c) Leisure
-(d) Lunch
-(e) Coauthor
-Figure 2: A visualization of the AUCS network where vertex colors represent research groups. Each network corresponds to a specific relationship among the vertices.
 Practical consideration.
@@
 Methods compared.
-Consistent with the simulation experiments, we compare our method ppDSC with ppDSC-1b, ppDSC-2b and Oracle.
+Consistent with the simulation experiments, we compare our method ppDSC with ppDSC-1b, ppDSC-2b, ppSC and Oracle.
 The detailed description of each method can be found in Section 5.
-Generally, we study how the performance of each method alters with the network number $L$ and the privacy parameters $q$ and $q^{\prime}$.
-Regarding the effect of $L$, we mean that we use some of the five local networks to construct the estimator to see how the number of used networks affects the clustering performance.
-Regarding the effect of $q,q^{\prime}$, we consider three cases, including $q,q^{\prime}$ both varying, $q$ varying but $q^{\prime}$ being fixed, and $q^{\prime}$ varying being $q$ being fixed, where the latter two cases aim to verify the different effect of $q$ and $q^{\prime}$ indicated by the theory.
-The detailed setups are as follows.
-  • Effect of $L$: $q=q^{\prime}=0.9$, and $L$ ranges from 2 to 5.
-  • Effect of $q,q^{\prime}$: $L=5$, and $q$ and $q^{\prime}$ are equal and range from 0.6 to 1.
-  • Effect of $q$: $L=5$, $q^{\prime}=0.95$, and $q$ ranges from 0.6 to 1.
-  • Effect of $q^{\prime}$: $L=5$, $q=0.95$, and $q^{\prime}$ ranges from 0.6 to 1.
+We study how the performance of each method alters with the network number $L$ and the privacy parameters $q$ and $q^{\prime}$.
+Specifically, we fix $q(q^{\prime})$ to be 0.95 and vary $q^{\prime}(q)$ to verify the different effect of $q$ and $q^{\prime}$ indicated by the theory.
+To save space, the detailed descriptions on the experimental set-up is relegated to Section C of the Supplementary Materials.
 Results.
@@
 We have the following observations.
-First, as $L$ increases, the misclassification rate of all four methods decreases, showing the benefits of aggregating the information of local networks.
+First, as $L$ increases, the misclassification rate of all five methods decreases, showing the benefits of aggregating the information of local networks.
 Note that when $L=5$, a small drop in terms of the clustering performance appears, partially because the fifth network (i.e., the coauthor network) is rather sparse and has a weak signal of communities.
-Second, as expected, as $q$ and $q^{\prime}$ increase, the clustering performance of all four methods becomes better.
+Second, as expected, as $q$ and $q^{\prime}$ increase, the clustering performance of all four privacy-preserving methods becomes better.
 We note that the clustering performance with $q^{\prime}$ fixed to be 0.95 (See Figure 4(c)) is better than that with $q$ fixed to be 0.95 (See Figure 4(d)), which shows that large $q^{\prime}$ is more desirable than large $q$.
@@
 This is consistent with the theoretical results.
-Third, consistent with the simulation results, the proposed method ppDSC is better than the biased method ppDSC-1b and ppDSC-2b and is close to the baseline method Oracle as $q$ and $q^{\prime}$ increase.
-(a) Whole
-(b) Work
-(c) Facebook
-(d) Leisure
-(e) Lunch
-(f) Coauthor
-Figure 3: A visualization of the network adjacency matrices of the AUCS dataset.
-The nodes are permutated according to their underlying research groups.
-(a) corresponds to the network adjacency matrix that contains all the edges of the five relationships.
-(b)-(f) correspond to the adjacency matrices of the five relationships among vertices, respectively.
+Third, consistent with the simulation results, the proposed method ppDSC is better than the biased method ppDSC-1b and ppDSC-2b and is close to the baseline methods ppSC and Oracle as $q$ and $q^{\prime}$ increase.
 (a) Effect of $L$
@@
 The effect of the privacy parameters $q,q^{\prime}$ ($q,q^{\prime}$ synchronously vary, $q^{\prime}$ fixed but $q$ varies, $q$ fixed but $q^{\prime}$ varies), and the number of networks $L$ are shown, respectively.
-The misclassification rate with respect to the research group labels are evaluated.
+The misclassification rate with respect to the research group labels is evaluated.
 7 Conclusion
@@
 In the toolbox, a new algorithm named ppDSC has been developed, and it is shown to be both communication-efficient and privacy-aware.
-We also propose a two-step bias-adjustment procedure to remove the biases caused by the RR perturbation and take the square of the private-preserved network matrices.
+We also propose a two-step bias-adjustment procedure to remove the biases caused by the RR perturbation and taking the square of the debiased adjacency matrices.
 The procedure is found to be effective both theoretically and numerically.
 Moreover, we have rigorously investigated the misclassification rate of the algorithm.
-We show how the privacy parameters affect the error bound.
-In particular, we demonstrate the effect of heterogeneity among population networks and the asymmetry roles of the privacy parameters $q$ and $q^{\prime}$ on the estimation accuracy through rigorous theoretical justifications.
-The numerical experiments well support our theoretical findings.
+In particular, we show that when the number of machines and the effective heterogeneity are well-controlled, the misclassification rate of our distributed algorithm matches its centralized counterpart.
+More interestingly, our result indicates the best choice of privacy parameters under a given privacy budget $\epsilon$.
+In addition, we demonstrate the effect of heterogeneity among population networks and the asymmetrical roles of the privacy parameters $q$ and $q^{\prime}$ on the estimation accuracy.
+The numerical experiments support our theoretical findings well.
 The proposed methodology can be extended to a variety of scenarios.
@@
 (2019); Arroyo et al.
+(2021).
+The method for computing the local eigenspaces can also be replaced with the tensor-based method, e.g., Jing et al.
 (2021).
@@
 Fourth, we focus on the distributed learning framework with only one round of communication.
-It is desirable to extend the results to the distributed framework with multiple rounds, say the communication-efficient distributed power method (Li et al., 2021).
+It is desirable to extend the results to the distributed framework with multiple rounds, say the communication-efficient distributed power method (Guo et al., 2024).
 Last but not least, the method and theory for estimating $K$ can be an interesting topic to explore in the future.
@@
 Appendix
-Appendix A provides the details of the bias-adjustment procedure.
-Appendix B contains the additional theoretical results and discussions.
-Appendix C contains the additional simulation experiment.
-Appendix D contains the technical proofs for Theorems 1-4 and Corollary 1.
-Several auxiliary lemmas are provided in Appendix E.
+Section A provides the details of the bias-adjustment procedure.
+Section B includes the additional theoretical results and discussions.
+Section C presents the additional simulation experiments and the results of real data analysis.
+Section D contains the technical proofs.
+Several auxiliary lemmas are provided in Section E.
 Appendix A Details of bias-adjustment
@@
 More importantly, as we aim to approximate $\sum_{l=1}^{L}P_{l}^{2}$, this debiasing step is crucial when $L$ is increasing.
-Appendix B Additional theoretical results and discussions
-B.1 Bound for $\overline{M}_{l}$
-We provide the bound for the single matrix $\overline{M}_{l}$.
-As we will see, due to the absence of $L$, the condition and the dominating term in the resulting bound compared with the centralized bound is changed.
-The results are summarized in Theorem B.1.
-Theorem B.1
-Suppose Assumption 2 holds, and $\epsilon\lesssim{\log}\,n$, the estimator $\overline{M}_{l}$ for $Q_{l}$ satisfies
-$\max_{l}\|\overline{M}_{l}-Q_{l}\|_{2}\leq C\max\left\{\frac{(q\rho+1-q^{\prime})^{1/2}n^{1/2}\rho\log^{1/2}(n)}{q+q^{\prime}-1},\,\frac{(q\rho+1-q^{\prime})\log n}{(q+q^{\prime}-1)^{2}}\right\},$ (B.1)
-with probability larger than $1-O(Ln^{-\gamma})$, where $\gamma,C>0$ are some fixed constants.
-The proof of Theorem B.1 is included in Appendix D.
-B.2 The effect of bias-adjustment
+Appendix B The effect of bias-adjustment
 We theoretically show the effect of bias-adjustment. Denote the non-debiased estimator by
@@
 then the non-debiased estimator $\tilde{M}$ of $\tilde{Q}$ satisfies
-$\scriptsize\|\tilde{M}-\tilde{Q}\|_{2}\lesssim(q+q^{\prime}-1)^{2}\mathcal{I}$ (B.2)
+$\|\tilde{M}-\tilde{Q}\|_{2}\lesssim(q+q^{\prime}-1)^{2}\mathcal{I}$ (B.1)
 with probability larger than $1-O((L+n)^{-\nu})$ for some constant $\nu>0$. Moreover, with the same probability,
@@
 Here, we consider the two-parameter SBMs that $P_{l}$’s are homogeneous, with the inter-community and cross-community probabilities being $a$ and $b$, respectively.
-In this regime, the minimum non-zero eigen-value of $(q+q^{\prime}-1)P_{l}+(1-q^{\prime})\textbf{1}_{n}\textbf{1}_{n}^{\intercal}$ is identical to that of $(q+q^{\prime}-1)P_{l}$, which implies that the error for the eigen-vector (counterpart of Corollary 1) and the misclassification error (counterpart of Theorem 4) totally depend on (B.3).
+In this regime, the minimum non-zero eigen-value of $(q+q^{\prime}-1)P_{l}+(1-q^{\prime})\textbf{1}_{n}\textbf{1}_{n}^{\intercal}$ is identical to that of $(q+q^{\prime}-1)P_{l}$, which implies that the error for the eigen-vector (counterpart of Theorem 4) and the misclassification error (counterpart of Theorem 5) totally depend on (B.2).
 As a result, we can also see the accuracy gain in eigen-vector estimation and clustering.
 Appendix C Additional numerical analysis and results
+C.1 Effect of heterogeneity
 We conduct the following simulation to show the effect of heterogeneity.
 Network-generation.
-In Theorem 2, we show that the eigen-gap (i.e., $\lambda_{{\min}}(Q)$) divided by the heterogeneity of connectivity matrices (i.e., $\mathcal{H}$ in Assumption 4) can affect the estimation accuracy of the proposed algorithm theoretically.
+In Theorem 2, we show that the eigen-gap (i.e., $\lambda_{{\min}}(Q)$) divided by the heterogeneity of connectivity matrices (i.e., $\mathcal{H}$ in Definition 2) can affect the estimation accuracy of the proposed algorithm theoretically.
 The heterogeneity can lead to either accuracy gain or loss, depending on the model structure of the network.
-To illustrate this phenomenon numerically, we consider the following two models for demonstrating the curses of heterogeneity and the blessings of heterogeneity, respectively.
-  • Model I (Curses of heterogeneity).
-We fix $K=2$, $L=12$, $q,q^{\prime}=0.8$ and let $n$ vary from 20 to 400.
+To illustrate this phenomenon numerically, we consider the following two models for demonstrating the effect of heterogeneity.
+  • Model I.
+We fix $K=2$, $L=m=12$, $q,q^{\prime}=0.8$ and let $n$ vary from 20 to 400.
 The communities are balanced.
@@
     For this model, one can imagine that larger $\alpha$ indicates larger heterogeneity but a similar level of signal strength (eigen-gap).
-Therefore, larger $\alpha$ leads to larger scaled heterogeneity (i.e., eigen-gap divided by heterogeneity; see the discussions after Theorem 2 for the details).
+Therefore, larger $\alpha$ leads to larger effective heterogeneity (i.e., eigen-gap divided by heterogeneity; see the discussions after Theorem 2 for the details).
 This is verified in Figure C.1(a).
-  • Model II (Blessings of heterogeneity).
-We fix $K=2$, $L=12$, $q,q^{\prime}=0.8$ and let $n$ vary from 20 to 400.
+  • Model II.
+We fix $K=2$, $L=m=12$, $q,q^{\prime}=0.8$ and let $n$ vary from 20 to 400.
 The communities are balanced.
@@
     For this model, larger $\alpha$ indicates larger heterogeneity but stronger signal strength (eigen-gap).
-Thus, the composite effects lead to larger $\alpha$ corresponding to smaller scaled heterogeneity.
+Thus, the composite effects lead to larger $\alpha$ corresponding to smaller effective heterogeneity.
 See Figure C.2(a) for the illustration.
@@
 The results are consistent with the theoretical results given in Theorem 2.
-For Model I, we see that large heterogeneity (i.e., large $\alpha$) leads to inferior performance of ppDSC, especially in terms of the projection distance, which is partially because that large $\alpha$ leads to large scaled heterogeneity (i.e., eigen-gap divided by heterogeneity) in Theorem 2; see Figure C.1.
-For Model II, the results are on the opposite side in that large heterogeneity (i.e., large $\alpha$) leads to performance gain of ppDSC, which is partially due to the fact that large $\alpha$ leads to large eigen-gap and small scaled heterogeneity; see Figure C.2.
-In addition, our estimator ppDSC is better than the biased ppDSC-1b and ppDSC-2b, and is closer to the Oracle estimator.
+For Model I, we see that large heterogeneity (i.e., large $\alpha$) leads to inferior performance of ppDSC, especially in terms of the projection distance, which is partially because large $\alpha$ leads to large effective heterogeneity (i.e., eigen-gap divided by heterogeneity) in Theorem 2; see Figure C.1.
+For Model II, the results are on the opposite side in that large heterogeneity (i.e., large $\alpha$) leads to performance gain of ppDSC, which is partially due to the fact that large $\alpha$ leads to large eigen-gap and small effective heterogeneity; see Figure C.2.
+In addition, our estimator ppDSC is better than the biased ppDSC-1b and ppDSC-2b, and is closer to the baseline estimators ppSC and Oracle estimator.
 (a) Population parameters
@@
 (c) Misclassification rate
-Figure C.1: Illustration for the curse of heterogeneity (Model I).
+Figure C.1: Illustration for the negative effect of heterogeneity (Model I).
 (a) corresponds to population parameters including the heterogeneity, eigen-gap and their ratios against $\alpha$.
-(b) and (c) correspond to the projection distance and misclassification rate of ppDSC, ppDSC-1b, ppDSC-2b and Oracle.
+(b) and (c) correspond to the projection distance and misclassification rate of ppDSC, ppDSC-1b, ppDSC-2b, ppSC and Oracle.
 (a) Population parameters
@@
 (c) Misclassification rate
-Figure C.2: Illustration for the blessing of heterogeneity (Model II).
+Figure C.2: Illustration for the positive effect of heterogeneity (Model II).
 (a) corresponds to population parameters including the heterogeneity, eigen-gap and their ratios against $\alpha$.
-(b) and (c) correspond to the projection distance and misclassification rate of ppDSC, ppDSC-1b, ppDSC-2b and Oracle.
+(b) and (c) correspond to the projection distance and misclassification rate of ppDSC, ppDSC-1b, ppDSC-2b, ppSC and Oracle.
+C.2 Effect of network sparsity
+We evaluate the effect of network sparsity on the performance of ppDSC under different privacy budgets.
+To incorporate the network sparsity, we generate $B_{l}$’s by $B_{l}:=0.8\gamma B^{(1)}$ for $l=1,...,\lfloor L/2\rfloor$ and $B_{l}:=0.6\gamma B^{(2)}$ for $l=\lfloor L/2\rfloor+1,...,L$ with varying $\gamma$.
+Other parameters are $n=210$ and $L=m=12$.
+The averaged results over 20 replications under different $q$’s are displayed in Figure C.3.
+We assume $q,q^{\prime}$ relates to $\epsilon$ via $q=q^{\prime}=\frac{{\rm e}^{\epsilon}}{1+{\rm e}^{\epsilon}}$.
+Under a given privacy budget, a denser network yields better clustering performance.
+In other words, under a given network sparsity, a denser network permits a wider range of privacy budget to attain a specified misclassification rate or projection distance.
+(a) Projection distance
+(b) Misclassification rate
+Figure C.3: The performance of ppDSC with varying sparsity under different $q,q^{\prime}$’s.
+C.3 Experimental details and additional figures of real data analysis
+Figures C.4 and C.5 show the visualization of the AUCS dataset in terms of networks and matrices, respectively.
+(a) Work
+(b) Facebook
+(c) Leisure
+(d) Lunch
+(e) Coauthor
+Figure C.4: A visualization of the AUCS network where vertex colors represent research groups. Each network corresponds to a specific relationship among the vertices.
+(a) Whole
+(b) Work
+(c) Facebook
+(d) Leisure
+(e) Lunch
+(f) Coauthor
+Figure C.5: A visualization of the network adjacency matrices of the AUCS dataset.
+The nodes are permutated according to their underlying research groups.
+(a) corresponds to the network adjacency matrix that contains all the edges of the five relationships.
+(b)-(f) correspond to the adjacency matrices of the five relationships among vertices, respectively.
+We next provide the details of the methods compared in Section 6.
+Methods compared in Section 6.
+Consistent with the simulation experiments, we compare our method ppDSC with ppDSC-1b, ppDSC-2b and Oracle.
+The detailed description of each method can be found in Section 5.
+Generally, we study how the performance of each method alters with the network number $L$ and the privacy parameters $q$ and $q^{\prime}$.
+Regarding the effect of $L$, we mean that we use some of the five local networks to construct the estimator to see how the number of used networks affects the clustering performance.
+Regarding the effect of $q,q^{\prime}$, we consider three cases, including $q,q^{\prime}$ both varying, $q$ varying but $q^{\prime}$ being fixed, and $q^{\prime}$ varying being $q$ being fixed, where the latter two cases aim to verify the different effect of $q$ and $q^{\prime}$ indicated by the theory.
+The detailed setups are as follows.
+  • Effect of $L$: $q=q^{\prime}=0.9$, and $L$ ranges from 2 to 5.
+  • Effect of $q,q^{\prime}$: $L=5$, and $q$ and $q^{\prime}$ are equal and range from 0.6 to 1.
+  • Effect of $q$: $L=5$, $q^{\prime}=0.95$, and $q$ ranges from 0.6 to 1.
+  • Effect of $q^{\prime}$: $L=5$, $q=0.95$, and $q^{\prime}$ ranges from 0.6 to 1.
 Appendix D Main proofs
 D.1 Proof of Theorem 1
-For $i\in\{1,...,L\}$, denote the RR-perturbed $A_{i}$ by ${\tilde{A}}_{i}$.
+For $l\in\{1,...,L\}$, denote the RR-perturbed $A_{i}$ by ${\tilde{A}}_{i}$.
 By Proposition 1, for each $i$, ${\tilde{A}}_{i}$ satisfies $\epsilon$-edge-DP.
@@
 Otherwise, we could work with $VZ$ for some $Z\in\mathcal{O}_{K}$ instead.
-We will make use of the following result to prove, which is modified from Charisopoulos et al. (2021) using the notation of this paper.
+We will make use of the following result to prove, which is modified from Charisopoulos et al. (2021) using the notation of this work.
 Proposition 3
@@
 and
-$\|\overline{M}_{l}-Q\|_{2}\leq\frac{\delta}{8},\quad{for\;all\;}l=1,...,L.$ (D.2)
+$\|\sum_{l\in S_{i}}\overline{M}_{l}/(L/m)-Q\|_{2}\leq\frac{\delta}{8},\quad{for\;all\;}i=1,...,m.$ (D.2)
 Then the output $\tilde{V}$ of Algorithm 1 satisfies that
-${\mathrm{dist}}(\tilde{V},V)\lesssim\frac{1}{\delta^{2}}\max_{l}\|\overline{M}_{l}-Q\|_{2}^{2}+\frac{1}{\delta}\|\frac{1}{L}\sum_{l=1}^{L}\overline{M}_{l}-Q\|_{2}.$
+${\mathrm{dist}}(\tilde{V},V)\lesssim\frac{1}{\delta^{2}}\max_{i\in[m]}\|\sum_{l\in S_{i}}{\overline{M}_{l}}/{(L/m)}-Q\|_{2}^{2}+\frac{1}{\delta}\|\frac{1}{L}\sum_{l=1}^{L}\overline{M}_{l}-Q\|_{2}.$
 We first have that
-$\|\overline{M}_{l}-Q\|_{2}$ $\leq\|\overline{M}_{l}-Q_{l}\|_{2}+\|Q_{l}-Q\|_{2}$
-$\leq\|\overline{M}_{l}-Q_{l}\|_{2}+C^{\prime}\mathcal{H}(n,\rho,L),$
+$\max_{i\in[m]}\|\sum_{l\in S_{i}}{\overline{M}_{l}}/{L/m}-Q\|_{2}$ $\leq\max_{i\in[m]}\|\sum_{l\in S_{i}}{(\overline{M}_{l}-Q_{l})}/{(L/m)}\|_{2}+\max_{i\in[m]}\|\sum_{l\in S_{i}}{Q_{l}}/{(L/m)}-Q\|_{2}$
+$\lesssim\max_{i\in[m]}\|\sum_{l\in S_{i}}{(\overline{M}_{l}-Q_{l})}/{(L/m)}\|_{2}+\mathcal{H}(n,\rho,L),$
 where the last inequality follows from
-$\|Q_{l}-Q\|_{2}$ $=\left\|\frac{P_{l}^{2}}{n}-\frac{1}{L}\sum_{i=1}^{L}\frac{P_{i}^{2}}{n}\right\|_{2}\leq\frac{1}{n}\|\Delta\|_{2}^{2}\|B_{l}\Delta^{2}B_{l}-\frac{1}{L}\sum_{i=1}^{L}B_{i}\Delta^{2}B_{i}\|_{2}$
-$\leq C\frac{1}{K}\|B_{l}\Delta^{2}B_{l}-\frac{1}{L}\sum_{i=1}^{L}B_{i}\Delta^{2}B_{i}\|_{2}\leq C^{\prime}\mathcal{H}(n,\rho,L),$
+$\max_{i\in[m]}\|\sum_{l\in S_{i}}{Q_{l}}/{(L/m)}-Q\|_{2}$ $=\max_{i\in[m]}\left\|\frac{1}{L/m}\sum_{l\in S_{i}}\frac{P_{l}^{2}}{n}-\frac{1}{L}\sum_{l=1}^{L}\frac{P_{l}^{2}}{n}\right\|_{2}$
+$\leq\max_{i\in[m]}\frac{1}{n}\|\Delta\|_{2}^{2}\left\|\frac{1}{L/m}\sum_{l\in S_{i}}B_{l}\Delta^{2}B_{l}-\frac{1}{L}\sum_{l\in[L]}B_{l}\Delta^{2}B_{l}\right\|_{2}$
+$\lesssim\max_{i\in[m]}\frac{1}{K}\left\|\frac{1}{L/m}\sum_{l\in S_{i}}B_{l}\Delta^{2}B_{l}-\frac{1}{L}\sum_{l\in[L]}B_{l}\Delta^{2}B_{l}\right\|_{2}\asymp\mathcal{H}(n,\rho,L),$
 by using the fact that
 $P_{l}^{2}=\Theta B_{l}\Theta^{\intercal}\Theta B_{l}\Theta^{\intercal}=\bar{\Theta}\Delta(B_{l}\Delta^{2}B_{l})\Delta\bar{\Theta}^{\intercal}$
-with $\Delta:={\operatorname{diag}}(\sqrt{n_{1}},...,\sqrt{n_{K}})$, $\bar{\Theta}:=\Theta\Delta^{-1}$, and $\bar{\Theta}^{\intercal}\bar{\Theta}=\mathbb{I}$, and Assumptions 1 and 4.
+with $\Delta:={\operatorname{diag}}(\sqrt{n_{1}},...,\sqrt{n_{K}})$, $\bar{\Theta}:=\Theta\Delta^{-1}$, and $\bar{\Theta}^{\intercal}\bar{\Theta}=\mathbb{I}$, and Assumption 1 and Definition 2.
 To meet (D.1) and (D.2), we let $\delta\asymp\lambda_{\min}(Q)$ and enforce
-$\|\overline{M}_{l}-Q_{l}\|_{2}\leq\frac{\lambda_{\min}(Q)}{16}\quad{and}\quad\mathcal{H}\leq\frac{\lambda_{\min}(Q)}{16C^{\prime}},$
-which follows by our condition in Theorem 2. We can thus use the results in Proposition 3 to obtain that,
-$\mathrm{dist}(\tilde{V},V)\lesssim{\max_{l}\frac{\|\overline{M}_{l}-Q_{l}\|_{2}^{2}}{\lambda_{\min}^{2}(Q)}}+{\frac{\mathcal{H}^{2}(n,\rho,L)}{\lambda_{\min}^{2}(Q)}}+\frac{\|\sum_{l}\overline{M}_{l}/L-Q\|_{2}}{\lambda_{\min}(Q)}.$
+$\max_{i\in[m]}\|\sum_{l\in S_{i}}(\overline{M}_{l}-Q_{l})/{(L/m)}\|_{2}\lesssim\frac{\lambda_{\min}(Q)}{16}\quad{and}\quad\mathcal{H}\lesssim\frac{\lambda_{\min}(Q)}{16},$
+which follows by our condition in Theorem 2 that $\mathcal{I}_{1}$ and $\mathcal{I}_{2}$ are smaller than a small constant $c$. We can thus use the results in Proposition 3 to obtain that,
+$\mathrm{dist}(\tilde{V},V)\lesssim{\max_{l}\frac{\|\sum_{l\in S_{i}}(\overline{M}_{l}-Q_{l})/{(L/m)}\|_{2}^{2}}{\lambda_{\min}^{2}(Q)}}+{\frac{\mathcal{H}^{2}(n,\rho,L)}{\lambda_{\min}^{2}(Q)}}+\frac{\|\sum_{l}\overline{M}_{l}/L-Q\|_{2}}{\lambda_{\min}(Q)}.$
 The results of Theorem 2 are proved. $\blacksquare$
@@
 $\|E_{3}-\sum_{l}\frac{1}{L}\frac{G_{l}(q^{\prime})^{2}}{(q+q^{\prime}-1)^{2}}\|_{2}\leq C\frac{n((q+q^{\prime}-1)\rho+1-q^{\prime})^{2}}{(q+q^{\prime}-1)^{2}}.$
-Consequently, $E_{1}$, $E_{2}$ and $E_{3}$ are the main effects and we obtained the results of Theorem 3. $\blacksquare$
-D.4 Proof of Theorem B.1
-The proof is similar to the that of Theorem 3. We here go through the main steps.
-Recalling (A.2) and (11), we can write
-$n\overline{M}_{l}-nQ_{l}:=n\overline{M}_{l}-nP_{l}^{2}:=E_{l,0}+E_{l,1}+E_{l,2}+E_{l,3}-\frac{G_{l}(q^{\prime})^{2}}{(q+q^{\prime}-1)^{2}}.$
-Next, we bound these error terms, respectively.
-For $E_{l,0}:={-P_{l}}\cdot{\operatorname{diag}}(P_{l})-{\operatorname{diag}}(P_{l})\cdot P_{l}+({\operatorname{diag}}(P_{l}))^{2}$, we can easily show that
-$\|E_{l,0}\|_{2}\leq cn\rho^{2}.$
-For $E_{l,1}:=X_{l}(P_{l}-{\operatorname{diag}}(P_{l}))+(P_{l}-{\operatorname{diag}}(P_{l}))X_{l}$, using Lemma 1 with the same Bernstein tail condition as those in the proof of Theorem 3, we can obtain that with probability larger than $1-n^{-\beta}$ for some $\beta>0$,
-$\|E_{l,1}\|_{2}\leq c\frac{(q\rho+1-q^{\prime})^{1/2}}{q+q^{\prime}-1}n^{3/2}\rho\log^{1/2}(n),$
-provided that $\log^{1/2}\,n\leq C(q\rho+1-q^{\prime})^{1/2}n$, which can be implied by $\log n\leq Cn(1-q^{\prime})=C{n}/({1+{\mathrm{e}}^{\epsilon}})$ by noting our condition $\epsilon\leq C{\log}\,n$.
-For $E_{l,2}:=X_{l}^{2}-{\operatorname{diag}}(X^{2}_{l})$, using Lemma 2 with the same Bernstein tail condition as the one used in the proof of Theorem 3, we observe that with probability larger than $1-O(n^{-1})$,
-$\|E_{l,2}\|_{2}\leq c\max\left\{vn\log(n),\,\sqrt{v}\sqrt{n}\log^{3/2}(n),\,\log^{2}(n)\right\}/(q+q^{\prime}-1)^{2},$ (D.5)
-where $v:=q\rho+1-q^{\prime}$.
-Note that by the condition that $\epsilon\leq C{\log}\,n$, we have ${\log}\,n\leq Cn(1-q^{\prime})\leq Cn(q\rho+1-q^{\prime})$, and thus the first term in (D.5) dominates and
-$\|E_{l,2}\|_{2}\leq c\frac{(q\rho+1-q^{\prime})n\log(n)}{(q+q^{\prime}-1)^{2}}.$
-For the last two error terms $E_{l,3}-\frac{G_{l}(q^{\prime})^{2}}{(q+q^{\prime}-1)^{2}}$, it is easy to see that
-$\|E_{l,3}-\frac{G_{l}(q^{\prime})^{2}}{(q+q^{\prime}-1)^{2}}\|_{2}\leq c\frac{(1-q^{\prime}+(q+q^{\prime}-1)\rho)^{2}}{(q+q^{\prime}-1)^{2}}\cdot n.$
-It is easy to see that $E_{l,0}\leq CE_{l,1}$. Now we show that $E_{3}\leq CE_{2}$, for which we only need to show
-$(1-q^{\prime}+(q+q^{\prime}-1)\rho)^{2}\leq C(q\rho+1-q^{\prime}){\log}\,n.$
-Actually, we have $(1-q^{\prime})^{2}\leq C\log\,n(1-q^{\prime})\leq C\log\,n(q\rho+1-q^{\prime})$, and $((q+q^{\prime}-1)\rho)^{2}\leq Cq\rho\log\,n\leq C(q\rho+1-q^{\prime}){\log}\,n$ by noting that $q>1/2$ and $q+q^{\prime}-1\leq 1$.
-As a result, $E_{l,1}$ and $E_{l,2}$ dominate.
-After using the union bound, we finally have
-$\max_{l}\|\overline{M}_{l}-Q_{l}\|_{2}\leq C\max\left\{\frac{(q\rho+1-q^{\prime})^{1/2}n^{1/2}\rho\log^{1/2}(n)}{q+q^{\prime}-1},\,\frac{(q\rho+1-q^{\prime})\log n}{(q+q^{\prime}-1)^{2}}\right\},$
-with probability larger than $1-O(Ln^{-\gamma})$ for some constant $\gamma>0$. The results of Theorem B.1 are arrived. $\blacksquare$
-D.5 Proof of Corollary 1
-The proof of Corollary 1 follows directly from that of Theorem 2.
+Consequently, $E_{1}$, $E_{2}$ and $E_{3}$ are the main effects. Noting (12), we obtain the results of Theorem 3. $\blacksquare$
+D.4 Proof of Theorem 4
+The proof of Theorem 4 follows from that of Theorem 2.
 First, we need to show that $Q$ is of rank $K$ and provide the lower bound of $\lambda_{K}(Q)$.
@@
 (D.4) implies that $Q$’s rank is $K$ and $\lambda_{K}(Q)\geq cn\rho^{2}$.
-Next, we specify the $\mathcal{I}_{1},\mathcal{I}_{2}$ and $\mathcal{I}_{3}$ and the corresponding condition in Theorem 2 that $\mathcal{I}_{1}$ and $\mathcal{I}_{2}$ need to be smaller enough.
-Recall the bound for the single matrix $\overline{M}_{l}$ in Theorem B.1, we find that when
-$\rho\gtrsim\frac{(q\rho+1-q^{\prime})^{1/2}}{q+q^{\prime}-1}\cdot\sqrt{\frac{\log(n)}{n}},$ (D.7)
-the bound in (B.1) reduces to the first term $\frac{(q\rho+1-q^{\prime})^{1/2}n^{1/2}\rho\log^{1/2}(n)}{q+q^{\prime}-1}$ which is automatically smaller than $\lambda_{K}(Q)$ provided with (D.7).
-As a result, $\mathcal{I}_{1}$ is specified under (D.7).
-For $\mathcal{I}_{2}$, we do not specify it but only require the condition that
-$\mathcal{H}\lesssim n\rho^{2}.$
-For $\mathcal{I}_{3}$, the bound directly follows from Theorem 3.
-As a result, the results of Corollary 1 have been arrived. $\blacksquare$
+Next, we specify the upper bounds for $\mathcal{I}_{1},\mathcal{I}_{2}$ and $\mathcal{I}_{3}$ and the corresponding condition in Theorem 2 that $\mathcal{I}_{1}$ and $\mathcal{I}_{2}$ need to be smaller enough.
+Similar to the proof of Theorem 3, we can show that with probability larger than $1-O(m(s+n)^{-\nu})$ for some positive constant $\nu$,
+$\small\max_{i}\|\sum_{l\in S_{i}}(\overline{M}_{l}-Q_{l})/|S_{i}|\|_{2}$ $\leq C\max\bigg{\{}\frac{(q\rho+1-q^{\prime})^{1/2}n^{1/2}\rho\log^{1/2}(s+n)}{(q+q^{\prime}-1)\sqrt{s}},$
+$\quad\quad\frac{(q\rho+1-q^{\prime})\log(s+n)}{(q+q^{\prime}-1)^{2}\sqrt{s}},\,\frac{((q+q^{\prime}-1)\rho+1-q^{\prime})^{2}}{(q+q^{\prime}-1)^{2}}\bigg{\}}.$ (D.6)
+When
+$\rho\lesssim\sqrt{\frac{\log(s+n)}{n}}\cdot\frac{(q\rho+1-q^{\prime})^{1/2}}{q+q^{\prime}-1},$ (D.7)
+the second term in the RHS of (D.4) dominates the first one. As a result,
+$\mathcal{I}_{1}$ $:=\frac{\max_{i}\|\sum_{l\in S_{i}}\overline{M}_{l}/|S_{i}|-Q\|_{2}^{2}}{\lambda_{\min}^{2}(Q)}$
+$\lesssim\max\bigg{\{}\frac{(q\rho+1-q^{\prime})^{2}}{(q+q^{\prime}-1)^{4}}\cdot\frac{\log^{2}(s+n)}{n^{2}\rho^{4}s},\,\frac{((q+q^{\prime}-1)\rho+1-q^{\prime})^{4}}{(q+q^{\prime}-1)^{4}n^{2}\rho^{4}}\bigg{\}}:=\mathcal{E}_{1}.$
+It is then easy to see that
+$n\sqrt{s}\cdot\frac{\rho^{2}(q+q^{\prime}-1)^{2}}{q\rho+1-q^{\prime}}\gtrsim{{\log(s+n)}}$
+and
+$\frac{(q+q^{\prime}-1)\rho+1-q^{\prime}}{\rho(q+q^{\prime}-1)}\lesssim n^{1/2},$
+are sufficient conditions for $\mathcal{I}_{1}\lesssim 1$.
+For $\mathcal{I}_{2}:=\frac{\mathcal{H}^{2}(n,\rho,L)}{n^{2}\rho^{4}}$, it can be implied by Assumptions 1, 2 and Definition 2 that, $\mathcal{H}\lesssim n\rho^{2}$, and thus $\mathcal{I}_{2}=\mathcal{E}_{2}\lesssim 1$ naturally.
+For $\mathcal{I}_{3}$, we have by Theorem 3 and (D.7) that,
+$\small\mathcal{I}_{3}\lesssim\max\bigg{\{}\frac{(q\rho+1-q^{\prime})\log(L+n)}{(q+q^{\prime}-1)^{2}\sqrt{L}n\rho^{2}},\,\frac{((q+q^{\prime}-1)\rho+1-q^{\prime})^{2}}{(q+q^{\prime}-1)^{2}n\rho^{2}}\bigg{\}}:=\mathcal{E}_{3}.$
+Finally, noting (12), we obtain the results of Theorem 4. $\blacksquare$
+D.5 Proof of Corollary 1
+Following Theorem 4, we only need to compare $\mathcal{E}_{1}$ and $\mathcal{E}_{2}$ with $\mathcal{E}_{3}$. When
+$m\lesssim\frac{n\rho^{2}(q+q^{\prime}-1)^{2}}{(q\rho+1-q^{\prime})\log(s+n)}\cdot L^{1/2},$
+the first term in $\mathcal{E}_{3}$ dominates the first term in $\mathcal{E}_{1}$.
+The second term in $\mathcal{E}_{3}$ dominates the second term in $\mathcal{E}_{1}$ because the former is just the squared version of the latter and they are both of smaller order than 1.
+And when
+$\mathcal{H}\lesssim\max\bigg{\{}\frac{n^{1/2}\rho\,(q\rho+1-q^{\prime})^{1/2}\log^{1/2}(L+n)}{(q+q^{\prime}-1)L^{1/4}},\,\frac{n^{1/2}\rho\,((q+q^{\prime}-1)\rho+1-q^{\prime})}{q+q^{\prime}-1}\bigg{\}},$
+$\mathcal{E}_{3}$ dominates $\mathcal{E}_{2}$. Hence we obtain the results of Theorem 4. $\blacksquare$
 D.6 Proof of Theorem 5
@@
 $=4\|\tilde{V}-VO\|_{\tiny F}^{2}$
-$\leq C\|\tilde{V}-VO\|_{2}^{2}\leq C\mathcal{S}^{2},$
-where $\mathcal{S}:=\mathcal{I}_{1}+\mathcal{I}_{2}+\mathcal{I}_{3}$ with $\mathcal{I}_{i}$’s defined in Corollary 1, the first inequality follows because we assume that ${\tilde{V}^{\prime}}$ is the global solution minimum of the following $k$-means objective and $VO$ is a feasible solution by Lemma 4,
+$\leq C\|\tilde{V}-VO\|_{2}^{2}\leq C\mathcal{E}_{3}^{2},$
+where $\mathcal{E}_{3}$ is defined in Theorem 4, the first inequality follows because we assume that ${\tilde{V}^{\prime}}$ is the global solution minimum of the following $k$-means objective and $VO$ is a feasible solution by Lemma 4,
 $(\tilde{\Theta},X)=\underset{{{\Theta}\in\mathbb{M}_{n,K},{X}\in\mathbb{R}^{K\times K}}}{{\mathrm{arg\;min}}}\;\|{\Theta}{X}-\tilde{V}\|_{\tiny F}^{2},$
-and the last inequality follows from Corollary 1 and the fact that $\min_{O\in\mathcal{O}_{K}}\|\tilde{V}-VO\|_{2}^{2}\lesssim\mathrm{dist}(\tilde{V},V)$.
+and the last inequality follows from Theorem 4 and the fact that $\min_{O\in\mathcal{O}_{K}}\|\tilde{V}-VO\|_{2}^{2}\lesssim\mathrm{dist}(\tilde{V},V)$.
 Then, we begin to bound the fraction of misclassified nodes. Define
@@
 $\|\hat{U}-UO\|_{F}\leq\frac{2\sqrt{2K}\|M-Q\|_{2}}{\gamma_{n}}.$
-Proof. See Lei and Rinaldo (2015); Chen et al. (2021b), among others. $\blacksquare$
+Proof. See Lei and Rinaldo (2015); Chen et al. (2021), among others. $\blacksquare$
 Lemma 4